### v1

# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

Abstract
Multi-objective optimization (MOO) in deep learning aims to simultaneously optimize multiple conflicting objectives, a challenge frequently encountered in areas like multi-task learning and multi-criteria learning. Recent advancements in gradient-based MOO methods have enabled the discovery of diverse types of solutions, ranging from a single balanced solution to finite or even infinite Pareto sets, tailored to user needs. These developments have broad applications across domains such as reinforcement learning, computer vision, recommendation systems, and large language models. This survey provides the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theories, and practical applications. By unifying various approaches and identifying critical challenges, it serves as a foundational resource for driving innovation in this evolving field. A comprehensive list of MOO algorithms in deep learning is available at https://github.com/Baijiong-Lin/Awesome-Multi-Objective-Deep-Learning.
Index Terms:
Multi-Objective Optimization, Multi-Task Learning, Pareto Set Learning, Deep Learning
1 Introduction
Multi-Objective Optimization (MOO) [1], which aims to optimize multiple potentially conflicting objectives simultaneously, has become an important topic in deep learning. A multi-objective optimization problem can be formulated as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$ (1)
where $m\geq 2$ is the number of objectives, ${\bm{\theta}}$ is the decision variable and ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable.
Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives. Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex. MOO is expected to achieve a solution that matches the user-defined preference. Hence, compared to single-objective optimization, MOO is more practical for complex real-world applications.
Originally arising from operations research, MOO has been extensively studied across various disciplines of science and engineering [2]. In recent years, there has been a surge of interest in applying MOO techniques to deep learning, leading to the reformulation or extension of many learning paradigms as MOO problems.
In deep learning, there are generally two primary types of MOO scenarios. One scenario is multi-task learning [3], where a model is trained to handle multiple tasks simultaneously. For example, in computer vision, we may desire a single network to perform various dense prediction tasks concurrently, such as segmentation, depth estimation, and surface normal prediction [4]. In reinforcement learning [5], we might aim for a model that performs well on several robotic manipulation tasks. In the training of Large Language Models (LLMs), we expect the model to excel at different tasks, such as mathematics, reasoning, language translation, and summarization. However, these tasks often compete for shared network resources, leading to conflicts that can degrade performance on individual tasks. By formulating the problem as an MOO, where the performance on each task is treated as a separate objective [6], we can systematically manage these conflicts and optimize all task objectives simultaneously.
Another scenario involves incorporating additional evaluation criteria, such as fairness, safety, harmfulness mitigation, and efficiency, into learning paradigms. Considering these objectives is essential for practical real-world applications but may conflict with the primary goal of maximizing performance. For example, in LLMs training [7], we aim for models that not only perform well on tasks like language understanding but also adhere to safety standards, avoid producing harmful content, and ensure fairness across different users. In hardware-aware Neural Architecture Search (NAS) [8], we aim to discover network architectures that achieve high accuracy while also being efficient in terms of computational resources, energy consumption, or parameter size. However, reducing model size or computational complexity often leads to decreased performance. By treating such problems as MOO, where each criterion is an objective, we provide a structured way to explore these trade-offs.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [Gradient-based Multi-objective Deep Learning, fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al. [12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SDMGrad [20]; IMTL-G [11]; Nash-MTL [21]; FairGrad [22]; Aligned-MTL [23]; GradNorm [24]; DB-MTL [12], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [25]; GradVac [26]; GradDrop [27], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Preference Vector-based Methods (§4.1), finite [PMTL [28]; EPO [29, 30]; WC-MGDA [31]; PMGDA [32]; FERERO [33]; GMOOAR [34]; UMOD [35], finite_work] ] [Preference Vector-free Methods (§4.2), finite [HVGrad [36, 37, 38]; MOO-SVGD [39]; F4M [40]; SoM [41]; MosT [42], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [43]; CPMTL [44]; PHN-HVI [45]; SUHNPF [46]; Tuan et al. [47], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [48]; YOTO [49]; Raychaudhuri et al. [50], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [51]; LORPMAN [52]; Chen and Kwok [53]; Panacea [54]; Tang et al. [55]; PaLoRA [56], infinite_work] ] ] ]
Figure 1: Taxonomy of existing MOO algorithms in deep learning.
1.1 Taxonomy
Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches based on the goal:
  • When the goal is to obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization. These methods focus on appropriately adjusting the contributions of each objective to achieve a solution that is not overly biased toward any one objective.
  • When the goal is to provide a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously. This allows users to select from a set of solutions that best meet their specific needs.
  • When the goal is to generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones. Algorithms in this category propose efficient structures to learn this subspace. During training, preference vectors are randomly sampled in each iteration. Once a preference is sampled, the optimization reduces to solving for either a single solution or a finite set, making it possible to use techniques from the first two approaches.
Figure 1 illustrates the taxonomy of current MOO algorithms in deep learning, serving as a guide to the methods discussed in this paper. By exploring these avenues, MOO provides a comprehensive framework for addressing complex optimization problems in deep learning, enabling the development of models that can effectively balance multiple competing objectives.
1.2 Comparison with Related Surveys
Some surveys have attempted to cover aspects of MOO. For example, Zhang and Yang [3] provide a comprehensive overview of traditional multi-task methods within the machine learning domain. Crawshaw [57] offers a summary of multi-task methods in deep learning up to the year 2020. Yu et al. [58] present a comprehensive survey that covers multi-task learning from traditional machine learning to deep learning. Eichfelder [59] discusses advancements in MOO from a general optimization perspective, without specific emphasis on deep learning. Wei et al. [60] focus on evolutionary MOO algorithms, but these algorithms are usually not suitable for deep neural networks due to their large parameter spaces and computational demands. Peitz and Hotegni [61] survey the MOO algorithms for deep learning but cover only a limited selection of gradient-based MOO methods and do not include theory and applications. Other surveys concentrate on applications of MOO in areas such as dense prediction tasks [4], recommend systems [62], and natural language processing [63].
In contrast, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications. To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
1.3 Organization of the Survey
The notations used in this survey are summarized in Table I. The rest of this survey is organized as follows:
  • Section 2 presents preliminaries in MOO, including key definitions and concepts;
  • Section 3 discusses methods for finding a single Pareto optimal solution, covering loss balancing and gradient balancing approaches;
  • Sections 4 and 5 focus on methods for identifying a finite and an infinite set of Pareto optimal solutions, respectively;
  • Section 6 delves into the theoretical analysis of convergence and generalization in gradient-based multi-objective optimization algorithms;
  • Section 7 showcases various applications in deep learning, including reinforcement learning, Bayesian optimization, computer vision, and large language models;
  • Section 8 offers useful resources for multi-objective deep learning, such as datasets and software tools;
  • Section 9 highlights challenges in the field and suggests promising directions for future research;
  • Section 10 summarizes this survey.
TABLE I: Summary of Notations.
Notation Description
${\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}$ Decision variable ${\bm{\theta}}$ in feasible set $\mathcal{K}$ with dimension $d$.
$m$ Number of objectives.
$n$ Number of finite Pareto solutions.
$K$ Number of iterations.
${\bm{f}}=[f_{1},\cdots,f_{m}]$ Objective function.
${\bm{\lambda}}=[\lambda_{1},\cdots,\lambda_{m}]$ Objective weight vector.
${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]$ Preference vector.
${\bm{z}}^{*}=[z_{1}^{*},\cdots,z_{m}^{*}]$ Ideal objective value.
${\bm{\theta}}^{(k)}$ Solution at $k$-th iteration.
${\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}$ Solution $1$ to Solution $n$ in a size-$n$ solution set.
${\bm{g}}^{(k)}_{i}$ The gradient vector of $i$-th objective at $k$-th iteration.
${\bm{G}}^{(k)}=[{\bm{g}}_{1}^{(k)},\cdots,{\bm{g}}_{m}^{(k)}]$ Gradient matrix at $k$-th iteration.
${\bm{d}}^{(k)}$ Updating direction of ${\bm{\theta}}$ at $k$-th iteration.
${\bm{\phi}}$ Parameters of Pareto set learning structures.
$\mathrm{HV}_{\bm{r}}(\cdot)$ Hypervolume with respect to reference point ${\bm{r}}$.
$[m]$ Index set $\{1,\cdots,m\}$.
$\Delta_{m-1}$ $(m-1)$-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
[…]
$\eta$ Step size for updating ${\bm{\theta}}$.
$\epsilon$ Error tolerance.
2 Preliminary: Key Concepts in MOO
In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to solve MOO problems in Section 2.2.
2.1 Pareto Concepts
Unlike single-objective optimization, solutions in MOO cannot be directly compared based on a single criterion, but are compared using the concept of dominance as follows. Note that without loss of generality, we consider multi-objective minimization (i.e., problem (1)) here.
Definition 1 ((strict) Pareto dominance [1])
A solution ${\bm{\theta}}^{(a)}$ dominates another solution ${\bm{\theta}}^{(b)}$ (denoted as ${\bm{\theta}}^{(a)}\preceq{\bm{\theta}}^{(b)}$) if and only if $f_{i}({\bm{\theta}}^{(a)})\leq f_{i}({\bm{\theta}}^{(b)})$ for all $i\in[m]$, and there exists at least one $i\in[m]$ such that $f_{i}({\bm{\theta}}^{(a)})<f_{i}({\bm{\theta}}^{(b)})$. A solution ${\bm{\theta}}^{(a)}$ strictly dominates another solution ${\bm{\theta}}^{(b)}$ if and only if $f_{i}({\bm{\theta}}^{(a)})<f_{i}({\bm{\theta}}^{(b)})$ for all $i\in[m]$.
Based on this definition, we further define Pareto optimality (PO), Pareto set, and Pareto front as follows.
Definition 2 ((weak) Pareto optimality [1])
A solution ${\bm{\theta}}^{*}$ is Pareto optimal if no other solution dominates it. A solution ${\bm{\theta}}^{*}$ is weakly Pareto optimal if no other solution strictly dominates it.
Definition 3 (Pareto set (PS) [1])
A PS is the set of all Pareto optimal solutions.
Definition 4 (Pareto front (PF) [1])
A PF is the set of all objective function values of the PO solutions.
Figure 2 illustrates the aforementioned definitions. The blue circles represent Pareto optimal solutions, the yellow circles represent weakly Pareto optimal solutions, and the red circles are solutions dominated by at least one other solution. The blue curve represents the Pareto front.
Figure 2: Illustration of Pareto optimality concepts. The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively. The blue curve represents the Pareto front.
Definition 5 (Preference vector)
${\bm{\alpha}}$ denotes a preference vector. The value of the entries of ${\bm{\alpha}}$ denote the preference of a specific objective. Throughout this paper, a preference vector ${\bm{\alpha}}$ is constrained on a simplex $\Delta_{m-1}$, where $\Delta_{m-1}=\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
In deep learning, the parameter ${\bm{f}}({\bm{\theta}})$ is typically optimized using gradient descent. In single-objective optimization, when the objective function is non-convex, the optimization process often reaches a stationary point. In the case of MOO, the solution generally converges to a Pareto stationary solution, which is defined as follows:
Definition 6 (Pareto stationary [64])
A solution ${\bm{\theta}}^{*}$ is called Pareto stationary if there exists a vector ${\bm{\lambda}}\in\Delta_{m-1}$ such that $\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}({\bm{\theta}}^{*})\|=0$.
Pareto stationarity is a necessary condition for achieving Pareto optimality. If all objectives are convex with $\lambda_{i}>0,\forall i$, it also serves as the Karush-Kuhn-Tucker (KKT) sufficient condition for Pareto optimality [64].
When evaluating the performance of a set of obtained solutions, the Hypervolume (HV) [65] indicator is one of the most widely used performance indicators, which is defined as follows.
Definition 7 (Hypervolume [65])
Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\cdots,{\bm{q}}^{(N)}\}$ and a reference point ${\bm{r}}$, the hypervolume of ${\mathbb{S}}$ is calculated by:
$\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$ (2)
where $\Lambda(\cdot)$ denotes the Lebesgue measure of a set.
HV provides a unary measure of both convergence and diversity of objective vectors. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
2.2 Scalarization Methods
Linear scalarization. The most straightforward approach to solving MOO problems (i.e., problem (1)) is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$ (3)
Linear scalarization is widely used in practice due to its simplicity. However, this method suffers from a significant limitations: It can only identify solutions on the convex part of the PF. For PFs with a concave shape, it typically finds only the endpoints.
Tchebycheff scalarization. Another popular scalarization function is Tchebycheff scalarization [66, 67], which is defined as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$ (4)
where $z_{i}^{*}$ are reference values, which can be set as the ideal value of the $i$-th objective. Unlike linear scalarization, Tchebycheff scalarization can explore the entire PF by traversing all preference vectors within the preference simplex. This ensures that both convex and concave regions of the PF are covered. The relationship between Tchebycheff scalarization and weak Pareto optimality can be formally described as follows:
Theorem 1 ([68])
A solution ${\bm{\theta}}^{*}$ is weakly Pareto optimal of the original MOO problem (problem (1)) if and only if there exists a preference vector ${\bm{\alpha}}$ such that ${\bm{\theta}}^{*}$ is an optimal solution of problem (4).
If ${\bm{\theta}}^{*}$ is unique for a given ${\bm{\alpha}}$, it is considered Pareto optimal. However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator. This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18]. Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, the problem (4) remains non-differentiable.
3 Finding a single Pareto optimal solution
In some scenarios, it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set. Here, all objectives are treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\cdots,\frac{1}{m}]$), and the goal is to produce a single model that effectively balances all objectives. Most approaches in multi-task learning (MTL) [3, 69, 57] belong to this category.
The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL. However, EW may cause some tasks to have unsatisfactory performance [70, 12]. Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
$\min_{{\bm{\theta}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}}),$ (5)
where $\lambda_{i}$ is the weight for the $i$-th objective $f_{i}({\bm{\theta}})$ and the uniform preference vector ${\bm{\alpha}}$ is omitted for simplicity. These methods can be categorized as loss balancing methods and gradient balancing approaches. Some representative loss/gradient balancing methods are illustrated in Figure 3 (adapted from [19, 11]).
3.1 Loss Balancing Methods
This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss [9], homeostatic uncertainty of loss [10], loss scale [11, 12], and validation loss [13, 14, 16, 15].
Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations for the corresponding objective. Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
$\lambda_{i}^{(k)}=\frac{m\exp(\omega^{(k-1)}_{i}/\gamma)}{\sum_{j=1}^{m}\exp(\omega_{j}^{(k-1)}/\gamma)},\quad\omega^{(k-1)}_{j}=\frac{f^{(k-1)}_{j}}{f^{(k-2)}_{j}},$ (6)
where $\gamma$ is the temperature and $f^{(k)}_{i}$ is the loss value of $i$-th objective at $k$-th iteration.
Kendall et al. [10] propose Uncertainty Weighting (UW) that considers the homoscedastic uncertainty of each objective and optimizes objective weights together with the model parameter as follows:
$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$ (7)
where ${\bm{s}}$ is learnable and $\log s_{i}$ is a regularization term.
Impartial Multi-Task Learning (IMTL) [11] balances losses across tasks, abbreviated as IMTL-L(oss). Specifically, IMTL-L encourages all objectives to have a similar loss scale by transforming each objective $f_{i}({\bm{\theta}})$ as $e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$. Similar to UW [10], IMTL-L simultaneously learns the transformation parameter ${\bm{s}}$ and the model parameter ${\bm{\theta}}$ as follows:
$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}\right).$ (8)
Similar to IMTL-L [11], Lin et al. [12] also aim to make all objective losses have a similar scale. They achieve it by performing a logarithm transformation on each objective (i.e., $\log f_{i}({\bm{\theta}})$). Moreover, they theoretically show that the transformation in IMTL-L is equivalent to the logarithm transformation when $s_{i}$ is the exact minimizer of $\min_{s_{i}}e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$ in each iteration. Hence, the logarithm transformation can recover the transformation in IMTL-L.
Ye et al. [13, 14] propose Multi-Objective Meta Learning (MOML) to adaptively tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ according to the validation performance by reformulating problem (5) as a multi-objective bi-level optimization problem:
$\min_{{\bm{\lambda}}}~{}\big{(}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\cdots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{)}^{\top}$ (9a)
$\ ~{}\mathrm{s.t.}\ ~{}{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$ (9b)
where $\mathcal{D}^{\text{tr}}_{i}$ and $\mathcal{D}^{\text{val}}_{i}$ are the training and validation datasets for the $i$-th objective, respectively. In each training iteration, when given objective weights $\{\lambda_{i}\}_{i=1}^{m}$, MOML first learns the model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the training dataset by solving the lower-level (LL) subproblem (9b) with $T$ iterations and then updates ${\bm{\lambda}}$ in the upper-level (UL) subproblem (9a) via minimizing the loss of the trained MTL model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the validation dataset of each objective. However, when solving the UL subproblem (9a), MOML needs to calculate the complex hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ which requires to compute many Hessian-vector products via the chain rule. Hence, the time and memory costs of MOML grow significantly fast with respect to the dimension of ${\bm{\theta}}$ and the number of LL iterations $T$.
Auto-$\lambda$ [15] has a similar formulation with problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem (i.e., $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$). Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach. Thus, it is more efficient than MOML [13, 14].
Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14]. It first reformulates the problem (9) as a constrained MOO problem via the value-function approach [71] and then proposes a multi-gradient aggregation method to solve it. Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$. Thus, it is significantly more efficient than MOML.
Random Weighting (RW) [17] randomly samples objective weights from a distribution (such as the standard normal distribution) and normalizes them into the simplex in each iteration. Compared to EW, RW has a higher probability of escaping local minima than EW due to the randomness from objective weights, resulting in better generalization performance.
Different from the above methods which are based on linear scalarization, Lin et al. [18] propose the Smooth Tchebycheff scalarization (STCH) to make the Tchebycheff scalarization a smooth function, which is defined as:
$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu):=\mu\log\left(\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right)\right),$ (10)
where $\mu>0$ is a smoothing parameter, and ${\bm{z}}^{*}$ is a reference point which is the same as in the Tchebycheff scalarization (problem (4)). STCH improves upon the original Tchebycheff scalarization by making $f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu)$ smooth when all objective functions $f_{i}({\bm{\theta}})$ are smooth, resulting in faster convergence. It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) the Pareto optimality property by appropriately choosing $\mu$.
3.2 Gradient Balancing Methods
Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ denoting the gradient of the $i$-th objective at the $k$-th iteration. This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration. Then, the model parameter is updated via ${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta{\bm{d}}^{(k)}$, where $\eta$ is a step size. To simplify notations, we omit the superscript in this section. Compared with loss balancing methods in Section 3.1, gradient balancing approaches usually achieve better performance.
Figure 3: Illustration of the update direction ${\bm{d}}$ of several representative methods: EW, DWA [9] from loss balancing methods, MGDA [6] from gradient weighting methods, and PCGrad [25] from gradient manipulation methods. They are illustrated in a two-objective learning problem (two specific gradients are labeled as ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$).
3.2.1 Gradient Weighting Methods
In most gradient weighting methods, ${\bm{d}}$ is computed by:
${\bm{d}}={\bm{G}}{\bm{\lambda}},$ (11)
where ${\bm{G}}=[{\bm{g}}_{1},\cdots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=(\lambda_{1},\cdots,\lambda_{m})^{\top}$ is the gradient weights.
Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [64], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses, formulated as follows:
$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))$
$\approx$ $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is computed as follows:
${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$ (13)
where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun [6] use the Frank-Wolfe algorithm [72] to solve problem (13).
Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a diagram $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant. Specifically, ${\bm{d}}$ is computed by solving the following problem:
$\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~{}~{}\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$ (14)
Problem (14) is equivalent to first compute ${\bm{\lambda}}$ by solving the following problem:
${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$ (15)
where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate ${\bm{d}}$ as follows:
${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}.$ (16)
Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
Similar to CAGrad [19], Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) [20] also constraints the update direction ${\bm{d}}$ stays not far away from a target direction (such as ${\bm{g}}_{0}$), while SDMGrad achieves it by leveraging an inner-product regularization ${\bm{g}}_{0}^{\top}{\bm{d}}$.
IMTL-G(rad) [11] aims to find ${\bm{d}}$ that has equal projections on all objective gradients. Thus, we have:
${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\qquad 2\leq i\leq m,$ (17)
where ${\bm{u}}_{i}=\nicefrac{{{\bm{g}}_{i}}}{{\left\|{\bm{g}}_{i}\right\|}}$. If constraining $\sum_{i=1}^{m}\lambda_{i}=1$, problem (17) has a closed-form solution of ${\bm{\lambda}}$:
${\bm{\lambda}}_{(2,\cdots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (18)
where ${\bm{\lambda}}_{(2,\cdots,m)}=(\lambda_{2},\cdots,\lambda_{m})^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\cdots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\cdots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
Nash-MTL [21] formulates problem (11) as a bargaining game [73, 74], where each objective is cast as a player and the utility function for each player is defined as ${({\bm{g}}_{i})}^{\top}{\bm{d}}$. Hence, ${\bm{d}}$ is computed by solving the following problem:
$\max_{{\bm{d}}}\sum_{i=1}^{m}\log\left({{\bm{g}}_{i}}^{\top}{\bm{d}}\right).$ (19)
The solution of problem (19) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is obtained by solving the following problem:
${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1},$ (20)
which can be approximately solved using a sequence of convex optimization problems.
FairGrad [22] then extends the more general formulation of Nash-MTL [21]. It computes ${\bm{\lambda}}$ by solving the following problem:
${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma},$ (21)
where $\gamma\geq 0$ is a constant. Different from Nash-MTL [21], FairGrad considers problem (21) as a simple constrained nonlinear least square problem.
Aligned-MTL [23] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$. Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed. Then, the eigenvalues $\{\sigma_{1},\cdots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained. Finally, ${\bm{\lambda}}$ is computed by:
${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$ (22)
where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\cdots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
Gradient Normalization (GradNorm) [24] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have a similar magnitude. It updates ${\bm{\lambda}}$ by solving the following problem via one-step gradient descent:
$\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$ (23)
where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is cast as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ denotes the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
Dual-Balancing Multi-Task Learning (DB-MTL) [12] also normalizes all objective gradients to the same magnitude. ${\bm{\lambda}}$ is computed as $\lambda_{i}=\nicefrac{{\gamma}}{{\left\|{\bm{g}}_{i}\right\|}}$, where $\gamma=\max_{i\in[m]}\left\|{\bm{g}}_{i}\right\|$ is a scaling factor controlling the update magnitude. Thus, compared with GradNorm [24], this method guarantees all objective gradients have the same norm in each iteration. Moreover, they observe that the choice of the update magnitude $\gamma$ significantly affects performance.
3.2.2 Gradient Manipulation Methods
Gradient manipulation methods usually correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ to address the conflicting gradient problem and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
Projecting Conflicting Gradients (PCGrad) [25] defines that two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ are conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$. Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients. Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$. Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as follows:
$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$ (24)
This process reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients. Finally, the update direction is calculated as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
While PCGrad [25] corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) [26] extends it to a more general and adaptive formulation. In each iteration, if $\phi_{ij}<\bar{\phi}_{ij}$, GradVac updates the corrected gradient $\hat{{\bm{g}}}_{i}$ as:
$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\left\|\hat{{\bm{g}}}_{i}\right\|\left(\bar{\phi}_{ij}\sqrt{1-\phi_{ij}^{2}}-\phi_{ij}\sqrt{1-\bar{\phi}_{ij}^{2}}\right)}{\left\|{\bm{g}}_{j}\right\|\sqrt{1-\bar{\phi}_{ij}^{2}}}{\bm{g}}_{j},$ (25)
where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$. $\bar{\phi}_{ij}$ is initialized with 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter. Note that GradVac is simplified to PCGrad if $\bar{\phi}_{ij}=0$.
Gradient Sign Dropout (GradDrop) [27] considers that conflicts arise from differences in the sign of gradient values. Thus, a probabilistic masking procedure is proposed to preserve gradients with consistent signs during each update.
3.2.3 Practical Speedup Strategy
Gradient balancing methods suffer from high computational and storage costs. Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$, which can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters. Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], SDMGrad [20], Nash-MTL [73], and FairGrad [22]) need to solve an optimization problem to obtain objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs. Hence, several strategies are proposed to alleviate this problem.
Sener and Koltun [6] use feature-level gradients (i.e., the gradient w.r.t the representations from the last shared layer) to replace parameter-level gradients (i.e., the gradient w.r.t shared parameters ${\bm{\theta}}$). Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs. This strategy is adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [23].
Liu et al. [19] randomly select a subset of objectives to calculate the update direction in each iteration, which is adopted by SDMGrad [20]. Navon et al. [21] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iteration instead of in each iteration. Although this strategy speedups the training process, they observe that it may cause a noticeable drop in performance. Liu et al. [75] consider optimizing the logarithm objective of MGDA [6] and propose a speedup strategy for it. Specifically, when solving problem (13) via one-step gradient descent, ${\bm{\lambda}}$ is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$. Note that,
$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$ (26)
$\approx\frac{1}{\eta}\left(f^{(k)}_{1}-f^{(k+1)}_{1},\cdots,f^{(k)}_{m}-f^{(k+1)}_{m}\right)^{\top},$ (27)
where $f^{(k)}_{i}$ is the loss value of the $i$-th objective in the $k$-th iteration. Hence, ${\bm{\lambda}}$ can be approximately updated using the change in losses without computing all objective gradients. Although this strategy significantly reduces computational and memory costs, it may cause performance degradation. Moreover, it is only applicable to MGDA-based methods.
3.3 Discussions
Loss balancing and gradient balancing methods can effectively mitigate conflicts among different objectives and achieve a Pareto optimal solution, which is important for MOO scenarios such as multi-task learning. However, they are limited to finding a single Pareto optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning. In Sections 4 and 5, we introduce some methods to achieve more diverse and controllable solutions.
The computational cost of loss balancing methods (except MOML [13, 14], Auto-$\lambda$ [15], and FORUM [16]) is almost the same as that of EW since they only need one backpropagation process in each training iteration. However, gradient balancing methods are much more computationally expensive due to the need of $m$ backpropagation processes to obtain each objective gradient, where $m$ is the number of objectives. Although some speedup strategies are proposed to alleviate this problem, they may cause a drop in performance, as detailed in Section 3.2.3. Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, as detailed in Section 6.1.
4 Finding a finite set of solutions
In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs. This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the PS. We first discuss preference vector-based methods in Section 4.1, followed by preference vector-free approaches in Section 4.2.
4.1 Preference Vector-Based Methods
Generally, preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\cdots,{\bm{\alpha}}^{(n)}\}$. These preference vectors partition the problem into $n$ subproblems, where each subproblem involves finding the solution that corresponds to a specific preference vector in the set. By solving $n$ subproblems, these methods obtain a set of $n$ solutions.
Pareto Multi-Task Learning (PMTL) [28] incorporates preference vectors as constraints on the objectives. For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to other preference vectors, formulated as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$
$~{}~{}\mathrm{s.t.}~{}~{}r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,\quad j\in[n].$ (28)
Let ${\bm{R}}$ represent the gradients of active constraints, i.e, ${\bm{R}}=\big{\{}\nabla r^{(j)}({\bm{\theta}})\big{\}}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ as the active constraint indices. Using a method akin to MGDA [6], PMTL derives a descent direction for problem (28) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$. The descent direction is:
${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$ (29)
where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\cdots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
${\bm{\lambda}},{\bm{\beta}}=\operatorname*{arg\,min}_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},$
[…]
A limitation of PMTL is its weak constraint that only requires the objective vector to lie within a sector, which lacks precise control over the Pareto position.
Exact Pareto Optimal search (EPO) [29, 30] addresses this by ensuring Pareto optimality and aligning the objective vector exactly with a given preference vector ${\bm{\alpha}}$. EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with ${\bm{\alpha}}$, achieving precise alignment and Pareto optimality. Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ denote the $i$-th column of ${\bm{C}}$. Denote ${\bm{a}}=[a_{1},\cdots,a_{m}]$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$. At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and ${\mathbb{J}}^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$. Based on these sets, it determines ${\bm{\lambda}}$ for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
${\bm{\lambda}}=\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
$\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right.$ (33)
Such updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector. However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (33)).
To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [31] considers the dual form of the Tchebycheff scalarization (problem (4)) of a linear programming problem. Optimizing the Tchebycheff function addresses the previously mentioned drawbacks of EPO. However, its convergence rate is relatively slow. Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) the smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the preference-based MGDA (PMGDA) [32] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient, rather than the gradient itself.
FERERO [33] captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint. It achieves fast convergence rates of $\mathcal{O}(\epsilon^{-1})$ for deterministic gradients and $\mathcal{O}(\epsilon^{-2})$ for stochastic gradients, where $\epsilon$ is the error tolerance.
The methods discussed above rely on a fixed set of preference vectors. However, in real-world applications, where the shape of the Pareto front is unknown, prefixed preference vectors do not always result in well-distributed solutions. To overcome this limitation, GMOOAR [34] formulates the problem as a bi-level optimization problem. In the upper level, preference vectors are optimized to maximize either the hypervolume or the uniformity of the solutions. In the lower level, solutions are optimized based on these preference vectors. This approach enables the algorithm to dynamically adjust the preference vectors based on the given optimization problem, ensuring the desired distribution of solutions is achieved. UMOD [35] introduces an approach that aims to maximize the minimum distance between solutions:
$\max\left(\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))\right),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (34)
where $\rho(\cdot,\cdot)$ denotes the Euclidean distance. It achieves desirable solution distributions with a theoretical guarantee: (1) for bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) as the number of solutions increases, objective vectors asymptotically distribute on the PF.
4.2 Preference Vector-Free Methods
Another line of research does not rely on a preference vector set. Instead, they directly optimize for a diverse set of Pareto optimal solutions.
As HV evaluates solution sets based on both convergence and diversity, it is commonly used for optimizing solution sets [76, 77, 78]. Let $\{{\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors. Using the chain rule, the hypervolume gradient [36, 37, 38] of the $i$-th solution ${\bm{\theta}}^{(i)}$ can be expressed as:
${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (35)
Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}-\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate. Term ${\bm{B}}$ can be easily estimated via backpropagation. The main challenge lies in estimating the first term $a_{i}$. Following Emmerich et al. [38], index sets are categorized based on differentiability: ${\mathbb{Z}}$ includes subvectors with zero partial gradients, ${\mathbb{U}}$ contains subvectors with undefined or indeterminate gradients, and ${\mathbb{P}}$ has subvectors with positive gradients. This classification leads to many zero-gradient terms, improving efficiency. Then, non-zero terms are computed using a fast dimension-sweeping algorithm [79, 80]. This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions. To further enhance convergence, the Newton-hypervolume method [81] incorporates second-order information for faster updates. Deist et al.  [36] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters as decision variables.
MOO-SVGD [39] proposes a different approach inspired by Stein Variational Gradient Descent (SVGD) [82] to maintain diversity while enhancing convergence. The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
${\bm{d}}^{(i)}=\underbrace{-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term1}}+$
$\qquad\qquad\qquad\qquad\mu\underbrace{\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term2}},$ (36)
where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\lambda_{i}^{*}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)). The function $\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$ represents a kernel function that operates on the objective vectors ${\bm{f}}({\bm{\theta}}^{(i)})$ and ${\bm{f}}({\bm{\theta}}^{(j)})$. A common choice for this kernel is the radial basis function (RBF) kernel. In this equation, term1 guides a solution to the PS, while term2 pushes the solutions away from each other to enhance solution diversity.
The methods discussed above focus on scenarios where the number of solutions $n$ typically exceeds the number of objectives $m$. Recent works [40, 42, 83, 41] explore cases with more objectives than solutions (i.e., $m>n$). Few for Many (F4M) [40, 83] identifies solutions using a Tchebycheff method to minimize the maximal value of each objective, formulated as:
$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\max_{i\in[m]}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (37)
Sum of Minimum (SoM) [41] minimizes the sum of the minimum values for each objective across solutions, formulated as:
$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (38)
The difference between F4M and SoM is that the former uses a ”max” aggregation in the outer loop, while the latter uses a ”sum” aggregation.
MosT [42] addresses this problem by finding a uniform Pareto set through a bilevel optimization framework. The outer loop uses a weighted MGDA to find Pareto optimal solutions, while the inner loop optimizes weights via an optimal transport problem for uniform distribution.
4.3 Discussions
Preference vector-based methods decompose the problem into multiple independent problems based on preference vectors. However, selecting suitable preference vectors can be challenging in real-world scenarios. On the other hand, preference vector-free methods avoid this challenge by directly optimize a set of solutions. While they produce a set of Pareto optimal solutions with interdependent information between solutions, they typically demand more memory and are more computationally intensive.
5 Finding an infinite set of solutions
While a finite set of solutions can only provide a discrete approximation of the Pareto front, many applications require the ability to obtain solutions corresponding to any user preference, effectively representing the entire Pareto set. Directly learning an infinite number of solutions individually is impractical.
Ma et al. [84] initially investigated deriving an infinite set of solutions by using a first-order expansion around discrete Pareto optimal solutions to approximate the PS. However, this method has several drawbacks: (1) the approximation error grows when the solutions are widely spaced, (2) first-order approximations perform poorly in high-dimensional objective spaces, and (3) the approach requires solving a linear programming problem for updates, which can be computationally challenging.
To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions, enabling the capture of the entire Pareto set. These methods rely on designing efficient network architectures and implementing effective training strategies.
In Sections 5.1.1, 5.1.2, and 5.1.3, we introduce various network structures designed to capture the entire Pareto set through preference-based mappings. Specifically, current algorithms can be categorized into three main types based on their network structures: Hypernetwork-Based, Preference-Conditioned Network-Based, and Model Combination-Based. An illustration of these methods is shown in Figure 4. Finally, in Section 5.2, we discuss training strategies for these network structures.
5.1 Network Structure
(a) Hypernetwork-based methods.
(b) Preference-conditioned network-based methods.
(c) Model combination-based methods.
Figure 4: Illustration of different structures to learn an infinite number of solutions.
5.1.1 Hypernetwork
Hypernetwork is a neural network that generates the parameters for another target network [85]. This idea has been leveraged by the Pareto Hypernetwork (PHN) [43] and Controllable Pareto Multi-Task Learning (CPMTL) [44] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters. The hypernetwork usually consists of several MLP layers, and this framework has been adopted in many subsequent approaches such as PHN-HVI [45] and SUHNPF [46]. Recently, Tuan et al. [47] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
A primary limitation of hypernetwork-based algorithms is their size. Since the output dimension matches the number of parameters in the target networks, hypernetworks are often much larger than the base networks, limiting their applicability to large models. Some methods address this limitation by employing hypernetworks with chunking [43, 44]. Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability. This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
5.1.2 Preference-Conditioned Network
In preference-conditioned network-based methods, instead of using a hypernetwork to generate weights, the original model is directly modified to incorporate preferences. COSMOS [48] suggests adding preferences as an additional input to the model by combining user preferences ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
However, this input-based conditioning has limited ability to produce diverse solutions. Some studies [49, 34, 50] employ the Feature-wise Linear Modulation (FiLM) layer [86] to condition the network. The FiLM layer works by applying an affine transformation to the feature maps. Specifically, given a feature map ${\bm{u}}$ with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference ${\bm{\alpha}}$. Then, the FiLM layer modifies the features as follows:
${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ (39)
where ${\bm{u}}_{c}$ represents the $c$-th channel of ${\bm{u}}$. This method achieves better conditioning ability compared to merely adding the preference to the input layer. Raychaudhuri et al. [50] also propose leveraging another network to predict the network architecture. This approach dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
5.1.3 Model Combination
Model combination-based methods construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity. PaMaL [51] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$ (40)
Although this method incorporates more parameters than preference-conditioned network-based methods, PaMaL enhances diversity, providing users with a broader range of choices.
Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives. To mitigate this issue, LORPMAN [52] proposes learning a full-rank based model, ${\bm{\theta}}_{0}$, and several low-rank models, ${\bm{B}}_{i}{\bm{A}}_{i}$, where ${\bm{B}}_{i}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times k}$. Given a user preference ${\bm{\alpha}}$, the composite model is given by:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$ (41)
where $s$ is a scaling factor that adjusts the impact of the low-rank components. This approach efficiently scales Pareto set learning to manage up to 40 objectives, achieving superior performance. A similar approach is also mentioned in [56]. To further enhance parameter efficiency, Chen and Kwok [53] introduce a preference-aware tensor multiplication:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$ (42)
where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ represents a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times d}$, ${\bm{B}}\in\mathbb{R}^{r\times k}$ are learnable matrices, and ×$\times_{u}$ represents the mode-$u$ tensor multiplication [87].
Furthermore, Zhong et al.[54] address scalability challenges for large-scale models, such as language models, using SVD decomposition:
${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$ (43)
where ${\bm{U}}\in\mathbb{R}^{d\times(r+m)}$, ${\bm{V}}\in\mathbb{R}^{(r+m)\times k}$, and ${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\cdots,\sigma_{r},s\alpha_{1},\cdots,s\alpha_{m})$.
Instead of training new models, Tang et al. [55] propose to utilize a mixture of experts [88] method to combine pre-existing models in order to achieve the Pareto set. Given $n$ models, ${\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{n}$, each trained on different datasets, an initial step involves integrating these models into a unified base model, ${\bm{\theta}}_{0}$. For some modules in the network, differences between the individual models and the base model, denoted as ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$, are maintained. These differences are treated as experts within the framework. Then, the mixture of experts method is applied, where a gating network, conditioned on user preferences, determines the weighting of these experts. It enables the generation of diverse models by applying varying weighted combinations of experts to align with specific user preferences.
5.2 Training Strategy
This section outlines the training approach for the parameters of the network structures introduced in Section 5.1, referred to as ${\bm{\phi}}$. Specifically, ${\bm{\phi}}$ refer to those of the Hypernetwork discussed in Section 5.1.1, the parameters of the preference-conditioned network described in Section 5.1.2, or the parameters of the base models or low-rank models presented in Section 5.1.3. Current algorithms follow similar approaches, with the core idea being to sample user preferences and optimize the structure parameters to generate networks that align with those preferences. The main idea is to sample user preferences and optimize the structure parameters to produce networks that align with these preferences. The general training objective can be written as:
$\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~{}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$ (44)
where $({\bm{x}},{\bm{y}})$ are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$. In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted. Below, we summarize the algorithms adopted by existing methods.
  • Scalarization: As outlined in Section 2.2, scalarization combines multiple objectives into one scalar objective. Linear scalarization is common in most algorithms, such as PHN [43], COSMOS [48], PAMAL [51], and LORPMAN [52]. Alternatives like Tchebycheff and smooth Tchebycheff scalarization [18] can also be adopted.
  • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization. PHN [43] employs the Exact Pareto Optimal (EPO) solver [29] to find Pareto optimal solutions aligned with preferences. CPMTL [44] uses a constrained approach inspired by PMTL [28].
  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can simultaneously optimize both the diversity and convergence of a solution set. PHN-HVI [45] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$. In each iteration, it samples a set of preference vectors and generates solutions based on these vectors. Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
5.3 Discussions
Selecting the appropriate structural framework and optimization algorithms is important for effectively learning an infinite Pareto set.
Hypernetwork-based approaches offer significant flexibility by generating a diverse set of models tailored to varying user preferences. However, a primary limitation of hypernetworks is their size. Since the output dimension directly corresponds to the number of parameters in the target network, hypernetworks can become substantially larger than the base models. This increase in size can limit their applicability, particularly for large-scale networks where computational and memory resources are constrained.
In contrast, preference-conditioned network-based methods are highly parameter-efficient. They introduce only a minimal number of additional parameters relative to the base model, making them well-suited for large-scale applications where adding extra parameters is costly. Despite their efficiency, these methods may struggle to provide sufficient diversity within the parameter space. This limitation can hinder their ability to approximate complex PS accurately, as the restricted capacity may fail to capture all the nuances of Pareto optimal solutions.
Model combination-based approaches serve as a middle ground between hypernetwork and preference-conditioned network methods. By integrating multiple models, they can offer a robust approximation of the PS while maintaining scalability for large models and handling multiple objectives effectively. These methods allow for parameter adjustments through the selection of appropriate ranks in low-rank approximations.
Regarding training strategies, most algorithms use linear scalarization for its simplicity and strong performance. However, when aligning solutions with user preferences is important, preference-dependent weighting schemes discussed in Section 4.1 are more suitable. Additionally, to achieve HV optimal solutions, HV-based optimization criteria discussed in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
6 Theories
6.1 Convergence of Gradient-Balancing Methods in Section 3.2
This section reviews the paper that provides a theoretical analysis of the convergence of gradient-balancing methods in Section 3.2. The convergence property is first examined under the deterministic gradient setting (i.e., full-batch gradient descent). Subsequently, the convergence property is analyzed under the stochastic gradient setting (i.e., stochastic gradient descent). Note that in non-convex MOO, convergence refers to reaching Pareto stationary, as defined in Definition 6.
6.1.1 Deterministic Gradient
When dealing with deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto stationary point with a convergence rate of the order -1/2 [89], where $K$ is the number of iterations. This rate is comparable to that of single-objective optimization [90]. The fundamental idea is to assess the reduction in each individual objective function. When the current solution is not Pareto stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization. Therefore, the method used in single-objective optimization can be applied to prove the convergence. Furthermore, other algorithms such as CAGrad [19], Nash-MTL [21], and PCGrad [25] also offer convergence analyses in similar approaches.
TABLE II: Convergence of existing stochastic gradient MOO algorithms. LS: $L$-smooth, BG: bounded gradient, BF: bounded function value, GS: generalized $L$-smooth, CA: conflict-avoidant.
Method Batch Size Assumption Sample Complexity Bounded CA Distance
SMG [91] ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
CR-MOGM [92] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoCo [93] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
PSMGD [94] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoDo [95] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SDMGrad [20] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SGSMGrad [96] ${\mathcal{O}}(1)$ GS ${\mathcal{O}}(\epsilon^{-2})$
6.1.2 Stochastic Gradient
Liu and Vicente [91] are the first to analyze the convergence of gradient-based MOO algorithms under stochastic gradients. They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto optimal solutions under the assumption of convex objective functions. However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence. This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
To overcome the limitations associated with increasing batch sizes, Zhou et al. [92] propose the Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM) method. It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient ${\bm{\lambda}}$. In CR-MOGM, the weight coefficient ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$ (45)
where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration. This smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction. Xu et al. [94] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM. They also match the convergence rate in stochastic single-objective optimization [97].
Another approach to address gradient bias is MoCo [93], which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right),$ (46)
where $\prod_{L_{i}}$ denotes the projection to set $\{{\bm{g}}\in\mathbb{R}^{d}~{}~{}|~{}~{}\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$. However, the analysis requires the number of iterations $K$ to be sufficiently large (on the order of ${\mathcal{O}}(m^{10})$). This high dependency on $m$ makes MoCo less practical for problems involving many objectives. Additionally, the convergence analysis of CR-MOGM and MoCo relies on the assumption that the objective functions have bounded values.
To address these limitations, Chen et al. [95] propose the Multi-objective gradient with double sampling algorithm (MoDo). MoDo mitigates bias in stochastic gradient-based MOO methods without requiring increasing batch sizes or assuming bounded function values. It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients. Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$ (47)
where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ denotes projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are the stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$. By reducing the bias in estimating the common descent direction, MoDo achieves convergence guarantees without bounded function value assumption. Additionally, it guarantees a bounded conflict-avoidant (CA) distance, which is the distance between the estimated update direction ${\bm{d}}$ and the conflict-avoidant direction defined in problem (12).
Similarly, Xiao et al. [20] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction, such as the average gradient of all objectives. In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated by:
${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta\left[{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})^{\top}\left({\bm{G}}^{(k)}({\bm{z}}_{2}^{(k)}){\bm{\lambda}}^{(k)}\right.\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\left.\left.+\gamma{\bm{g}}_{0}({\bm{z}}^{(k)}_{2})\right)\right]\right),$ (48)
where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$ (e.g., the weighted sum of gradients). By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction. Zhang et al. [96] further consider the convergence under generalized $L$-smoothness [98] and without bounded gradient assumption. A summary of convergence of existing stochastic gradient MOO algorithms is provided in Table II.
6.2 Generalization
The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties. Cortes et al. [99] analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations. Sukenik et al. [100] consider the generalization for a broader class of scalarization methods. They show that the generalization bounds for individual objectives extend to MOO with scalarization. Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
Another line of research investigates the Tchebycheff scalarization method, which focuses on the sample complexity required to achieve generalization error within $\epsilon$ of the optimal objective value. Formally, given a set of $m$ distributions $\{{\mathcal{D}}_{i}\}_{i=1}^{m}$ and a hypothesis class ${\mathcal{H}}$, the goal is to find a (possibly randomized) hypothesis $h$ such that
$\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$ (49)
where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss on ${\mathcal{D}}_{i}$ with hypothesis $h$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance. It is first formulated by Awasthi et al. [101] as an open problem in 2023 and has since been addressed by several works [102, 103]. Haghtalab et al. are the first to show the lower bound for sample complexity of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$. By using a boosting framework, Peng [102] gives an algorithm that achieves an upper bound for sample complexity of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal. In a concurrent work, Zhang et al. [103] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$, addressing the open problem raised by Awasthi et al. [101].
In the online setting, Liu et al. [104] provide an adaptive online mirror descent algorithm that achieves regret on the order of $O\left(\frac{mv}{\sqrt{K}}\right)$, where $K$ is the number of iterations. Using a plain online-to-batch conversion scheme, this algorithm leads to an $O\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$. It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
Moreover, MoDo [95] introduces a different perspective by leveraging algorithm stability to derive the first set of algorithm-dependent generalization error bounds of gradient balancing algorithms.
7 Applications
7.1 Reinforcement Learning
Reinforcement Learning (RL) [105] involves training agents to make sequential decisions by maximizing cumulative rewards. Multi-Objective Reinforcement Learning (MORL) extends single-objective RL by utilizing a vector-valued reward function, i.e., ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$. MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
$\min_{{\bm{\theta}}}{\bm{f}}({\bm{\theta}}):=\mathbb{E}_{a\sim\pi_{\bm{\theta}}(s)}\sum_{t=1}^{\infty}\gamma^{t}{\bm{r}}(s_{t},a_{t}),$ (50)
where $t$ is the time step and $\gamma^{t}$ is the discount factor.
Radius Algorithm (RA) [106] employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem. However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions. To address this, various methods focus on improving the configuration of preference vectors. For instance, Optimal Linear Support (OLS) [107, 108] adds new preference vectors dynamically during each iteration. Moreover, Prediction-guided MORL (PGMORL) [109] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
Although Lu et al. [5] demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization. CAPQL [110] introduces a scalarization method that combines linear scalarization with an entropy term. PPA [111] incorporates a term to align objective values more closely with preference vectors. Several studies [112, 113, 114, 115, 116, 117] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto optimal policies. Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
Another research direction focuses on learning the entire Pareto set [118, 119, 120, 121, 122, 123]. In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and the preference vector ${\bm{\alpha}}$. Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
7.2 Bayesian Optimization
Bayesian Optimization (BO) [124] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model, typically a Gaussian Process, to approximate the objective function and guide the search. BO balances exploration (sampling uncertain regions) and exploitation (sampling promising regions) by optimizing an acquisition function like Expected Improvement (EI) [125] or Upper Confidence Bound (UCB) [126].
When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO). MOBO facilitates the exploration of trade-offs between different objectives and aims to approximate the Pareto set of optimal solutions.
In MOBO, acquisition functions must be designed to manage the balance between exploration and exploitation across multiple objectives. Based on this, MOBO methods can be classified into three main categories:
  1. 1.
    Scalarization-based Methods. These methods convert multi-objective problems into single-objective ones using scalarization functions discussed in Section 2.2 [127, 128]. Zhang and Golovin [129] propose using the hypervolume indicator for scalarization. Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [130], PAL [131, 132], and MOBO-RS [133] extend single-objective acquisitions to multi-objective optimization, leading to improved performance.
  2. 2.
    Hypervolume Improvement-Based Methods. These methods extend Expected Improvement (EI)  [125] to Expected Hypervolume Improvement (EHVI) [134], which selects candidate points to enhance the current solution set’s hypervolume. Many algorithms are proposed to improve the efficiency and allow parallel candidate generation  [135, 136, 137, 138, 139, 140, 141, 142].
  3. 3.
    Information Theoretic-Based Methods. These methods aim to maximize information gain about the PF by selecting candidate points. PESMO [143] reduces uncertainty in the posterior over the Pareto optimal input set. Extensions like MESMO [144] and Pareto-Frontier Entropy Search [145] refine this by focusing on output space and the PF. Joint Entropy Search (JES)[146, 147] measures combined information gain across input and output spaces. USeMO[148] offers an alternative uncertainty measure based on hyper-rectangle volumes.
The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set. Lin et al. [149] propose learning a continuous approximation of Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1). By adjusting preferences, decision-makers can flexibly explore any trade-off region within the approximate Pareto set. Subsequent improvements have been proposed to leverage diffusion model [150] and stein variational gradient descent [151] for efficient solution generation, as well as incorporate data augmentation for Pareto set model building [152].
7.3 Computer Vision
The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has successfully applied in autonomous driving [153, 154]. Since the encoder in computer vision usually contains large parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in partial tasks. Hence, many methods are proposed to address the issue from the perspective of MOO, which use some scalarization methods to balance multiple losses [155, 156, 157, 158, 159]. A comprehensive survey on multi-task dense prediction is provided in [4].
Expect the dense prediction task, MOO is also useful in other computer vision tasks, such as point cloud [160, 161, 162], medical image denoising [163], and pose estimation [164].
7.4 Neural Architecture Search
Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently [165]. Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy. For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, latency, and energy consumption. These works are mainly based on gradient-based NAS methods (like DARTS [166]) and formulate it as a multi-objective optimization problem. Among them, [8, 167, 168, 169, 170] employ some scalarization methods to identify a single solution, where task and efficiency objectives are combined with fixed weights. However, altering the weights for objectives needs a complete rerun of the search, which is computationally intensive. Therefore, Sukthanker et al. [171] propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
7.5 Recommendation System
Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty and diversity, serendipity, popularity, and others [172]. These factors naturally frame the problem as a multi-objective optimization optimization. Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives [173, 174, 175]. In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model [176, 177, 178, 179]. Additionally, some research efforts aim to identify finite Pareto sets [180] and infinite Pareto sets [181, 182], enabling the provision of personalized recommender models that cater to varying user preferences.
7.6 Large Language Models (LLMs)
Recently, there has been a growing trend of incorporating multi-objective optimization into the training of large language models (LLMs), including multi-task fine-tuning and multi-objective alignment.
Due to the powerful transferability of LLMs, users can fine-tune LLMs to specific downstream tasks or scenarios. This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment. MFTCoder [183] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks. CoBa [184] studies the multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects. MORLHF [185] and MODPO [186] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models. To avoid the LLM retraining, some methods separately train multiple LLMs for each preference dimension and merge their parameters [187, 188] or output logits [189] to deal with different preference requirements in the inference. However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs. To improve the efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [190, 191, 192] or model parameters [54].
7.7 Miscellaneous
In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [193, 194], federated learning [195, 196, 197], long-tailed learning [198, 199, 200], continual learning [201], diffusion models [202, 203, 204, 205], neural combinatorial optimization [206, 119, 207, 208, 209], and GFlowNets [210, 211]. These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
8 Resources
8.1 Benchmark Datasets
In this section, we introduce some benchmark datasets for multi-objective deep learning. They are summarized in Table III and listed in detail as follows.
  • NYUv2 dataset [212] is for indoor scene understanding. It has three tasks (i.e., semantic segmentation, depth estimation, and surface normal prediction) with $795$ training and $654$ testing samples.
  • Taskonomy dataset [213] is obtained from 3D scans of about $600$ buildings and contains $26$ tasks (such as semantic segmentation, depth estimation, surface normal prediction, keypoint detection, and edge detection) with about $4$ million samples.
  • Cityscapes dataset [214] is for urban scene understanding. It has two tasks (i.e., semantic segmentation and depth estimation) with $2,975$ training and $500$ testing samples.
  • QM9 dataset [215] is for molecular property prediction with $11$ tasks. Each task performs regression on one property. It contains $130,000$ samples.
  • Office-31 dataset [216] contains $4,110$ images from three domains (tasks): Amazon, DSLR, and Webcam. Each task has 31 classes.
  • Office-Home dataset [217] contains $15,500$ images from four domains (tasks): artistic images, clipart, product images, and real-world images. Each task has $65$ object categories collected under office and home settings.
  • CelebA dataset [218] is a large-scale face attributes dataset, containing $202,599$ face images. Each image has $40$ attribute annotation. Hence, this dataset includes $40$ tasks and each of them is a binary classification problem for one attribute.
  • CIFAR-100 dataset [219] an image classification with $100$ classes, containing $50,000$ training and $10,000$ testing image samples. These $100$ classes are divided into $20$ tasks, each of which is a $5$-class image classification problem.
  • MT10 and MT50 from Meta-World benchmark [220] are two multi-task reinforcement learning settings, containing $10$ and $50$ robot manipulation tasks, respectively.
  • XTREME benchmark [221] aims to evaluate the cross-lingual generalization abilities of multilingual representations. It contains $9$ tasks in $40$ languages, including $2$ classification tasks, $2$ structure prediction tasks, $3$ question-answering tasks, and $2$ sentence retrieval tasks.
  • PKU-SafeRLHF dataset [222] focuses on safety alignment in LLMs, containing $82,118$ question-answering pairs with the annotations of helpfulness and harmlessness.
  • UltraFeedBack dataset [223] contains $63,967$ prompts, each corresponding to four responses from different LLMs. Each response has $4$ aspects of annotations, namely instruction-following, truthfulness, honesty, and helpfulness, generated by GPT-4.
  • HelpSteer2 dataset [224] contains $21,362$ samples. Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 00 and $4$ where higher means better for each attribute.
TABLE III: Summary of benchmark datasets in multi-objective deep learning.
Dataset Description #Objectives #Samples
NYUv2 [212] indoor scene understanding $3$ $1,449$
Taskonomy [213] indoor scene understanding $26$ $\approx 4$M
Cityscapes [214] urban scene understanding $2$ $3,475$
QM9 [215] molecular property prediction $11$ $130$K
Office-31 [216] image classification $3$ $4,110$
Office-Home [217] image classification $4$ $15,500$
CelebA [218] image classification $40$ $\approx 202$K
CIFAR-100 [219] image classification $20$ $60$K
MT10/MT50 [220] reinforcement learning $10/50$
XTREME [221] multilingual learning $9$ $\approx 597$K
PKU-SafeRLHF [222] two-dimensional preference data $2$ $\approx 82$K
UltraFeedBack [223] multi-dimensional preference data $4$ $\approx 64$K
HelpSteer2 [224] multi-attributes preference data $5$ $\approx 21$K
8.2 Software
In this section, we introduce two open-source libraries for gradient-based multi-objective deep learning: LibMTL [^1] https://github.com/median-research-group/LibMTL [225] and LibMOON [^2] https://github.com/xzhang2523/libmoon [226]. These two libraries provide a unified environment for users to implement or fairly evaluate different MOO methods. Moreover, they are built on PyTorch [227] and designed in a modular fashion, allowing users to flexibly develop novel methods or apply existing MOO methods to new application scenarios. LibMTL [225] mainly focuses on finding a single Pareto optimal solution and supports over $19$ methods introduced in Section 3. Meanwhile, it also supports $6$ benchmark datasets (i.e., NYUv2, Cityscapes, QM9, Office-31, Office-Home, and XTREME datasets) introduced in Section 8.1. LibMOON [226] mainly focuses on exploring the whole Pareto set. It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5). These methods address a variety of tasks, including classification, fairness, and distribution matching.
9 Challenges and Future Directions
Despite significant progress in applying MOO in deep learning, both in learning a single model and in learning a Pareto set of models, several challenges remain.
9.1 Theoretical Understanding
While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received much less attention. For instance, Chen et al. [95], as discussed in Section 6.2, provide the first algorithm-dependent generalization analysis for MOO. Extending this work to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets. Zhang et al. [228] established a generalization bound for hypervolume-based Pareto set learning. However, the effect of various network architectures (Section 5.1) on approximating Pareto sets and convergence rate remain unclear. Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
9.2 Reducing Gradient Balancing Costs
While gradient balancing methods in Section 3.2 often show strong performance in various applications, they come with significant computational overhead. Despite the introduction of some practical speedup strategies (discussed in Section 3.2.3), existing approaches remain insufficiently efficient. A deeper understanding of the optimization differences between gradient balancing, linear scalarization, and loss balancing is crucial. This insight could facilitate the integration of gradient balancing with linear scalarization and loss balancing, reducing computational overhead significantly and enabling its application in large-scale training scenarios.
9.3 Dealing with Large Number of Objectives
Some real-world problems involve handling a large number of objectives, which poses significant challenges for current Pareto set learning algorithms in Section 5. As the number of objectives grows, the preference vector space expands exponentially, making it challenging for existing random sampling-based techniques to effectively learn the mapping between preference vectors and solutions. Developing efficient sampling strategies for high-dimensional preference spaces remains a critical research problem. Additionally, exploring methods to reduce or merge objectives based on their properties could be a promising approach to minimize the total number of objectives.
9.4 Distributed Training
Current MOO algorithms in deep learning are limited to single GPUs or machines, but scaling them to multi-GPU and distributed environments is increasingly important as models and datasets grow. This scaling introduces unique challenges compared to single-objective optimization. Gradient-balancing algorithms require efficient gradient distribution and communication across GPUs. Learning a set of Pareto optimal solutions involves synchronizing multiple models across nodes with minimal overhead. Additionally, when data for different objectives is distributed across devices and cannot be shared (e.g., due to privacy or memory constraints), collaborative computation of solutions or Pareto sets without direct data sharing becomes an important challenge.
9.5 Advancements in Large Language Models (LLMs)
As discussed in Section 7.6, current research in MOO for LLMs largely focuses on the RLHF stage. Expanding MOO techniques to other stages in the LLM lifecycle, such as pre-training and inference, is a valuable direction for future research. Addressing these challenges could lead to models that are better aligned with user needs across their entire development processes. Additionally, user preferences are currently modeled as a preference vector. However, this representation may oversimplify more complex user preferences on LLMs. Exploring advanced methods to capture and represent these intricate preferences could enhance the effectiveness and personalization of LLMs.
9.6 Application in More Deep Learning Scenarios
While MOO has already been utilized in various deep learning scenarios, as highlighted in Section 7, there remain numerous unexplored areas within this field. Multi-objective characteristics are inherently present in most deep learning problems, as methods are typically developed or models evaluated based on multiple criteria. Consequently, trade-offs often arise naturally. Leveraging MOO methods to effectively address these trade-offs presents a significant opportunity for further exploration.
10 Conclusion
In this paper, we provide the first comprehensive review of gradient-based multi-objective optimization in deep learning, covering algorithms, theoretical foundations, practical applications, benchmark datasets, and software. By unifying diverse approaches and systematically identifying key challenges, this paper serves as a foundational resource to drive future advancements in this important and evolving field.

### v2

@@
 A multi-objective optimization problem can be formulated as:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$ (1)
+$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$ (1)
 where $m\geq 2$ is the number of objectives, ${\bm{\theta}}$ is the decision variable and ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable.
+Figure 1: An overview of multi-objective deep learning.
 Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives.
-Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex.
+Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex.
 MOO is expected to achieve a solution that matches the user-defined preference.
@@
 for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al.
-[12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SDMGrad [20]; IMTL-G [11]; Nash-MTL [21]; FairGrad [22]; Aligned-MTL [23]; GradNorm [24]; DB-MTL [12], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [25]; GradVac [26]; GradDrop [27], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Preference Vector-based Methods (§4.1), finite [PMTL [28]; EPO [29, 30]; WC-MGDA [31]; PMGDA [32]; FERERO [33]; GMOOAR [34]; UMOD [35], finite_work] ] [Preference Vector-free Methods (§4.2), finite [HVGrad [36, 37, 38]; MOO-SVGD [39]; F4M [40]; SoM [41]; MosT [42], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [43]; CPMTL [44]; PHN-HVI [45]; SUHNPF [46]; Tuan et al.
-[47], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [48]; YOTO [49]; Raychaudhuri et al.
+[12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SMG [20]; CR-MOGM [21]; MoCo [22]; PSMGD [23]; MoDo [24]; SDMGrad [25]; SGSMGrad [26]; IMTL-G [11]; Nash-MTL [27]; FairGrad [28]; UPGrad [29]; Aligned-MTL [30]; GradNorm [31]; DB-MTL [12]; FAMO [32], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [33]; GradVac [34]; GradDrop [35], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL [36]; EPO [37, 38]; WC-MGDA [39]; PMGDA [40]; FERERO [41]; GMOOAR [42]; UMOD [43], finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV [44, 45, 46, 47]; MOO-SVGD [48]; F4M [49]; SoM [50]; MosT [51], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [52]; CPMTL [53]; PHN-HVI [54]; SUHNPF [55]; Hyper-Trans [56], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [57]; YOTO [58]; Raychaudhuri et al.
 [59], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [60]; LORPMAN [61]; Chen and Kwok [62]; Panacea [63]; Tang et al.
@@
 1.1 Taxonomy
-Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches based on the goal:
-  • When the goal is to obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization.
-These methods focus on appropriately adjusting the contributions of each objective to achieve a solution that is not overly biased toward any one objective.
-  • When the goal is to provide a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously.
+Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches:
+  • To obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization.
+These methods focus on adjusting the weight of each objective to achieve a solution that is not overly biased toward any one objective.
+  • To obtain a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously.
 This allows users to select from a set of solutions that best meet their specific needs.
-  • When the goal is to generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones.
+  • To generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones.
 Algorithms in this category propose efficient structures to learn this subspace.
@@
 Other surveys concentrate on applications of MOO in areas such as dense prediction tasks [4], recommend systems [71], and natural language processing [72].
-In contrast, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications.
+Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications.
 To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
+TABLE I: Summary of Notations.
+Notation Description
+${\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}$ Decision variable ${\bm{\theta}}$ in feasible set $\mathcal{K}$ with dimension $d$.
+$m$ Number of objectives.
+$n$ Number of finite Pareto solutions.
+$K$ Number of iterations.
+${\bm{f}}=[f_{1},\dots,f_{m}]^{\top}$ Objective function.
+${\bm{\lambda}}=[\lambda_{1},\dots,\lambda_{m}]^{\top}$ Objective weight vector.
+${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}$ Preference vector.
+${\bm{z}}^{*}=[z_{1}^{*},\dots,z_{m}^{*}]^{\top}$ Ideal objective value.
+${\bm{\theta}}^{(k)}$ Solution at $k$-th iteration.
+${\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}$ Solution $1$ to Solution $n$ in a size-$n$ solution set.
+${\bm{g}}^{(k)}_{i}$ The gradient vector of $i$-th objective at $k$-th iteration.
+${\bm{G}}^{(k)}=[{\bm{g}}_{1}^{(k)},\dots,{\bm{g}}_{m}^{(k)}]$ Gradient matrix at $k$-th iteration.
+${\bm{d}}^{(k)}$ Updating direction of ${\bm{\theta}}$ at $k$-th iteration.
+${\bm{\phi}}$ Parameters of Pareto set learning structures.
+$\mathrm{HV}_{\bm{r}}(\cdot)$ Hypervolume with respect to reference point ${\bm{r}}$.
+$[m]$ Index set $\{1,\dots,m\}$.
+$\Delta_{m-1}$ $(m-1)$-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
+[…]
+$\eta$ Step size for updating ${\bm{\theta}}$.
+$\epsilon$ Error tolerance.
 1.3 Organization of the Survey
-The notations used in this survey are summarized in Table I. The rest of this survey is organized as follows:
+The notations used in this survey are summarized in Table I. An overview of MOO in deep learning is provided in Figure 1. The rest of this survey is organized as follows:
   • Section 2 presents preliminaries in MOO, including key definitions and concepts;
@@
   • Section 10 summarizes this survey.
-TABLE I: Summary of Notations.
-Notation Description
-${\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}$ Decision variable ${\bm{\theta}}$ in feasible set $\mathcal{K}$ with dimension $d$.
-$m$ Number of objectives.
-$n$ Number of finite Pareto solutions.
-$K$ Number of iterations.
-${\bm{f}}=[f_{1},\cdots,f_{m}]$ Objective function.
-${\bm{\lambda}}=[\lambda_{1},\cdots,\lambda_{m}]$ Objective weight vector.
-${\bm{\alpha}}=[\alpha_{1},\cdots,\alpha_{m}]$ Preference vector.
-${\bm{z}}^{*}=[z_{1}^{*},\cdots,z_{m}^{*}]$ Ideal objective value.
-${\bm{\theta}}^{(k)}$ Solution at $k$-th iteration.
-${\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}$ Solution $1$ to Solution $n$ in a size-$n$ solution set.
-${\bm{g}}^{(k)}_{i}$ The gradient vector of $i$-th objective at $k$-th iteration.
-${\bm{G}}^{(k)}=[{\bm{g}}_{1}^{(k)},\cdots,{\bm{g}}_{m}^{(k)}]$ Gradient matrix at $k$-th iteration.
-${\bm{d}}^{(k)}$ Updating direction of ${\bm{\theta}}$ at $k$-th iteration.
-${\bm{\phi}}$ Parameters of Pareto set learning structures.
-$\mathrm{HV}_{\bm{r}}(\cdot)$ Hypervolume with respect to reference point ${\bm{r}}$.
-$[m]$ Index set $\{1,\cdots,m\}$.
-$\Delta_{m-1}$ $(m-1)$-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
-[…]
-$\eta$ Step size for updating ${\bm{\theta}}$.
-$\epsilon$ Error tolerance.
 2 Preliminary: Key Concepts in MOO
-In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to solve MOO problems in Section 2.2.
+In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2.
+An illustration of some Pareto concepts is in Appendix A.
 2.1 Pareto Concepts
@@
 A solution ${\bm{\theta}}^{*}$ is Pareto optimal if no other solution dominates it. A solution ${\bm{\theta}}^{*}$ is weakly Pareto optimal if no other solution strictly dominates it.
-Definition 3 (Pareto set (PS) [1])
-A PS is the set of all Pareto optimal solutions.
-Definition 4 (Pareto front (PF) [1])
-A PF is the set of all objective function values of the PO solutions.
-Figure 2 illustrates the aforementioned definitions.
-The blue circles represent Pareto optimal solutions, the yellow circles represent weakly Pareto optimal solutions, and the red circles are solutions dominated by at least one other solution.
-The blue curve represents the Pareto front.
-Figure 2: Illustration of Pareto optimality concepts.
-The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively.
-The blue curve represents the Pareto front.
+Definition 3 (Pareto set (PS) and Pareto front (PF) [1])
+A PS is the set of all PO solutions. A PF is the set of all objective function values of the PO solutions.
 Definition 4 (Preference vector)
@@
 Throughout this paper, a preference vector ${\bm{\alpha}}$ is constrained on a simplex $\Delta_{m-1}$, where $\Delta_{m-1}=\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
-In deep learning, the parameter ${\bm{f}}({\bm{\theta}})$ is typically optimized using gradient descent.
+In deep learning, the parameter ${\bm{\theta}}$ is typically optimized using gradient descent.
 In single-objective optimization, when the objective function is non-convex, the optimization process often reaches a stationary point.
@@
 Definition 6 (Hypervolume [74])
-Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\cdots,{\bm{q}}^{(N)}\}$ and a reference point ${\bm{r}}$, the hypervolume of ${\mathbb{S}}$ is calculated by:
+Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\dots,{\bm{q}}^{(N)}\}$ and a reference point ${\bm{r}}$, the hypervolume of ${\mathbb{S}}$ is calculated by:
 $\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$ (2)
 where $\Lambda(\cdot)$ denotes the Lebesgue measure of a set.
-HV provides a unary measure of both convergence and diversity of objective vectors. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
+HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
 2.2 Scalarization Methods
 Linear scalarization.
-The most straightforward approach to solving MOO problems (i.e., problem (1)) is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
+The most straightforward way to solve MOO problems is to reformulate them as single-objective optimization problems weighted by the given preference vector ${\bm{\alpha}}$ as follows:
 $\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$ (3)
 Linear scalarization is widely used in practice due to its simplicity.
-However, this method suffers from a significant limitations: It can only identify solutions on the convex part of the PF.
-For PFs with a concave shape, it typically finds only the endpoints.
-Tchebycheff scalarization. Another popular scalarization function is Tchebycheff scalarization [66, 67], which is defined as:
+However, this method suffers from a limitation: it can only identify solutions on the convex part of the PF.
+For PFs with a concave shape, using LS only finds the endpoints of a PF.
+Tchebycheff scalarization.
+Another way to convert an MOP to a single one is to use the Tchebycheff scalarization function [75, 76], defined as:
 $\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$ (4)
-where $z_{i}^{*}$ are reference values, which can be set as the ideal value of the $i$-th objective.
-Unlike linear scalarization, Tchebycheff scalarization can explore the entire PF by traversing all preference vectors within the preference simplex.
+where $z_{i}^{*}$ are reference values, usually set as the ideal value of the $i$-th objective.
+Unlike linear scalarization, Tchebycheff scalarization can explore the entire PF by traversing all preference vectors within the simplex.
 This ensures that both convex and concave regions of the PF are covered.
@@
 Theorem 1 ([77])
-A solution ${\bm{\theta}}^{*}$ is weakly Pareto optimal of the original MOO problem (problem (1)) if and only if there exists a preference vector ${\bm{\alpha}}$ such that ${\bm{\theta}}^{*}$ is an optimal solution of problem (4).
+A solution ${\bm{\theta}}^{*}$ of the original MOO problem (1) is weakly Pareto optimal if and only if there exists a preference vector ${\bm{\alpha}}$ such that ${\bm{\theta}}^{*}$ is an optimal solution of problem (4).
 If ${\bm{\theta}}^{*}$ is unique for a given ${\bm{\alpha}}$, it is considered Pareto optimal.
@@
 This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18].
-Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, the problem (4) remains non-differentiable.
+Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) remains non-differentiable.
 3 Finding a Single Pareto Optimal Solution
-In some scenarios, it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set.
-Here, all objectives are treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\cdots,\frac{1}{m}]$), and the goal is to produce a single model that effectively balances all objectives.
-Most approaches in multi-task learning (MTL) [3, 69, 57] belong to this category.
+In some scenarios such as multi-task learning (MTL) [3, 78, 66], it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set.
+Here, all objectives are usually treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\dots,\frac{1}{m}]^{\top}$), and the goal is to produce a single model that effectively balances all objectives.
 The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL.
@@
 This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss [9], homeostatic uncertainty of loss [10], loss scale [11, 12], and validation loss [13, 14, 16, 15].
-Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations for the corresponding objective.
+Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations.
 Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
@@
 $\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$ (7)
-where ${\bm{s}}$ is learnable and $\log s_{i}$ is a regularization term.
+where ${\bm{s}}=[s_{1},\dots,s_{m}]^{\top}$ is learnable and $\log s_{i}$’s are regularization terms.
 Impartial Multi-Task Learning (IMTL) [11] balances losses across tasks, abbreviated as IMTL-L(oss).
@@
 Ye et al.
-[13, 14] propose Multi-Objective Meta Learning (MOML) to adaptively tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ according to the validation performance by reformulating problem (5) as a multi-objective bi-level optimization problem:
-$\min_{{\bm{\lambda}}}~{}\big{(}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\cdots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{)}^{\top}$ (9a)
+[13, 14] propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights ${\bm{\lambda}}$ based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
+$\min_{{\bm{\lambda}}}~{}\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$ (9a)
 $\ ~{}\mathrm{s.t.}\ ~{}{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$ (9b)
 where $\mathcal{D}^{\text{tr}}_{i}$ and $\mathcal{D}^{\text{val}}_{i}$ are the training and validation datasets for the $i$-th objective, respectively.
-In each training iteration, when given objective weights $\{\lambda_{i}\}_{i=1}^{m}$, MOML first learns the model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the training dataset by solving the lower-level (LL) subproblem (9b) with $T$ iterations and then updates ${\bm{\lambda}}$ in the upper-level (UL) subproblem (9a) via minimizing the loss of the trained MTL model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the validation dataset of each objective.
+In each training iteration, when given objective weights $\{\lambda_{i}\}_{i=1}^{m}$, MOML first learns the model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the training dataset by solving lower-level (LL) subproblem (9b) with $T$ iterations and then updates ${\bm{\lambda}}$ in the upper-level (UL) subproblem (9a) via minimizing the loss of the trained MTL model ${\bm{\theta}}^{*}({\bm{\lambda}})$ on the validation dataset of each objective.
 However, when solving the UL subproblem (9a), MOML needs to calculate the complex hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ which requires to compute many Hessian-vector products via the chain rule.
 Hence, the time and memory costs of MOML grow significantly fast with respect to the dimension of ${\bm{\theta}}$ and the number of LL iterations $T$.
-Auto-$\lambda$ [15] has a similar formulation with problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem (i.e., $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$).
+Auto-$\lambda$ [15] modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$.
 Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach.
@@
 Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14].
-It first reformulates the problem (9) as a constrained MOO problem via the value-function approach [71] and then proposes a multi-gradient aggregation method to solve it.
+FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach [80] and then proposes a multi-gradient aggregation method to solve it.
 Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$.
@@
 Different from the above methods which are based on linear scalarization, Lin et al.
-[18] propose the Smooth Tchebycheff scalarization (STCH) to make the Tchebycheff scalarization a smooth function, which is defined as:
-$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu):=\mu\log\left(\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right)\right),$ (10)
-where $\mu>0$ is a smoothing parameter, and ${\bm{z}}^{*}$ is a reference point which is the same as in the Tchebycheff scalarization (problem (4)).
-STCH improves upon the original Tchebycheff scalarization by making $f_{\bm{\alpha}}^{\mathrm{STche}}({\bm{\theta}},\mu)$ smooth when all objective functions $f_{i}({\bm{\theta}})$ are smooth, resulting in faster convergence.
-It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) the Pareto optimality property by appropriately choosing $\mu$.
+[18] propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
+$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right),$ (10)
+where $\mu>0$ is a smoothing parameter, and ${\bm{z}}^{*}$ is a reference point as in Tchebycheff scalarization (problem (4)).
+STCH improves upon the original Tchebycheff scalarization by making $f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu)$ smooth when all objective functions $f_{i}({\bm{\theta}})$’s are smooth, resulting in faster convergence.
+It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) Pareto optimality by appropriately choosing $\mu$.
 3.2 Gradient Balancing Methods
-Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ denoting the gradient of the $i$-th objective at the $k$-th iteration.
+Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ be the gradient of the $i$-th objective at the $k$-th iteration.
 This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration.
@@
 To simplify notations, we omit the superscript in this section.
-Compared with loss balancing methods in Section 3.1, gradient balancing approaches usually achieve better performance.
+Compared with loss balancing methods in Section 3.1, the gradient balancing approach usually achieves better performance.
 Figure 3: Illustration of the update direction ${\bm{d}}$ of several representative methods: EW, DWA [9] from loss balancing methods, MGDA [6] from gradient weighting methods, and PCGrad [33] from gradient manipulation methods.
@@
 ${\bm{d}}={\bm{G}}{\bm{\lambda}},$ (11)
-where ${\bm{G}}=[{\bm{g}}_{1},\cdots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=(\lambda_{1},\cdots,\lambda_{m})^{\top}$ is the gradient weights.
-Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [64], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses, formulated as follows:
-$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))$
-$\approx$ $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
-where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is computed as follows:
+where ${\bm{G}}=[{\bm{g}}_{1},\dots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=[\lambda_{1},\dots,\lambda_{m}]^{\top}$ is the gradient weights.
+Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [73], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses.
+This is formulated as:
+$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
+where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is computed as:
 ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$ (13)
 where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun [6] use the Frank-Wolfe algorithm [81] to solve problem (13).
-Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a diagram $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant.
+Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant.
 Specifically, ${\bm{d}}$ is computed by solving the following problem:
 $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~{}~{}\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$ (14)
-Problem (14) is equivalent to first compute ${\bm{\lambda}}$ by solving the following problem:
+Problem (14) is equivalent to first computing ${\bm{\lambda}}$ by solving the following problem:
 ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$ (15)
-where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate ${\bm{d}}$ as follows:
-${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}.$ (16)
+where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate the update direction ${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}$.
 Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
-Similar to CAGrad [19], Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) [20] also constraints the update direction ${\bm{d}}$ stays not far away from a target direction (such as ${\bm{g}}_{0}$), while SDMGrad achieves it by leveraging an inner-product regularization ${\bm{g}}_{0}^{\top}{\bm{d}}$.
+MGDA [6] and its variant CAGrad [19] can converge to a Pareto stationary point in the deterministic setting with full-gradient computations.
+However, in deep learning, where stochastic gradients are used, many works [20, 21, 23, 22, 24, 25, 26] explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
 IMTL-G(rad) [11] aims to find ${\bm{d}}$ that has equal projections on all objective gradients. Thus, we have:
-${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\qquad 2\leq i\leq m,$ (17)
+${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\quad 2\leq i\leq m,$ (16)
 where ${\bm{u}}_{i}=\nicefrac{{{\bm{g}}_{i}}}{{\left\|{\bm{g}}_{i}\right\|}}$. If constraining $\sum_{i=1}^{m}\lambda_{i}=1$, problem (16) has a closed-form solution of ${\bm{\lambda}}$:
-${\bm{\lambda}}_{(2,\cdots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (18)
-where ${\bm{\lambda}}_{(2,\cdots,m)}=(\lambda_{2},\cdots,\lambda_{m})^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\cdots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\cdots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
-Nash-MTL [21] formulates problem (11) as a bargaining game [73, 74], where each objective is cast as a player and the utility function for each player is defined as ${({\bm{g}}_{i})}^{\top}{\bm{d}}$.
+${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (17)
+where ${\bm{\lambda}}_{(2,\dots,m)}=[\lambda_{2},\dots,\lambda_{m}]^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\dots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\dots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
+Nash-MTL [27] formulates problem (11) as a bargaining game [82, 83], where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$.
 Hence, ${\bm{d}}$ is computed by solving the following problem:
@@
 which can be approximately solved using a sequence of convex optimization problems.
-FairGrad [22] then extends the more general formulation of Nash-MTL [21]. It computes ${\bm{\lambda}}$ by solving the following problem:
-${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma},$ (21)
-where $\gamma\geq 0$ is a constant. Different from Nash-MTL [21], FairGrad considers problem (21) as a simple constrained nonlinear least square problem.
+FairGrad [28] extends Nash-MTL [27] by computing ${\bm{\lambda}}$ through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant.
+Unlike Nash-MTL [27], FairGrad treats the problem as a constrained nonlinear least squares problem.
+Similarly, UPGrad [29] calculates ${\bm{\lambda}}$ by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}{\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}$, a quadratic programming (QP) problem.
 Aligned-MTL [30] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$.
 Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed.
-Then, the eigenvalues $\{\sigma_{1},\cdots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained.
+Then, the eigenvalues $\{\sigma_{1},\dots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained.
 Finally, ${\bm{\lambda}}$ is computed by:
 ${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$ (20)
-where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\cdots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
-Gradient Normalization (GradNorm) [24] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have a similar magnitude.
+where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\dots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
+Gradient Normalization (GradNorm) [31] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have similar magnitudes.
 It updates ${\bm{\lambda}}$ by solving the following problem via one-step gradient descent:
 $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$ (21)
-where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is cast as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ denotes the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
+where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is treated as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ is the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
 Dual-Balancing Multi-Task Learning (DB-MTL) [12] also normalizes all objective gradients to the same magnitude.
@@
 3.2.2 Gradient Manipulation Methods
-Gradient manipulation methods usually correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ to address the conflicting gradient problem and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
-Projecting Conflicting Gradients (PCGrad) [25] defines that two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ are conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$.
+To address the conflicting gradient problem, gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
+Projecting Conflicting Gradients (PCGrad) [33] considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$.
 Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients.
 Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$.
-Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as follows:
+Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as:
 $\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$ (22)
-This process reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients.
-Finally, the update direction is calculated as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
+This reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients.
 While PCGrad [33] corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) [34] extends it to a more general and adaptive formulation.
@@
 where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$.
-$\bar{\phi}_{ij}$ is initialized with 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter.
-Note that GradVac is simplified to PCGrad if $\bar{\phi}_{ij}=0$.
+$\bar{\phi}_{ij}$ is initialized to 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter.
+Note that GradVac simplifies to PCGrad when $\bar{\phi}_{ij}=0$.
 Gradient Sign Dropout (GradDrop) [35] considers that conflicts arise from differences in the sign of gradient values.
@@
 Gradient balancing methods suffer from high computational and storage costs.
-Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$, which can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters.
-Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], SDMGrad [20], Nash-MTL [73], and FairGrad [22]) need to solve an optimization problem to obtain objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs.
+Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t.
+the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$.
+This can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters.
+Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], Nash-MTL [27], and FairGrad [28]) need to solve an optimization problem to obtain the objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs.
 Hence, several strategies are proposed to alleviate this problem.
-Sener and Koltun [6] use feature-level gradients (i.e., the gradient w.r.t the representations from the last shared layer) to replace parameter-level gradients (i.e., the gradient w.r.t shared parameters ${\bm{\theta}}$).
+Sener and Koltun [6] use feature-level gradients (i.e., gradients w.r.t.
+the representations from the last shared layer) to replace the parameter-level gradients (i.e., gradients w.r.t.
+the shared parameters ${\bm{\theta}}$).
 Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs.
-This strategy is adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [23].
+This strategy is also adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [30].
 Liu et al.
-[19] randomly select a subset of objectives to calculate the update direction in each iteration, which is adopted by SDMGrad [20].
+[19] randomly select a subset of objectives to calculate the update direction in each iteration.
 Navon et al.
-[21] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iteration instead of in each iteration.
-Although this strategy speedups the training process, they observe that it may cause a noticeable drop in performance.
+[27] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iterations instead of updating in each iteration.
+Although this strategy speeds up training, they observe that it may cause a noticeable drop in performance.
 Liu et al.
-[75] consider optimizing the logarithm objective of MGDA [6] and propose a speedup strategy for it.
+[32] consider optimizing the logarithm of the MGDA objective [6] and propose a speedup strategy.
 Specifically, when solving problem (13) via one-step gradient descent, ${\bm{\lambda}}$ is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$.
-Note that,
-$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$ (26)
-$\approx\frac{1}{\eta}\left(f^{(k)}_{1}-f^{(k+1)}_{1},\cdots,f^{(k)}_{m}-f^{(k+1)}_{m}\right)^{\top},$ (27)
+Note that
+$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$
+$\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right]^{\top},$ (24)
 where $f^{(k)}_{i}$ is the loss value of the $i$-th objective in the $k$-th iteration.
@@
 3.3 Discussions
-Loss balancing and gradient balancing methods can effectively mitigate conflicts among different objectives and achieve a Pareto optimal solution, which is important for MOO scenarios such as multi-task learning.
-However, they are limited to finding a single Pareto optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning.
-In Sections 4 and 5, we introduce some methods to achieve more diverse and controllable solutions.
+Loss balancing and gradient balancing methods effectively mitigate conflicts among different objectives and achieve a Pareto-optimal solution, which is important for MOO scenarios such as multi-task learning.
+However, they are limited to finding a single Pareto-optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning.
+In Sections 4 and 5, we introduce methods to achieve more diverse and controllable solutions.
 The computational cost of loss balancing methods (except MOML [13, 14], Auto-$\lambda$ [15], and FORUM [16]) is almost the same as that of EW since they only need one backpropagation process in each training iteration.
@@
 Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, as detailed in Section 6.1.
+Figure 4: Distributions of solutions obtained by using representative methods: PTML [36] and EPO [37, 38] are preference vector-based methods, while GradHV [47] and MOO-SVGD [48] are methods not using preference vectors.
+The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
 4 Finding a Finite Set of Solutions
 In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs.
-This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the PS.
-We first discuss preference vector-based methods in Section 4.1, followed by preference vector-free approaches in Section 4.2.
-4.1 Preference Vector-Based Methods
-Generally, preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\cdots,{\bm{\alpha}}^{(n)}\}$.
+This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the Pareto set.
+We first discuss methods based on preference vectors (Section 4.1), followed by approaches that do not require the use of preference vectors (Section 4.2).
+Solutions obtained by some representative methods on LSMOP1 [84] are shown in Figure 4, with additional details and more results provided in Appendix B.
+4.1 Methods Based on Preference Vectors
+Preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\dots,{\bm{\alpha}}^{(n)}\}$.
 These preference vectors partition the problem into $n$ subproblems, where each subproblem involves finding the solution that corresponds to a specific preference vector in the set.
 By solving $n$ subproblems, these methods obtain a set of $n$ solutions.
-Pareto Multi-Task Learning (PMTL) [28] incorporates preference vectors as constraints on the objectives.
-For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to other preference vectors, formulated as:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$
+Pareto Multi-Task Learning (PMTL) [36], inspired by the idea of decomposition-based MOO algorithms [85], incorporates preference vectors as constraints on the objectives.
+For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to the other preference vectors:
+$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$
 $~{}~{}\mathrm{s.t.}~{}~{}r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,\quad j\in[n].$ (25)
-Let ${\bm{R}}$ represent the gradients of active constraints, i.e, ${\bm{R}}=\big{\{}\nabla r^{(j)}({\bm{\theta}})\big{\}}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ as the active constraint indices.
+Let ${\bm{R}}$ be the matrix containing gradients of the active constraints, i.e, ${\bm{R}}=\big{[}\nabla r^{(j)}({\bm{\theta}})\big{]}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ containing indices of the active constraints.
 Using a method akin to MGDA [6], PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$.
@@
 ${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$ (26)
-where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\cdots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
+where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
 ${\bm{\lambda}},{\bm{\beta}}=$ $\operatorname*{arg\,min}_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},$
 […]
-A limitation of PMTL is its weak constraint that only requires the objective vector to lie within a sector, which lacks precise control over the Pareto position.
-Exact Pareto Optimal search (EPO) [29, 30] addresses this by ensuring Pareto optimality and aligning the objective vector exactly with a given preference vector ${\bm{\alpha}}$.
+A limitation of PMTL is its weak constraint that only requires the objective vector to lie within sectors, which lacks precise control over the position of Pareto solutions.
+Exact Pareto Optimal search (EPO) [37, 38] is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors.
 EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with ${\bm{\alpha}}$, achieving precise alignment and Pareto optimality.
-Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ denote the $i$-th column of ${\bm{C}}$.
-Denote ${\bm{a}}=[a_{1},\cdots,a_{m}]$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$.
+Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ be the $i$-th column of ${\bm{C}}$.
+Let ${\bm{a}}=[a_{1},\dots,a_{m}]^{\top}$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$.
 At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and ${\mathbb{J}}^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$.
 Based on these sets, it determines ${\bm{\lambda}}$ for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
-${\bm{\lambda}}=\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
+${\bm{\lambda}}=$ $\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}~{}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
 $\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right.$ (30)
-Such updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector.
+This updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector.
 However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (30)).
-To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [31] considers the dual form of the Tchebycheff scalarization (problem (4)) of a linear programming problem.
-Optimizing the Tchebycheff function addresses the previously mentioned drawbacks of EPO.
+To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [39] considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem.
+Optimizing Tchebycheff function addresses the previously mentioned drawbacks of EPO.
 However, its convergence rate is relatively slow.
-Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) the smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the preference-based MGDA (PMGDA) [32] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient, rather than the gradient itself.
+Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the Preference-based MGDA (PMGDA) [40] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient.
 FERERO [41] captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint.
@@
 The methods discussed above rely on a fixed set of preference vectors.
-However, in real-world applications, where the shape of the Pareto front is unknown, prefixed preference vectors do not always result in well-distributed solutions.
+However, in real-world applications where the shape of the Pareto front is unknown, a predefined set of preference vectors may not always result in well-distributed solutions.
 To overcome this limitation, GMOOAR [42] formulates the problem as a bi-level optimization problem.
-In the upper level, preference vectors are optimized to maximize either the hypervolume or the uniformity of the solutions.
+In the upper level, preference vectors are optimized to maximize either the hypervolume or uniformity of the solutions.
 In the lower level, solutions are optimized based on these preference vectors.
-This approach enables the algorithm to dynamically adjust the preference vectors based on the given optimization problem, ensuring the desired distribution of solutions is achieved.
+This approach enables the algorithm to dynamically adjust preference vectors based on the given optimization problem, ensuring the desired solution distribution.
 UMOD [43] introduces an approach that aims to maximize the minimum distance between solutions:
-$\max\left(\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))\right),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (34)
+$\max\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (31)
 where $\rho(\cdot,\cdot)$ denotes the Euclidean distance.
-It achieves desirable solution distributions with a theoretical guarantee: (1) for bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) as the number of solutions increases, objective vectors asymptotically distribute on the PF.
-4.2 Preference Vector-Free Methods
-Another line of research does not rely on a preference vector set. Instead, they directly optimize for a diverse set of Pareto optimal solutions.
-As HV evaluates solution sets based on both convergence and diversity, it is commonly used for optimizing solution sets [76, 77, 78].
-Let $\{{\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors.
-Using the chain rule, the hypervolume gradient [36, 37, 38] of the $i$-th solution ${\bm{\theta}}^{(i)}$ can be expressed as:
-${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (35)
-Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}-\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate.
-Term ${\bm{B}}$ can be easily estimated via backpropagation.
+UMOD achieves desirable solution distributions with theoretical guarantees: (1) For bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) As the number of solutions increases, the objective vectors asymptotically distribute on the PF.
+4.2 Methods without Using Preference Vectors
+Another line of research, directly optimize for a diverse set of Pareto-optimal solutions, does not rely on a given set of preference vectors.
+As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO [86].
+In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) [47, 44, 45, 46] have been developed.
+Let $\{{\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors.
+These algorithm first calculate the hypervolume gradient of the $i$-th solution ${\bm{\theta}}^{(i)}$ as follows:
+${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (32)
+Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}+\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate.
+The term ${\bm{B}}$ in Equation (32) can be easily estimated via backpropagation.
 The main challenge lies in estimating the first term $a_{i}$.
@@
 Deist et al.
-[36] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters as decision variables.
-MOO-SVGD [39] proposes a different approach inspired by Stein Variational Gradient Descent (SVGD) [82] to maintain diversity while enhancing convergence.
+[44] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
+Inspired by Stein Variational Gradient Descent (SVGD) [90], MOO-SVGD [48] proposes a different approach to maintain diversity while enhancing convergence.
 The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
-${\bm{d}}^{(i)}=\underbrace{-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term1}}+$
-$\qquad\qquad\qquad\qquad\mu\underbrace{\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term2}},$ (36)
-where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\lambda_{i}^{*}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)).
-The function $\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$ represents a kernel function that operates on the objective vectors ${\bm{f}}({\bm{\theta}}^{(i)})$ and ${\bm{f}}({\bm{\theta}}^{(j)})$.
-A common choice for this kernel is the radial basis function (RBF) kernel.
-In this equation, term1 guides a solution to the PS, while term2 pushes the solutions away from each other to enhance solution diversity.
-The methods discussed above focus on scenarios where the number of solutions $n$ typically exceeds the number of objectives $m$.
-Recent works [40, 42, 83, 41] explore cases with more objectives than solutions (i.e., $m>n$).
-Few for Many (F4M) [40, 83] identifies solutions using a Tchebycheff method to minimize the maximal value of each objective, formulated as:
-$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\max_{i\in[m]}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (37)
-Sum of Minimum (SoM) [41] minimizes the sum of the minimum values for each objective across solutions, formulated as:
-$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (38)
-The difference between F4M and SoM is that the former uses a ”max” aggregation in the outer loop, while the latter uses a ”sum” aggregation.
-MosT [42] addresses this problem by finding a uniform Pareto set through a bilevel optimization framework.
-The outer loop uses a weighted MGDA to find Pareto optimal solutions, while the inner loop optimizes weights via an optimal transport problem for uniform distribution.
+${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$
+$\qquad\qquad+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$ (33)
+where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)).
+$\kappa(\cdot,\cdot)$ represents a kernel function, often chosen as the radial basis function (RBF) kernel.
+The first term in (33) guides the solution towards the PS, while the second term pushes the solutions away from each other to enhance the diversity.
+The methods discussed above focus on scenarios where the number of solutions $n$ exceeds the number of objectives $m$.
+Recent works [49, 51, 91, 50] explore the case with more objectives than solutions (i.e., $m>n$).
+These approaches assign each objective to a solution ${\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}$ that achieves the smallest objective value for that particular objective.
+They then minimize either the maximum or the sum of these smallest objective values.
+Few for Many (F4M) [49, 91] uses the maximum aggregation, formulated as:
+$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$ (34)
+Sum of Minimum (SoM) [50] uses the sum aggregation, formulated as:
+$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (35)
+MosT [51] frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
 4.3 Discussions
-Preference vector-based methods decompose the problem into multiple independent problems based on preference vectors.
-However, selecting suitable preference vectors can be challenging in real-world scenarios.
-On the other hand, preference vector-free methods avoid this challenge by directly optimize a set of solutions.
-While they produce a set of Pareto optimal solutions with interdependent information between solutions, they typically demand more memory and are more computationally intensive.
+This section reviews two primary approaches for obtaining multiple Pareto optimal solutions.
+Preference vector-based methods decompose an MOO problem into multiple independent subproblems based on predefined preference vectors.
+These subproblems can be solved independently, which enhances memory efficiency and enables parallel computation.
+However, in real-world applications, determining suitable preference vectors can be challenging due to the complex and often unpredictable nature of the Pareto front.
+Furthermore, since the subproblems are solved independently, they cannot benefit from shared information between one another.
+On the other hand, methods not using preference vectors bypass the challenge of selecting preference vectors by directly optimizing a set of solutions.
+These methods generate a set of Pareto-optimal solutions that inherently share interdependent information.
+However, this approach typically requires higher memory resources and incurs greater computational costs, making it less efficient in resource-constrained scenarios.
 5 Finding an Infinite Set of Solutions
@@
 Ma et al.
-[84] initially investigated deriving an infinite set of solutions by using a first-order expansion around discrete Pareto optimal solutions to approximate the PS.
+[92] initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions.
 However, this method has several drawbacks: (1) The approximation error grows when the solutions are widely spaced, (2) First-order approximations perform poorly in high-dimensional objective spaces, and (3) The approach requires solving a linear programming problem for updates, which can be computationally challenging.
-To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions, enabling the capture of the entire Pareto set.
+To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions directly, enabling the capture of the entire PS.
 These methods rely on designing efficient network architectures and implementing effective training strategies.
-In Sections 5.1.1, 5.1.2, and 5.1.3, we introduce various network structures designed to capture the entire Pareto set through preference-based mappings.
-Specifically, current algorithms can be categorized into three main types based on their network structures: Hypernetwork-Based, Preference-Conditioned Network-Based, and Model Combination-Based.
+In Section 5.1, we introduce various network structures designed to capture the entire PS through preference-based mappings.
+Specifically, current algorithms mainly use three types of network structures: (1) Hypernetwork in Section 5.1.1, (2) Preference-Conditioned Network in Section 5.1.2, and (3) Model Combination in Section 5.1.3.
 An illustration of these methods is shown in Figure 5.
-Finally, in Section 5.2, we discuss training strategies for these network structures.
+Section 5.2 discusses training strategies for these network structures.
 5.1 Network Structure
-(a) Hypernetwork-based methods.
-(b) Preference-conditioned network-based methods.
-(c) Model combination-based methods.
+(a) Methods based on the hypernetwork.
+(b) Methods based on preference-conditioned network.
+(c) Methods based on model combination.
 Figure 5: Illustration of different structures to learn an infinite number of solutions.
@@
 This idea has been leveraged by the Pareto Hypernetwork (PHN) [52] and Controllable Pareto Multi-Task Learning (CPMTL) [53] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters.
-The hypernetwork usually consists of several MLP layers, and this framework has been adopted in many subsequent approaches such as PHN-HVI [45] and SUHNPF [46].
+The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI [54] and SUHNPF [55].
 Recently, Tuan et al.
 [56] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
-A primary limitation of hypernetwork-based algorithms is their size.
-Since the output dimension matches the number of parameters in the target networks, hypernetworks are often much larger than the base networks, limiting their applicability to large models.
+A primary limitation of hypernetwork-based algorithms is their sizes.
+Since the output dimension matches the number of parameters in the target network, hypernetworks are often much larger than the base networks, limiting their applicability to large models.
 Some methods address this limitation by employing hypernetworks with chunking [52, 53].
@@
 5.1.2 Preference-Conditioned Network
-In preference-conditioned network-based methods, instead of using a hypernetwork to generate weights, the original model is directly modified to incorporate preferences.
-COSMOS [48] suggests adding preferences as an additional input to the model by combining user preferences ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
-However, this input-based conditioning has limited ability to produce diverse solutions.
+Instead of using a hypernetwork to generate weights, the original model can be directly modified to incorporate preferences.
+COSMOS [57] suggests adding preferences as an additional input to the model by combining user preference ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
+However, such input-based conditioning has limited ability to produce diverse solutions.
 Some studies [58, 42, 59] employ the Feature-wise Linear Modulation (FiLM) layer [94] to condition the network.
@@
 Specifically, given a feature map ${\bm{u}}$ with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference ${\bm{\alpha}}$.
-Then, the FiLM layer modifies the features as follows:
-${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ (39)
-where ${\bm{u}}_{c}$ represents the $c$-th channel of ${\bm{u}}$.
+Then, the FiLM layer modifies the features as ${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ where ${\bm{u}}_{c}$ is the $c$-th channel of ${\bm{u}}$.
 This method achieves better conditioning ability compared to merely adding the preference to the input layer.
@@
 [59] also propose leveraging another network to predict the network architecture.
-This approach dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
+It dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
 5.1.3 Model Combination
-Model combination-based methods construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity.
+Methods based on model combination construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity.
 PaMaL [60] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
 ${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$ (36)
-Although this method incorporates more parameters than preference-conditioned network-based methods, PaMaL enhances diversity, providing users with a broader range of choices.
+Although this incorporates more parameters than methods based on preference-conditioned networks, PaMaL enhances diversity and provides users with a broader range of choices.
 Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives.
-To mitigate this issue, LORPMAN [52] proposes learning a full-rank based model, ${\bm{\theta}}_{0}$, and several low-rank models, ${\bm{B}}_{i}{\bm{A}}_{i}$, where ${\bm{B}}_{i}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times k}$.
-Given a user preference ${\bm{\alpha}}$, the composite model is given by:
+To mitigate this issue, LORPMAN [61] proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$.
+Given user preference ${\bm{\alpha}}$, the composite model is given by:
 ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$ (37)
@@
 ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$ (38)
-where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ represents a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times d}$, ${\bm{B}}\in\mathbb{R}^{r\times k}$ are learnable matrices, and ×$\times_{u}$ represents the mode-$u$ tensor multiplication [87].
-Furthermore, Zhong et al.[54] address scalability challenges for large-scale models, such as language models, using SVD decomposition:
+where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and ×$\times_{u}$ denotes mode-$u$ tensor multiplication [95].
+Zhong et al.[63] address scalability challenges for large-scale models, such as language models, using SVD:
 ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$ (39)
-where ${\bm{U}}\in\mathbb{R}^{d\times(r+m)}$, ${\bm{V}}\in\mathbb{R}^{(r+m)\times k}$, and ${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\cdots,\sigma_{r},s\alpha_{1},\cdots,s\alpha_{m})$.
+where ${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices.
+${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, where $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars.
 Instead of training new models, Tang et al.
-[55] propose to utilize a mixture of experts [88] method to combine pre-existing models in order to achieve the Pareto set.
-Given $n$ models, ${\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{n}$, each trained on different datasets, an initial step involves integrating these models into a unified base model, ${\bm{\theta}}_{0}$.
-For some modules in the network, differences between the individual models and the base model, denoted as ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$, are maintained.
+[64] propose to utilize a mixture of experts [96] to combine pre-existing models in order to achieve the Pareto set.
+Given $n$ models, ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}$, each trained on different datasets, they are first integrated to a unified base model ${\bm{\theta}}_{0}$.
+For some modules in the network, differences between the individual models and the base model (i.e., ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$’s) are maintained.
 These differences are treated as experts within the framework.
-Then, the mixture of experts method is applied, where a gating network, conditioned on user preferences, determines the weighting of these experts.
-It enables the generation of diverse models by applying varying weighted combinations of experts to align with specific user preferences.
+Then, a mixture of experts is applied, where a gating network conditioned on user preferences determines the weighting of these experts.
+It enables the generation of diverse models by applying different weighted combinations of experts to align with specific user preferences.
 5.2 Training Strategy
-This section outlines the training approach for the parameters of the network structures introduced in Section 5.1, referred to as ${\bm{\phi}}$.
-Specifically, ${\bm{\phi}}$ refer to those of the Hypernetwork discussed in Section 5.1.1, the parameters of the preference-conditioned network described in Section 5.1.2, or the parameters of the base models or low-rank models presented in Section 5.1.3.
-Current algorithms follow similar approaches, with the core idea being to sample user preferences and optimize the structure parameters to generate networks that align with those preferences.
-The main idea is to sample user preferences and optimize the structure parameters to produce networks that align with these preferences.
-The general training objective can be written as:
+This section outlines the training approach for the parameters (denoted ${\bm{\phi}}$) of the network structures introduced in Section 5.1.
+Specifically, ${\bm{\phi}}$ refers to the parameters of the Hypernetwork in Section 5.1.1, parameters of the preference-conditioned network in Section 5.1.2, or parameters of the base models or low-rank models in Section 5.1.3.
+Current algorithms follow a similar approach: sample user preferences and optimize the structure parameters to generate networks that align with those preferences.
+The training objective can be written as:
 $\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~{}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$ (40)
-where $({\bm{x}},{\bm{y}})$ are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$.
+where $({\bm{x}},{\bm{y}})$’s are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot,\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$.
 In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted.
-Below, we summarize the algorithms adopted by existing methods.
+Below, we summarize the algorithms adopted by the existing methods.
   • Scalarization: As outlined in Section 2.2, scalarization combines multiple objectives into one scalar objective.
@@
   • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization.
-PHN [43] employs the Exact Pareto Optimal (EPO) solver [29] to find Pareto optimal solutions aligned with preferences.
+PHN [52] employs the Exact Pareto Optimal (EPO) solver [37] to find Pareto-optimal solutions aligned with preferences.
 CPMTL [53] uses a constrained approach inspired by PMTL [36].
-  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can simultaneously optimize both the diversity and convergence of a solution set.
+  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can optimize both diversity and convergence of a solution set.
 PHN-HVI [54] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$.
@@
 Hypernetwork-based approaches offer significant flexibility by generating a diverse set of models tailored to varying user preferences.
-However, a primary limitation of hypernetworks is their size.
-Since the output dimension directly corresponds to the number of parameters in the target network, hypernetworks can become substantially larger than the base models.
-This increase in size can limit their applicability, particularly for large-scale networks where computational and memory resources are constrained.
-In contrast, preference-conditioned network-based methods are highly parameter-efficient.
-They introduce only a minimal number of additional parameters relative to the base model, making them well-suited for large-scale applications where adding extra parameters is costly.
-Despite their efficiency, these methods may struggle to provide sufficient diversity within the parameter space.
-This limitation can hinder their ability to approximate complex PS accurately, as the restricted capacity may fail to capture all the nuances of Pareto optimal solutions.
-Model combination-based approaches serve as a middle ground between hypernetwork and preference-conditioned network methods.
-By integrating multiple models, they can offer a robust approximation of the PS while maintaining scalability for large models and handling multiple objectives effectively.
-These methods allow for parameter adjustments through the selection of appropriate ranks in low-rank approximations.
+However, their main drawback is size: the output dimension matches the target network’s parameters, making hypernetworks significantly larger than base models.
+This limits their applicability for large-scale networks or resource-constrained scenarios.
+In contrast, methods based on preference-conditioned networks are highly parameter-efficient.
+They introduce a minimal number of additional parameters relative to the base model, making them well-suited for large-scale applications where adding extra parameters is costly.
+However, their limited parameter space may reduce diversity, hindering their ability to approximate complex PS accurately.
+Methods based on model combinations serve as a middle ground between hypernetwork and preference-conditioned network methods.
+By integrating multiple models, they offer a good approximation of the PS while maintaining scalability for large models.
+Furthermore, some methods allow for parameter adjustments through the selection of appropriate ranks in low-rank approximations.
 Regarding training strategies, most algorithms use linear scalarization for its simplicity and strong performance.
-However, when aligning solutions with user preferences is important, preference-dependent weighting schemes discussed in Section 4.1 are more suitable.
-Additionally, to achieve HV optimal solutions, HV-based optimization criteria discussed in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
+However, when aligning solutions with user preferences is important, preference-dependent weighting schemes in Section 4.1 are more suitable.
+Additionally, to achieve HV-optimal solutions, HV-based optimization criteria in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
 6 Theories
 6.1 Convergence of Gradient-Balancing Methods in Section 3.2
-This section reviews the paper that provides a theoretical analysis of the convergence of gradient-balancing methods in Section 3.2.
-The convergence property is first examined under the deterministic gradient setting (i.e., full-batch gradient descent).
-Subsequently, the convergence property is analyzed under the stochastic gradient setting (i.e., stochastic gradient descent).
-Note that in non-convex MOO, convergence refers to reaching Pareto stationary, as defined in Definition 6.
+This section reviews the convergence analysis for gradient-balancing methods in Section 3.2.
+This is first examined under the deterministic gradient setting (i.e., full-batch gradient).
+Subsequently, it is analyzed under the stochastic gradient setting.
+Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 5).
 6.1.1 Deterministic Gradient
-When dealing with deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto stationary point with a convergence rate of the order -1/2 [89], where $K$ is the number of iterations.
+With the use of deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ [97], where $K$ is the number of iterations.
 This rate is comparable to that of single-objective optimization [98].
 The fundamental idea is to assess the reduction in each individual objective function.
-When the current solution is not Pareto stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization.
+When the current solution is not Pareto-stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization [97].
 Therefore, the method used in single-objective optimization can be applied to prove the convergence.
 Furthermore, other algorithms such as CAGrad [19], Nash-MTL [27], and PCGrad [33] also offer convergence analyses in similar approaches.
-TABLE II: Convergence of existing stochastic gradient MOO algorithms. LS: $L$-smooth, BG: bounded gradient, BF: bounded function value, GS: generalized $L$-smooth, CA: conflict-avoidant.
-Method Batch Size Assumption Sample Complexity Bounded CA Distance
+TABLE II: Convergence of existing stochastic gradient MOO algorithms.
+“LS” and “GS” denote the objectives are $L$-smooth and generalized $L$-smooth, respectively.
+“BG” and “BF” represent the bounded gradient and bounded function value assumptions, respectively.
+“Complexity” denotes the sample complexity achieving $\epsilon$-accurate Pareto stationary point.
+“Bounded CA” denotes the bounded conflict-avoidant distance.
+Method Batch Size Assumption Complexity Bounded CA
 SMG [20] ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
@@
 6.1.2 Stochastic Gradient
-Liu and Vicente [91] are the first to analyze the convergence of gradient-based MOO algorithms under stochastic gradients.
-They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto optimal solutions under the assumption of convex objective functions.
+Liu and Vicente [20] are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients.
+They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions.
 However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence.
 This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
-To overcome the limitations associated with increasing batch sizes, Zhou et al.
-[92] propose the Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM) method.
+To overcome this limitation, Zhou et al.
+[21] propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM).
 It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient ${\bm{\lambda}}$.
-In CR-MOGM, the weight coefficient ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
+In CR-MOGM, ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
 ${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$ (41)
 where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration.
-This smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction.
+Smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction.
 Xu et al.
-[94] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM.
+[23] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM.
 They also match the convergence rate in stochastic single-objective optimization [99].
 Another approach to address gradient bias is MoCo [22], which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
-$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right),$ (46)
-where $\prod_{L_{i}}$ denotes the projection to set $\{{\bm{g}}\in\mathbb{R}^{d}~{}~{}|~{}~{}\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$.
-However, the analysis requires the number of iterations $K$ to be sufficiently large (on the order of ${\mathcal{O}}(m^{10})$).
+$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right),$ (42)
+where $\prod_{L_{i}}$ is the projection to the set $\{{\bm{g}}\in\mathbb{R}^{d}~{}~{}|~{}~{}\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$.
+However, the analysis requires the number of iterations $K$ to be sufficiently large (in the order of ${\mathcal{O}}(m^{10})$).
 This high dependency on $m$ makes MoCo less practical for problems involving many objectives.
@@
 [24] propose the Multi-objective gradient with double sampling algorithm (MoDo).
-MoDo mitigates bias in stochastic gradient-based MOO methods without requiring increasing batch sizes or assuming bounded function values.
+MoDo mitigates bias in stochastic gradient-based MOO methods without assuming bounded function values.
 It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients.
@@
 ${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$ (43)
-where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ denotes projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are the stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$.
-By reducing the bias in estimating the common descent direction, MoDo achieves convergence guarantees without bounded function value assumption.
-Additionally, it guarantees a bounded conflict-avoidant (CA) distance, which is the distance between the estimated update direction ${\bm{d}}$ and the conflict-avoidant direction defined in problem (12).
+where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ is the projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$.
+By reducing the bias in estimating the common descent direction, MoDo achieves convergence guarantees without the bounded function value assumption.
+Additionally, it guarantees a bounded conflict-avoidant (CA) distance, which is the distance between the estimated update direction and the CA direction defined in problem (12).
 Similarly, Xiao et al.
-[20] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction, such as the average gradient of all objectives.
-In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated by:
+[25] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives).
+In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated as:
 ${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta\left[{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})^{\top}\left({\bm{G}}^{(k)}({\bm{z}}_{2}^{(k)}){\bm{\lambda}}^{(k)}\right.\right.\right.$
 $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\left.\left.+\gamma{\bm{g}}_{0}({\bm{z}}^{(k)}_{2})\right)\right]\right),$ (44)
-where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$ (e.g., the weighted sum of gradients).
+where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$.
 By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction.
@@
 [26] further consider the convergence under generalized $L$-smoothness [100] and without bounded gradient assumption.
-A summary of convergence of existing stochastic gradient MOO algorithms is provided in Table II.
+A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table II.
 6.2 Generalization
@@
 Sukenik et al.
-[100] consider the generalization for a broader class of scalarization methods.
+[102] consider a broader class of scalarization methods.
 They show that the generalization bounds for individual objectives extend to MOO with scalarization.
 Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
-Another line of research investigates the Tchebycheff scalarization method, which focuses on the sample complexity required to achieve generalization error within $\epsilon$ of the optimal objective value.
+Another line of research investigates Tchebycheff scalarization, which focuses on the sample complexity required to achieve a generalization error within $\epsilon$ of the optimal objective value.
 Formally, given a set of $m$ distributions $\{{\mathcal{D}}_{i}\}_{i=1}^{m}$ and a hypothesis class ${\mathcal{H}}$, the goal is to find a (possibly randomized) hypothesis $h$ such that
 $\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$ (45)
-where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss on ${\mathcal{D}}_{i}$ with hypothesis $h$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance.
+where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss of hypothesis $h$ on ${\mathcal{D}}_{i}$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance.
 It is first formulated by Awasthi et al.
@@
 Haghtalab et al.
-are the first to show the lower bound for sample complexity of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$.
-By using a boosting framework, Peng [102] gives an algorithm that achieves an upper bound for sample complexity of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal.
+are the first to show the sample complexity lower bound of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$.
+By using a boosting framework, Peng [104] gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal.
 In a concurrent work, Zhang et al.
-[103] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$, addressing the open problem raised by Awasthi et al.
-[101].
+[105] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
 In the online setting, Liu et al.
-[104] provide an adaptive online mirror descent algorithm that achieves regret on the order of $O\left(\frac{mv}{\sqrt{K}}\right)$, where $K$ is the number of iterations.
-Using a plain online-to-batch conversion scheme, this algorithm leads to an $O\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$.
+[106] provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations.
+Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ [105].
 It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
-Moreover, MoDo [95] introduces a different perspective by leveraging algorithm stability to derive the first set of algorithm-dependent generalization error bounds of gradient balancing algorithms.
+MoDo [24] introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
 7 Applications
@@
 Reinforcement Learning (RL) [107] involves training agents to make sequential decisions by maximizing cumulative rewards.
-Multi-Objective Reinforcement Learning (MORL) extends single-objective RL by utilizing a vector-valued reward function, i.e., ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$.
+Multi-Objective Reinforcement Learning (MORL) [108] extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$.
 MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
@@
 For instance, Optimal Linear Support (OLS) [110, 111] adds new preference vectors dynamically during each iteration.
-Moreover, Prediction-guided MORL (PGMORL) [109] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
+Prediction-guided MORL (PGMORL) [112] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
 Although Lu et al.
@@
 PPA [114] incorporates a term to align objective values more closely with preference vectors.
-Several studies [112, 113, 114, 115, 116, 117] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto optimal policies.
+Several studies [115, 116, 117, 118, 119, 120] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies.
 Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
 Another research direction focuses on learning the entire Pareto set [121, 122, 123, 124, 125, 126].
-In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and the preference vector ${\bm{\alpha}}$.
+In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and preference vector ${\bm{\alpha}}$.
 Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
 7.2 Bayesian Optimization
-Bayesian Optimization (BO) [124] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model, typically a Gaussian Process, to approximate the objective function and guide the search.
-BO balances exploration (sampling uncertain regions) and exploitation (sampling promising regions) by optimizing an acquisition function like Expected Improvement (EI) [125] or Upper Confidence Bound (UCB) [126].
-When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO).
-MOBO facilitates the exploration of trade-offs between different objectives and aims to approximate the Pareto set of optimal solutions.
-In MOBO, acquisition functions must be designed to manage the balance between exploration and exploitation across multiple objectives.
+Bayesian Optimization (BO) [127] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model.
+When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) [128].
+In MOBO, acquisition functions are designed to manage the balance between exploration and exploitation across multiple objectives.
 Based on this, MOBO methods can be classified into three main categories:
@@
     Scalarization-based Methods.
-These methods convert multi-objective problems into single-objective ones using scalarization functions discussed in Section 2.2 [127, 128].
+These methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 [128, 129].
 Zhang and Golovin [130] propose using the hypervolume indicator for scalarization.
-Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [130], PAL [131, 132], and MOBO-RS [133] extend single-objective acquisitions to multi-objective optimization, leading to improved performance.
+Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [131], PAL [132, 133], and MOBO-RS [134] extend single-objective acquisitions to multi-objective optimization.
   2. 2.
@@
 PESMO [145] reduces uncertainty in the posterior over the Pareto optimal input set.
-Extensions like MESMO [144] and Pareto-Frontier Entropy Search [145] refine this by focusing on output space and the PF.
-Joint Entropy Search (JES)[146, 147] measures combined information gain across input and output spaces.
+Some extensions [146, 147, 148, 149] refine this by considering output space.
 USeMO[150] offers an alternative uncertainty measure based on hyper-rectangle volumes.
@@
 Lin et al.
-[149] propose learning a continuous approximation of Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1).
-By adjusting preferences, decision-makers can flexibly explore any trade-off region within the approximate Pareto set.
-Subsequent improvements have been proposed to leverage diffusion model [150] and stein variational gradient descent [151] for efficient solution generation, as well as incorporate data augmentation for Pareto set model building [152].
+[151] propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1).
+Subsequent improvements have been proposed to leverage diffusion models [152] and Stein variational gradient descent [153] for efficient solution generation, as well as incorporating data augmentation for Pareto set model building [154].
 7.3 Computer Vision
-The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has successfully applied in autonomous driving [153, 154].
-Since the encoder in computer vision usually contains large parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in partial tasks.
-Hence, many methods are proposed to address the issue from the perspective of MOO, which use some scalarization methods to balance multiple losses [155, 156, 157, 158, 159].
-A comprehensive survey on multi-task dense prediction is provided in [4].
-Expect the dense prediction task, MOO is also useful in other computer vision tasks, such as point cloud [160, 161, 162], medical image denoising [163], and pose estimation [164].
+The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving [155, 156].
+Since the encoder in computer vision usually contains large number of parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in some of the tasks.
+Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses [157, 158, 159, 160, 161].
+A comprehensive survey on multi-task dense prediction is in [4].
+Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud [162, 163, 164], medical image denoising [165], and pose estimation [166].
 7.4 Neural Architecture Search
@@
 Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy.
-For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, latency, and energy consumption.
+For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, and latency.
 These works are mainly based on gradient-based NAS methods (like DARTS [168]) and formulate it as a multi-objective optimization problem.
-Among them, [8, 167, 168, 169, 170] employ some scalarization methods to identify a single solution, where task and efficiency objectives are combined with fixed weights.
-However, altering the weights for objectives needs a complete rerun of the search, which is computationally intensive.
+Among them, [8, 169, 170, 171, 172] employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights.
+However, altering the weights of the objectives needs a complete rerun of the search, which is computationally intensive.
 Therefore, Sukthanker et al.
@@
 7.5 Recommendation System
-Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty and diversity, serendipity, popularity, and others [172].
+Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others [174].
 These factors naturally frame the problem as a multi-objective optimization optimization.
@@
 MFTCoder [185] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks.
-CoBa [184] studies the multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
+CoBa [186] studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
 Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects.
 MORLHF [187] and MODPO [188] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models.
-To avoid the LLM retraining, some methods separately train multiple LLMs for each preference dimension and merge their parameters [187, 188] or output logits [189] to deal with different preference requirements in the inference.
+To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters [189, 190] or output logits [191] to deal with different preference requirements.
 However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs.
-To improve the efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [190, 191, 192] or model parameters [54].
+To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [192, 193, 194] or model parameters [63].
 7.7 Miscellaneous
-In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [193, 194], federated learning [195, 196, 197], long-tailed learning [198, 199, 200], continual learning [201], diffusion models [202, 203, 204, 205], neural combinatorial optimization [206, 119, 207, 208, 209], and GFlowNets [210, 211].
+In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [195, 196], federated learning [197, 198, 199], long-tailed learning [200, 201, 202], continual learning [203], diffusion models [204, 205, 206, 207], neural combinatorial optimization [208, 122, 209, 210, 211], GFlowNets [212, 213], and physics informed neural networks (PINNs) [214, 215].
 These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
 8 Resources
-8.1 Benchmark Datasets
-In this section, we introduce some benchmark datasets for multi-objective deep learning. They are summarized in Table III and listed in detail as follows.
+In this section, we introduce some benchmark datasets and open-source libraries for gradient-based MOO in deep learning. Benchmark datasets are summarized in Table III, with details in Appendix C.
+Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL [216] and LibMOON [^2] https://github.com/xzhang2523/libmoon [217], provide unified environments for implementing and fairly evaluating MOO methods.
+Both are built on PyTorch [218] and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios.
+LibMTL [216] mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets.
+LibMOON [217] mainly focuses on exploring the whole Pareto set.
+It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
+TABLE III: Summary of benchmark datasets in multi-objective deep learning.
+Dataset Description #Objectives #Samples
+NYUv2 [219] indoor scene understanding $3$ $1,449$
+Taskonomy [220] indoor scene understanding $26$ $\approx 4$M
+Cityscapes [221] urban scene understanding $2$ $3,475$
+QM9 [222] molecular property prediction $11$ $130$K
+Office-31 [223] image classification $3$ $4,110$
+Office-Home [224] image classification $4$ $15,500$
+CelebA [225] image classification $40$ $\approx 202$K
+CIFAR-100 [226] image classification $20$ $60$K
+MT10/MT50 [227] reinforcement learning $10/50$
+XTREME [228] multilingual learning $9$ $\approx 597$K
+PKU-SafeRLHF [229] two-dimensional preference data $2$ $\approx 82$K
+UltraFeedBack [230] multi-dimensional preference data $4$ $\approx 64$K
+HelpSteer2 [231] multi-attributes preference data $5$ $\approx 21$K
+9 Challenges and Future Directions
+Despite significant progress in applying MOO in deep learning, both in learning a single model and in learning a Pareto set of models, several challenges remain.
+Theoretical Understanding. While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
+Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received less attention.
+For instance, Chen et al.
+[24], as discussed in Section 6.2, provide the algorithm-dependent generalization analysis.
+Extending this to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
+For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets.
+Zhang et al.
+[232] established a generalization bound for HV-based Pareto set learning.
+However, the effect of various network architectures (Section 5.1) on approximating Pareto sets remain unclear.
+Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
+Reducing Gradient Balancing Costs.
+While gradient balancing methods in Section 3.2 often show strong performance in various applications, they come with significant computational overhead.
+Despite the introduction of some practical speedup strategies (discussed in Section 3.2.3), existing approaches remain insufficiently efficient.
+A deeper understanding of the optimization differences between gradient balancing, linear scalarization, and loss balancing is crucial.
+This insight could facilitate the integration of gradient balancing with linear scalarization and loss balancing, reducing computational overhead significantly and enabling its application in large-scale training scenarios.
+Dealing with Large Number of Objectives.
+Some real-world problems involve handling a large number of objectives, which poses significant challenges for current Pareto set learning algorithms in Section 5.
+As the number of objectives grows, the preference vector space expands exponentially, making it challenging for existing random sampling-based techniques to effectively learn the mapping between preference vectors and solutions.
+Developing efficient sampling strategies for high-dimensional preference spaces remains a important research problem.
+Additionally, exploring methods to reduce or merge objectives based on their properties can be a promising approach to minimize the total number of objectives.
+Distributed Training.
+Current MOO algorithms in deep learning are limited to single GPUs or machines, but scaling them to multi-GPU and distributed environments is increasingly important as models and datasets grow.
+This scaling introduces unique challenges compared to single-objective optimization.
+Gradient-balancing algorithms require efficient gradient distribution and communication across GPUs.
+Learning a set of Pareto-optimal solutions involves synchronizing multiple models across nodes with minimal overhead.
+Additionally, when data for different objectives is distributed across devices and cannot be shared (e.g., due to privacy constraints), collaborative computation of solutions or Pareto sets without direct data sharing becomes an important challenge.
+Advancements in Large Language Models (LLMs).
+As discussed in Section 7.6, current research in MOO for LLMs largely focuses on the RLHF stage.
+Expanding MOO techniques to other stages in the LLM lifecycle, such as pre-training and inference, is a valuable direction for future research.
+Addressing these challenges can lead to models that are better aligned with user needs across their entire development processes.
+Additionally, user preferences are currently modeled as a preference vector.
+However, this representation may oversimplify more complex user preferences on LLMs.
+Exploring advanced methods to capture and represent these intricate preferences can enhance the effectiveness and personalization of LLMs.
+Application in More Deep Learning Scenarios.
+While MOO has already been utilized in various deep learning scenarios, as highlighted in Section 7, there remain numerous unexplored areas within this field.
+Multi-objective characteristics are inherently present in most deep learning problems, as models are typically developed or evaluated based on multiple criteria.
+Consequently, trade-offs often arise naturally.
+Leveraging MOO methods to effectively address these trade-offs presents an opportunity for further exploration.
+10 Conclusion
+In this paper, we provide the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theoretical foundations, applications, benchmark datasets, and software.
+By unifying diverse approaches and systematically identifying key challenges, this paper serves as a foundational resource to drive future advancements in this important and evolving field.
+Appendix A Illustration of Key Concepts in MOO
+Figure 6 illustrates the Pareto concepts on the two-objective problem.
+The blue circles represent Pareto optimal solutions (defined in Definition 2), which indicates that no solution can improve one objective without worsening the other.
+The blue curve connecting the blue circles denotes the Pareto front, as defined in Definition 3.
+The yellow circles indicate weak Pareto optimal solutions since they can be improved in one objective without negatively impacting the other.
+For example, comparing ${\bm{\theta}}_{2}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{2}$ is a weak Pareto optimal solution since it can be improved in the second objective without affecting the first.
+The red circles represent solutions that are not Pareto optimal because there exists another solution that strictly Pareto dominates them.
+For instance, comparing ${\bm{\theta}}_{3}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{3}$ is not a Pareto optimal solution as ${\bm{\theta}}_{1}$ outperforms it in both objectives.
+Figure 6: Illustration of Pareto optimality concepts.
+The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively.
+The blue curve represents the Pareto front.
+Appendix B Additional Details and Further Results for the Experiment in Fig. 4
+To demonstrate the differences in the solution distributions produced by various methods, we consider a commonly used benchmark problem, LSMOP1 [84].
+Each algorithm is run for 5000 iterations.
+The solutions obtained by PTML [36], EPO [37, 38], GradHV [47], and MOO-SVGD [48] are shown in Figure 4 in the main paper.
+Additionally, the solutions generated by MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18] are shown in Figure 7.
+From Figures 4 and 7, we can observe that MGDA, which does not incorporate preferences, exhibits poor diversity.
+The constraint-based approach in PTML produces solutions that are better aligned with the preference vector but still fall short of exact alignment.
+EPO and PMGDA achieves exact alignment with the preference vector.
+Some methods that do not use a preference vector, such as GradHV and MOO-SVGD, are also able to produce diverse solutions.
+Figure 7: Distributions of solutions obtained by using MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18].
+The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
+Appendix C Details of Benchmark Datasets
+In this section, we list the details of benchmark datasets in Section 8.
   • NYUv2 dataset [219] is for indoor scene understanding.
@@
 Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 00 and $4$ where higher means better for each attribute.
-TABLE III: Summary of benchmark datasets in multi-objective deep learning.
-Dataset Description #Objectives #Samples
-NYUv2 [212] indoor scene understanding $3$ $1,449$
-Taskonomy [213] indoor scene understanding $26$ $\approx 4$M
-Cityscapes [214] urban scene understanding $2$ $3,475$
-QM9 [215] molecular property prediction $11$ $130$K
-Office-31 [216] image classification $3$ $4,110$
-Office-Home [217] image classification $4$ $15,500$
-CelebA [218] image classification $40$ $\approx 202$K
-CIFAR-100 [219] image classification $20$ $60$K
-MT10/MT50 [220] reinforcement learning $10/50$
-XTREME [221] multilingual learning $9$ $\approx 597$K
-PKU-SafeRLHF [222] two-dimensional preference data $2$ $\approx 82$K
-UltraFeedBack [223] multi-dimensional preference data $4$ $\approx 64$K
-HelpSteer2 [224] multi-attributes preference data $5$ $\approx 21$K
-8.2 Software
-In this section, we introduce two open-source libraries for gradient-based multi-objective deep learning: LibMTL [^1] https://github.com/median-research-group/LibMTL [225] and LibMOON [^2] https://github.com/xzhang2523/libmoon [226].
-These two libraries provide a unified environment for users to implement or fairly evaluate different MOO methods.
-Moreover, they are built on PyTorch [227] and designed in a modular fashion, allowing users to flexibly develop novel methods or apply existing MOO methods to new application scenarios.
-LibMTL [225] mainly focuses on finding a single Pareto optimal solution and supports over $19$ methods introduced in Section 3.
-Meanwhile, it also supports $6$ benchmark datasets (i.e., NYUv2, Cityscapes, QM9, Office-31, Office-Home, and XTREME datasets) introduced in Section 8.1.
-LibMOON [226] mainly focuses on exploring the whole Pareto set.
-It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
-These methods address a variety of tasks, including classification, fairness, and distribution matching.
-9 Challenges and Future Directions
-Despite significant progress in applying MOO in deep learning, both in learning a single model and in learning a Pareto set of models, several challenges remain.
-9.1 Theoretical Understanding
-While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
-Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received much less attention.
-For instance, Chen et al.
-[95], as discussed in Section 6.2, provide the first algorithm-dependent generalization analysis for MOO.
-Extending this work to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
-For Pareto set learning algorithms, there is limited theoretical understanding of how effectively existing algorithms can approximate Pareto sets.
-Zhang et al.
-[228] established a generalization bound for hypervolume-based Pareto set learning.
-However, the effect of various network architectures (Section 5.1) on approximating Pareto sets and convergence rate remain unclear.
-Additionally, it is uncertain how effective current training strategies (Section 5.2) are in approximating the Pareto set.
-9.2 Reducing Gradient Balancing Costs
-While gradient balancing methods in Section 3.2 often show strong performance in various applications, they come with significant computational overhead.
-Despite the introduction of some practical speedup strategies (discussed in Section 3.2.3), existing approaches remain insufficiently efficient.
-A deeper understanding of the optimization differences between gradient balancing, linear scalarization, and loss balancing is crucial.
-This insight could facilitate the integration of gradient balancing with linear scalarization and loss balancing, reducing computational overhead significantly and enabling its application in large-scale training scenarios.
-9.3 Dealing with Large Number of Objectives
-Some real-world problems involve handling a large number of objectives, which poses significant challenges for current Pareto set learning algorithms in Section 5.
-As the number of objectives grows, the preference vector space expands exponentially, making it challenging for existing random sampling-based techniques to effectively learn the mapping between preference vectors and solutions.
-Developing efficient sampling strategies for high-dimensional preference spaces remains a critical research problem.
-Additionally, exploring methods to reduce or merge objectives based on their properties could be a promising approach to minimize the total number of objectives.
-9.4 Distributed Training
-Current MOO algorithms in deep learning are limited to single GPUs or machines, but scaling them to multi-GPU and distributed environments is increasingly important as models and datasets grow.
-This scaling introduces unique challenges compared to single-objective optimization.
-Gradient-balancing algorithms require efficient gradient distribution and communication across GPUs.
-Learning a set of Pareto optimal solutions involves synchronizing multiple models across nodes with minimal overhead.
-Additionally, when data for different objectives is distributed across devices and cannot be shared (e.g., due to privacy or memory constraints), collaborative computation of solutions or Pareto sets without direct data sharing becomes an important challenge.
-9.5 Advancements in Large Language Models (LLMs)
-As discussed in Section 7.6, current research in MOO for LLMs largely focuses on the RLHF stage.
-Expanding MOO techniques to other stages in the LLM lifecycle, such as pre-training and inference, is a valuable direction for future research.
-Addressing these challenges could lead to models that are better aligned with user needs across their entire development processes.
-Additionally, user preferences are currently modeled as a preference vector.
-However, this representation may oversimplify more complex user preferences on LLMs.
-Exploring advanced methods to capture and represent these intricate preferences could enhance the effectiveness and personalization of LLMs.
-9.6 Application in More Deep Learning Scenarios
-While MOO has already been utilized in various deep learning scenarios, as highlighted in Section 7, there remain numerous unexplored areas within this field.
-Multi-objective characteristics are inherently present in most deep learning problems, as methods are typically developed or models evaluated based on multiple criteria.
-Consequently, trade-offs often arise naturally.
-Leveraging MOO methods to effectively address these trade-offs presents a significant opportunity for further exploration.
-10 Conclusion
-In this paper, we provide the first comprehensive review of gradient-based multi-objective optimization in deep learning, covering algorithms, theoretical foundations, practical applications, benchmark datasets, and software.
-By unifying diverse approaches and systematically identifying key challenges, this paper serves as a foundational resource to drive future advancements in this important and evolving field.

### v3

@@
 Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond
-Abstract
-Multi-objective optimization (MOO) in deep learning aims to simultaneously optimize multiple conflicting objectives, a challenge frequently encountered in areas like multi-task learning and multi-criteria learning.
-Recent advancements in gradient-based MOO methods have enabled the discovery of diverse types of solutions, ranging from a single balanced solution to finite or even infinite Pareto sets, tailored to user needs.
-These developments have broad applications across domains such as reinforcement learning, computer vision, recommendation systems, and large language models.
-This survey provides the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theories, and practical applications.
-By unifying various approaches and identifying critical challenges, it serves as a foundational resource for driving innovation in this evolving field.
-A comprehensive list of MOO algorithms in deep learning is available at https://github.com/Baijiong-Lin/Awesome-Multi-Objective-Deep-Learning.
-Index Terms:
+Abstract.
+Many modern deep learning applications require balancing multiple objectives that are often conflicting.
+Examples include multi-task learning, fairness-aware learning, and the alignment of Large Language Models (LLMs).
+This leads to multi-objective deep learning, which tries to find optimal trade-offs or Pareto-optimal solutions by adapting mathematical principles from the field of Multi-Objective Optimization (MOO).
+However, directly applying gradient-based MOO techniques to deep neural networks presents unique challenges, including high computational costs, optimization instability, and the difficulty of effectively incorporating user preferences.
+This paper provides a comprehensive survey of gradient-based techniques for multi-objective deep learning.
+We systematically categorize existing algorithms based on their outputs: (i) methods that find a single, well-balanced solution, (ii) methods that generate a finite set of diverse Pareto-optimal solutions, and (iii) methods that learn a continuous Pareto set of solutions.
+In addition to this taxonomy, the survey covers theoretical analyses, key applications, practical resources, and highlights open challenges and promising directions for future research.
+A comprehensive list of multi-objective deep learning algorithms is available at https://github.com/Baijiong-Lin/Awesome-Multi-Objective-Deep-Learning.
 Multi-Objective Optimization, Multi-Task Learning, Pareto Set Learning, Deep Learning
-1 Introduction
-Multi-Objective Optimization (MOO) [1], which aims to optimize multiple potentially conflicting objectives simultaneously, has become an important topic in deep learning.
-A multi-objective optimization problem can be formulated as:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$ (1)
-where $m\geq 2$ is the number of objectives, ${\bm{\theta}}$ is the decision variable and ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable.
-Figure 1: An overview of multi-objective deep learning.
-Different from traditional single-objective optimization, which focuses on improving a single metric, MOO aims to obtain different solutions that achieve different trade-offs among objectives.
-Particularly, in MOO, users can specify their preferences regarding the relative importance of objectives, typically represented by a preference vector ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where each $\alpha_{i}$ indicates the importance of the $i$-th objective and $\Delta_{m-1}$ denotes a $(m-1)$-dimensional simplex.
-MOO is expected to achieve a solution that matches the user-defined preference.
-Hence, compared to single-objective optimization, MOO is more practical for complex real-world applications.
-Originally arising from operations research, MOO has been extensively studied across various disciplines of science and engineering [2].
-In recent years, there has been a surge of interest in applying MOO techniques to deep learning, leading to the reformulation or extension of many learning paradigms as MOO problems.
-In deep learning, there are generally two primary types of MOO scenarios.
-One scenario is multi-task learning [3], where a model is trained to handle multiple tasks simultaneously.
-For example, in computer vision, we may desire a single network to perform various dense prediction tasks concurrently, such as segmentation, depth estimation, and surface normal prediction [4].
-In reinforcement learning [5], we might aim for a model that performs well on several robotic manipulation tasks.
-In the training of Large Language Models (LLMs), we expect the model to excel at different tasks, such as mathematics, reasoning, language translation, and summarization.
-However, these tasks often compete for shared network resources, leading to conflicts that can degrade performance on individual tasks.
-By formulating the problem as an MOO, where the performance on each task is treated as a separate objective [6], we can systematically manage these conflicts and optimize all task objectives simultaneously.
-Another scenario involves incorporating additional evaluation criteria, such as fairness, safety, harmfulness mitigation, and efficiency, into learning paradigms.
-Considering these objectives is essential for practical real-world applications but may conflict with the primary goal of maximizing performance.
-For example, in LLMs training [7], we aim for models that not only perform well on tasks like language understanding but also adhere to safety standards, avoid producing harmful content, and ensure fairness across different users.
-In hardware-aware Neural Architecture Search (NAS) [8], we aim to discover network architectures that achieve high accuracy while also being efficient in terms of computational resources, energy consumption, or parameter size.
-However, reducing model size or computational complexity often leads to decreased performance.
-By treating such problems as MOO, where each criterion is an objective, we provide a structured way to explore these trade-offs.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Machine learningccs: Computing methodologies Multi-task learning
+1. Introduction
+Traditional deep learning often focuses on optimizing a single learning objective, such as minimizing the prediction error or maximizing the likelihood.
+However, many real-world applications require balancing multiple, often conflicting, objectives.
+For instance, a computer vision system might need to perform the tasks of segmentation, depth estimation, and surface normal prediction simultaneously (Vandenhende et al., 2021), thus moving from single-task learning to multi-task learning (Zhang and Yang, 2022).
+Similarly, Large Language Models (LLMs) are expected to excel at diverse tasks like reasoning and coding, while also being safe, fair, and harmless (Wang et al., 2023b).
+Multi-Objective Optimization (MOO) (Miettinen, 1999), a field originating from operations research (Ehrgott, 2005), provides a formal framework for navigating these trade-offs.
+It has been widely studied across science and engineering, with diverse applications such as in finance (Steuer and Na, 2003), engineering design (Marler and Arora, 2004), and transportation planning (Tzeng et al., 2005).
+Recently, there has been a surge of interests in adapting MOO for deep learning, re-framing many popular learning problems as multi-objective optimization problems.
+Examples include multi-task learning where each task performance is considered as an objective (Sener and Koltun, 2018), fairness-aware learning where accuracy is one objective and fairness metrics constitute the other objectives (Martinez et al., 2020), LLM alignment where each alignment criterion (such as helpfulness, harmlessness, and honesty) represents a separate objective (Wang et al., 2023b), and federated learning where the performance on each client is treated as an individual objective (Hu et al., 2022).
+All these problems can be formally expressed as:
+(1) $\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~~{\bm{f}}({\bm{\theta}})$ $:=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$
+where $m\geq 2$ is the number of objectives, $\bm{\theta}$} is the decision variable, ${\mathcal{K}}\subset{\mathbb{R}}^{d}$ is the feasible set of the decision variable, and each $f_{i}:{\mathcal{K}}\to{\mathbb{R}}$ represents an individual objective function to be minimized.
+Unlike single-objective optimization, which focuses on finding a single best solution, MOO recognizes that no single solution is optimal for all objectives simultaneously.
+Instead, MOO aims to identify a set of solutions that represent different trade-offs between objectives, collectively known as the Pareto set (Miettinen, 1999).
+The Pareto set consists of all solutions where improving any objective would necessarily worsen at least one other objective, representing the best possible trade-offs available.
+In real-world applications, users often have varying preferences for these trade-offs.
+For instance, in the development of LLMs, a customer service application might emphasize harmlessness and safety, whereas an educational tool might prioritize reasoning and factual accuracy.
+To address this variability, users can specify their preferences using a vector, ${\bm{\alpha}}=[\alpha_{1},\dots,\alpha_{m}]^{\top}\in\Delta_{m-1}$, where $\Delta_{m-1}=\{{\bm{\alpha}}\in{\mathbb{R}}^{m}_{+}:\sum_{i=1}^{m}\alpha_{i}=1\}$ is the probability simplex and each $\alpha_{i}$ represents the importance assigned to the $i$-th objective.
+MOO approaches can accommodate these user-defined preferences, enabling the discovery of solutions that align with individual needs.
+MOO methods can be broadly divided into gradient-free and gradient-based approaches.
+Gradient-free methods, such as evolutionary algorithms (Deb et al., 2002; Zhang and Li, 2007) and particle swarm optimization (Coello et al., 2004), explore the solution space using population-based sampling.
+While effective for traditional low-dimensional MOO problems (Deb, 2011), they struggle with deep neural networks, which often have millions or even billions of parameters.
+Due to the curse of dimensionality, the parameter space becomes too large to explore efficiently without gradient information, leading to slow or failed convergence.
+This makes gradient-based methods (Désidéri, 2012; Mukai, 1980; Fliege and Svaiter, 2000; Liu et al., 2021b) preferable for deep neural networks, as they efficiently guide the search using objective gradients.
+However, gradient-based MOO faces several significant challenges in deep learning.
+Firstly, incorporating user preferences is difficult because directly weighting objectives often produces unaligned solutions.
+Secondly, training even one neural network is computationally expensive, posing a major challenge to efficiently approximating the entire Pareto set of networks.
+Finally, the use of mini-batch optimization in deep learning introduces high variance, which can further destabilize the optimization process.
+These challenges have motivated extensive research efforts to develop specialized MOO techniques for deep learning.
+This paper provides a comprehensive survey of these advances, systematically categorizing existing approaches and identifying key open problems in the field.
+1.1. Comparison with Related Surveys on MOO
+Some surveys have attempted to cover aspects of MOO.
+For example, Zhang and Yang (Zhang and Yang, 2022) provide a comprehensive overview of traditional multi-task methods within the machine learning domain.
+Crawshaw (Crawshaw, 2020) offers a summary of multi-task methods in deep learning up to the year 2020.
+Yu et al.
+(Yu et al., 2024) present a comprehensive survey that covers multi-task learning from traditional machine learning to deep learning.
+Eichfelder (Eichfelder, 2021) discusses advancements in MOO from a general optimization perspective, without specific emphasis on deep learning.
+Wei et al.
+(Wei et al., 2021) focus on evolutionary MOO algorithms, but these algorithms are usually not suitable for deep neural networks due to their large parameter spaces and computational demands.
+Peitz and Hotegni (Peitz and Hotegni, 2024) survey the MOO algorithms for deep learning but cover only a limited selection of gradient-based MOO methods and do not include theory and applications.
+Other surveys concentrate on applications of MOO in areas such as dense prediction tasks (Vandenhende et al., 2021), recommend systems (Zhang et al., 2023b), and natural language processing (Chen et al., 2024b).
+Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications, as the overview shown in Figure 1.
+To the best of our knowledge, it is the first survey paper focusing on the gradient-based MOO methods in deep learning.
+Figure 1. An overview of gradient-based multi-objective deep learning.
 {forest}
-for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA [9]; UW [10]; IMTL-L [11]; Lin et al.
-[12]; MOML [13, 14]; Auto-$\lambda$ [15]; FORUM [16]; RW [17]; STCH [18], single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA [6]; CAGrad [19]; SMG [20]; CR-MOGM [21]; MoCo [22]; PSMGD [23]; MoDo [24]; SDMGrad [25]; SGSMGrad [26]; IMTL-G [11]; Nash-MTL [27]; FairGrad [28]; UPGrad [29]; Aligned-MTL [30]; GradNorm [31]; DB-MTL [12]; FAMO [32], single_work] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad [33]; GradVac [34]; GradDrop [35], single_work] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL [36]; EPO [37, 38]; WC-MGDA [39]; PMGDA [40]; FERERO [41]; GMOOAR [42]; UMOD [43], finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV [44, 45, 46, 47]; MOO-SVGD [48]; F4M [49]; SoM [50]; MosT [51], finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN [52]; CPMTL [53]; PHN-HVI [54]; SUHNPF [55]; Hyper-Trans [56], infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS [57]; YOTO [58]; Raychaudhuri et al.
-[59], infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL [60]; LORPMAN [61]; Chen and Kwok [62]; Panacea [63]; Tang et al.
-[64]; PaLoRA [65], infinite_work] ] ] ]
-Figure 2: Taxonomy of existing MOO algorithms in deep learning.
-1.1 Taxonomy
-Instead of the traditional approach of directly summing all objectives into a single scalar function, MOO algorithms in deep learning can be categorized into three main approaches:
-  • To obtain a single, well-balanced solution, techniques like loss balancing and gradient balancing are employed to resolve conflicts between objectives during optimization.
-These methods focus on adjusting the weight of each objective to achieve a solution that is not overly biased toward any one objective.
-  • To obtain a finite set of Pareto optimal solutions, methods either divide the problem into subproblems using preference vectors or directly optimize for multiple solutions simultaneously.
-This allows users to select from a set of solutions that best meet their specific needs.
-  • To generate an infinite set of Pareto optimal solutions, the focus shifts to learning a subspace of solutions rather than individual ones.
-Algorithms in this category propose efficient structures to learn this subspace.
-During training, preference vectors are randomly sampled in each iteration.
-Once a preference is sampled, the optimization reduces to solving for either a single solution or a finite set, making it possible to use techniques from the first two approaches.
-Figure 2 illustrates the taxonomy of current MOO algorithms in deep learning, serving as a guide to the methods discussed in this paper.
-By exploring these avenues, MOO provides a comprehensive framework for addressing complex optimization problems in deep learning, enabling the development of models that can effectively balance multiple competing objectives.
-1.2 Comparison with Related Surveys
-Some surveys have attempted to cover aspects of MOO.
-For example, Zhang and Yang [3] provide a comprehensive overview of traditional multi-task methods within the machine learning domain.
-Crawshaw [66] offers a summary of multi-task methods in deep learning up to the year 2020.
-Yu et al.
-[67] present a comprehensive survey that covers multi-task learning from traditional machine learning to deep learning.
-Eichfelder [68] discusses advancements in MOO from a general optimization perspective, without specific emphasis on deep learning.
-Wei et al.
-[69] focus on evolutionary MOO algorithms, but these algorithms are usually not suitable for deep neural networks due to their large parameter spaces and computational demands.
-Peitz and Hotegni [70] survey the MOO algorithms for deep learning but cover only a limited selection of gradient-based MOO methods and do not include theory and applications.
-Other surveys concentrate on applications of MOO in areas such as dense prediction tasks [4], recommend systems [71], and natural language processing [72].
-Different with the previous surveys, this survey focuses on recent gradient-based MOO methods in deep learning, encompassing approaches for finding a single solution, a finite set of solutions, and an infinite set of solutions, and covering algorithms, theories, and applications.
-To the best of our knowledge, it is the first survey paper to focus on the gradient-based MOO methods in deep learning.
-TABLE I: Summary of Notations.
+for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, , [ Gradient-based Multi-objective Deep Learning , fill=gray!45, parent [Finding a Single Solution (§3), for tree=single [Loss Balancing Methods (§3.1), single [DWA (Liu et al., 2019a); UW (Kendall et al., 2018); IMTL-L (Liu et al., 2021a); Lin et al.
+(Lin et al., 2023); MOML (Ye et al., 2021, 2024b); Auto-$\lambda$ (Liu et al., 2022); FORUM (Ye et al., 2024a); RW (Lin et al., 2022c); STCH (Lin et al., 2024b), single_work] ] [Gradient Balancing Methods (§3.2), single [Gradient Weighting Methods (§3.2.1), single [MGDA (Sener and Koltun, 2018); CAGrad (Liu et al., 2021b); SMG (Liu and Vicente, 2021); CR-MOGM (Zhou et al., 2022); MoCo (Fernando et al., 2023); PSMGD (Xu et al., 2025b); MoDo (Chen et al., 2023a); SDMGrad (Xiao et al., 2023); SGSMGrad (Zhang et al., 2025); IMTL-G (Liu et al., 2021a); Nash-MTL (Navon et al., 2022); FairGrad (Ban and Ji, 2024); UPGrad (Quinton and Rey, 2024); Aligned-MTL (Senushkin et al., 2023); GradNorm (Chen et al., 2018); DB-MTL (Lin et al., 2023); FAMO (Liu et al., 2023), single_work2] ] [Gradient Manipulation Methods (§3.2.2), single [PCGrad (Yu et al., 2020a); GradVac (Wang et al., 2021b); GradDrop (Chen et al., 2020), single_work2] ] ] ] [Finding a Finite Set of Solutions (§4), for tree=finite [Methods Based on Preference Vectors (§4.1), finite [PMTL (Lin et al., 2019b); EPO (Mahapatra and Rajan, 2020, 2021); WC-MGDA (Momma et al., 2022); PMGDA (Zhang et al., 2024b); FERERO (Chen et al., 2024a); GMOOAR (Chen and Kwok, 2022); UMOD (Zhang et al., 2024a), finite_work] ] [Methods without Using Preference Vectors (§4.2), finite [GradHV (Deist et al., 2021, 2020; Emmerich et al., 2007; Wang et al., 2017); MOO-SVGD (Liu et al., 2021d); F4M (Lin et al., 2025c); SoM (Ding et al., 2024); MosT (Li et al., 2025), finite_work] ] ] [Finding an Infinite Set of Solutions (§5), for tree=infinite [Hypernetwork-based Methods (§5.1.1), infinite [PHN (Navon et al., 2021); CPMTL (Lin et al., 2020); PHN-HVI (Hoang et al., 2023); SUHNPF (Gupta et al., 2022); Hyper-Trans (Tuan et al., 2024), infinite_work] ] [Preference-Conditioned Network-based Methods (§5.1.2), infinite [COSMOS (Ruchte and Grabocka, 2021); YOTO (Dosovitskiy and Djolonga, 2020); Raychaudhuri et al.
+(Raychaudhuri et al., 2022), infinite_work] ] [Model Combination-based Methods (§5.1.3), infinite [PaMaL (Dimitriadis et al., 2023); LORPMAN (Chen and Kwok, 2024a); Chen and Kwok (Chen and Kwok, 2024b); Panacea (Zhong et al., 2024); Tang et al.
+(Tang et al., 2024); PaLoRA (Dimitriadis et al., 2025), infinite_work] ] ] ]
+Figure 2. Taxonomy of existing gradient-based multi-objective deep learning algorithms.
+1.2. Organization of the Survey
+In this survey, we categorize gradient-based multi-objective deep learning algorithms based on the type of output they obtain: (Figure 2): (i) Algorithms that obtain a single, well-balanced model that is not overly biased toward any one objective: This is achieved through techniques like loss balancing and gradient balancing, which adjust the weight of each objective to resolve conflicts during optimization.
+(ii) Algorithms that obtain a finite set of Pareto-optimal solutions that allows users to select from multiple options based on their specific needs: This is accomplished by either dividing the problem into subproblems using preference vectors or directly optimizing for multiple solutions simultaneously.
+(iii) Algorithms that obtain a continuous set of Pareto-optimal models that can be generated on-demand for any given preference: This is achieved by learning a solution subspace using efficient structures, where preference vectors are randomly sampled during training and the optimization adapts techniques from single or finite solution methods.
+The rest of this survey is organized as follows: Section 2 presents MOO preliminaries, including MOO definitions and concepts; Section 3 discusses methods for finding a single Pareto-optimal solution, covering loss balancing and gradient balancing approaches; Section 4 focuses on methods for identifying a set of finite Pareto-optimal solutions; Section 5 covers methods for identifying a set of infinite Pareto optimal solutions; Section 6 delves into the theoretical analysis of convergence and generalization in gradient-based multi-objective optimization algorithms; Section 7 showcases various applications in deep learning, including reinforcement learning, Bayesian optimization, computer vision, neural architecture search, recommender systems, and large language models; Section 8 offers useful resources on datasets and software tools for multi-objective deep learning; Section 9 highlights challenges in the field and suggests promising directions for future research; Section 10 summarizes this survey.
+1.3. Notations
+The notations used in this survey are summarized in Table 1.
+Scalars are represented by non-bold letters; vectors and matrices are denoted by boldface type.
+Subscripts are used as indices for elements, and superscripts typically denote iteration numbers or indices within a solution set.
+Table 1. Summary of Notations.
 Notation Description
@@
 $[m]$ Index set $\{1,\dots,m\}$.
-$\Delta_{m-1}$ $(m-1)$-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
-[…]
+$\Delta_{m-1}$ (m1)(m-1)-D simplex $\{{\bm{\alpha}}|\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0,i\in[m]\}$.
+$\|\cdot\|$ $\ell_{2}$ norm.
 $\eta$ Step size for updating $\bm{\theta}$}.
 $\epsilon$ Error tolerance.
-1.3 Organization of the Survey
-The notations used in this survey are summarized in Table I. An overview of MOO in deep learning is provided in Figure 1. The rest of this survey is organized as follows:
-  • Section 2 presents preliminaries in MOO, including key definitions and concepts;
-  • Section 3 discusses methods for finding a single Pareto optimal solution, covering loss balancing and gradient balancing approaches;
-  • Sections 4 and 5 focus on methods for identifying a finite and an infinite set of Pareto optimal solutions, respectively;
-  • Section 6 delves into the theoretical analysis of convergence and generalization in gradient-based multi-objective optimization algorithms;
-  • Section 7 showcases various applications in deep learning, including reinforcement learning, Bayesian optimization, computer vision, and large language models;
-  • Section 8 offers useful resources for multi-objective deep learning, such as datasets and software tools;
-  • Section 9 highlights challenges in the field and suggests promising directions for future research;
-  • Section 10 summarizes this survey.
-2 Preliminary: Key Concepts in MOO
-In this section, we first introduce some Pareto concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2.
-An illustration of some Pareto concepts is in Appendix A.
-2.1 Pareto Concepts
+2. Preliminary: Multi-Objective Optimization
+In this section, we first introduce some MOO concepts in Section 2.1 and then review two classical methods (i.e., linear scalarization and Tchebycheff scalarization) to find Pareto solutions of MOO problems in defined in Section 2.2.
+2.1. Key Concepts in Multi-Objective Optimization
 Unlike single-objective optimization, solutions in MOO cannot be directly compared based on a single criterion, but are compared using the concept of dominance as follows.
 Note that without loss of generality, we consider multi-objective minimization (i.e., problem (1)) here.
-Definition 1 ((strict) Pareto dominance [1])
+Definition 2.1 ((strict) Pareto dominance (Miettinen, 1999)).
 A solution ${\bm{\theta}}^{(a)}$ dominates another solution ${\bm{\theta}}^{(b)}$ (denoted as ${\bm{\theta}}^{(a)}\preceq{\bm{\theta}}^{(b)}$) if and only if $f_{i}({\bm{\theta}}^{(a)})\leq f_{i}({\bm{\theta}}^{(b)})$ for all $i\in[m]$, and there exists at least one $i\in[m]$ such that $f_{i}({\bm{\theta}}^{(a)})<f_{i}({\bm{\theta}}^{(b)})$.
@@
 Based on this definition, we further define Pareto optimality (PO), Pareto set, and Pareto front as follows.
-Definition 2 ((weak) Pareto optimality [1])
+Definition 2.2 ((weak) Pareto optimality (Miettinen, 1999)).
 A solution $\bm{\theta}$$^{*}$ is Pareto optimal if no other solution dominates it. A solution $\bm{\theta}$$^{*}$ is weakly Pareto optimal if no other solution strictly dominates it.
-Definition 3 (Pareto set (PS) and Pareto front (PF) [1])
+Definition 2.3 (Pareto set (PS) and Pareto front (PF) (Miettinen, 1999)).
 A PS is the set of all PO solutions. A PF is the set of all objective function values of the PO solutions.
-Definition 4 (Preference vector)
+Figure 3 illustrates the Pareto concepts on the two-objective problem.
+The blue circles represent Pareto optimal solutions (defined in Definition 2.2), which indicates that no solution can improve one objective without worsening the other.
+The blue curve connecting the blue circles denotes the Pareto front, as defined in Definition 2.3.
+The yellow circles indicate weak Pareto optimal solutions since they can be improved in one objective without negatively impacting the other.
+For example, comparing ${\bm{\theta}}_{2}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{2}$ is a weak Pareto optimal solution since it can be improved in the second objective without affecting the first.
+The red circles represent solutions that are not Pareto optimal because there exists another solution that strictly Pareto dominates them.
+For instance, comparing ${\bm{\theta}}_{3}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{3}$ is not a Pareto optimal solution as ${\bm{\theta}}_{1}$ outperforms it in both objectives.
+Figure 3.
+Illustration of Pareto optimality concepts.
+The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively.
+The blue curve represents the Pareto front.
+Definition 2.4 (Preference vector).
 $\bm{\alpha}$} denotes a preference vector.
@@
 In the case of MOO, the solution generally converges to a Pareto stationary solution, which is defined as follows:
-Definition 5 (Pareto stationary [73])
+Definition 2.5 (Pareto stationary (Désidéri, 2012)).
 A solution $\bm{\theta}$$^{*}$ is called Pareto stationary if there exists a vector ${\bm{\lambda}}\in\Delta_{m-1}$ such that $\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}({\bm{\theta}}^{*})\|=0$.
 Pareto stationarity is a necessary condition for achieving Pareto optimality.
-If all objectives are convex with $\lambda_{i}>0,\forall i$, it also serves as the Karush-Kuhn-Tucker (KKT) sufficient condition for Pareto optimality [73].
-When evaluating the performance of a set of obtained solutions, the Hypervolume (HV) [74] indicator is one of the most widely used performance indicators, which is defined as follows.
-Definition 6 (Hypervolume [74])
+If all objectives are convex with $\lambda_{i}>0,\forall i$, it also serves as the Karush-Kuhn-Tucker (KKT) sufficient and necessary condition for Pareto optimality (Désidéri, 2012).
+When evaluating the performance of a set of obtained solutions, the Hypervolume (HV) (Zitzler and Thiele, 1998) indicator is one of the most widely used performance indicators, which is defined as follows.
+Definition 2.6 (Hypervolume (Zitzler and Thiele, 1998)).
 Given a solution set ${\mathbb{S}}=\{{\bm{q}}^{(1)},\dots,{\bm{q}}^{(N)}\}$ and a reference point $\bm{r}$}, the hypervolume of $\mathbb{S}$} is calculated by:
-$\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$ (2)
+(2) $\mathrm{HV}_{\bm{r}}({\mathbb{S}})=\Lambda({{\bm{p}}\;|\;\exists{\bm{q}}\in{\mathbb{S}}:{\bm{q}}\preceq{\bm{p}}\preceq{\bm{r}}}),$
 where $\Lambda(\cdot)$ denotes the Lebesgue measure of a set.
-HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are diverse and their objective values are close to the PF.
-2.2 Scalarization Methods
-Linear scalarization.
+HV provides a unary measure of both convergence and diversity of a solution set. A larger HV indicates that the obtained solutions are more diverse and closer the PF in the objective space.
+2.2. Scalarization Methods
+Linear Scalarization (LS).
 The most straightforward way to solve MOO problems is to reformulate them as single-objective optimization problems weighted by the given preference vector $\bm{\alpha}$} as follows:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$ (3)
-Linear scalarization is widely used in practice due to its simplicity.
+(3) $\min_{{\bm{\theta}}\in{\mathcal{K}}}g_{\bm{\alpha}}^{\mathrm{LS}}({\bm{\theta}}):=\sum_{i=1}^{m}\alpha_{i}f_{i}({\bm{\theta}}).$
+LS is widely used in practice due to its simplicity.
 However, this method suffers from a limitation: it can only identify solutions on the convex part of the PF.
@@
 Tchebycheff Scalarization.
-Another way to convert an MOP to a single one is to use the Tchebycheff scalarization function [75, 76], defined as:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}}f_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$ (4)
+Another way to convert a MOO problem to a single objective optimization problem is to use the Tchebycheff scalarization function (Bowman Jr, 1976; Steuer and Choo, 1983), defined as:
+(4) $\min_{{\bm{\theta}}\in{\mathcal{K}}}g_{\bm{\alpha}}^{\mathrm{Tche}}({\bm{\theta}}):=\max_{i\in[m]}\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*}),$
 where $z_{i}^{*}$ are reference values, usually set as the ideal value of the $i$-th objective.
@@
 The relationship between Tchebycheff scalarization and weak Pareto optimality can be formally described as follows:
-Theorem 1 ([77])
+Theorem 2.7 ((Choo and Atkins, 1983)).
 A solution $\bm{\theta}$$^{*}$ of the original MOO problem (1) is weakly Pareto optimal if and only if there exists a preference vector $\bm{\alpha}$} such that $\bm{\theta}$$^{*}$ is an optimal solution of problem (4).
@@
 However, Tchebycheff scalarization poses challenges for gradient-based deep learning due to the nonsmooth nature of the $\max(\cdot)$ operator.
-This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance [18].
-Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) remains non-differentiable.
-3 Finding a Single Pareto Optimal Solution
-In some scenarios such as multi-task learning (MTL) [3, 78, 66], it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set.
+This nonsmoothness leads to non-differentiability and a slow convergence rate of $\mathcal{O}(1/\epsilon^{2})$, where $\epsilon$ represents the error tolerance (Lin et al., 2024b).
+Even when all objectives $\{f_{i}\}_{i=1}^{m}$ are differentiable, problem (4) is still non-differentiable.
+3. Finding a Single Pareto Optimal Solution
+In some scenarios such as multi-task learning (MTL) (Zhang and Yang, 2022; Ruder, 2017; Crawshaw, 2020), it is sufficient to identify a single Pareto optimal solution without the need to explore the entire Pareto set.
 Here, all objectives are usually treated as equally important (i.e., the preference vector is ${\bm{\alpha}}=[\frac{1}{m},\dots,\frac{1}{m}]^{\top}$), and the goal is to produce a single model that effectively balances all objectives.
-The most straightforward method is the linear scalarization method (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) [3] in MTL.
-However, EW may cause some tasks to have unsatisfactory performance [79, 12].
+The most straightforward method is linear scalarization (i.e., problem (3)) with $\alpha_{i}=\frac{1}{m}$, which is known as Equal Weighting (EW) (Zhang and Yang, 2022) in MTL.
+However, EW may cause some tasks to have unsatisfactory performance (Standley et al., 2020; Lin et al., 2023).
 Therefore, to improve the performance, many methods have been proposed to dynamically tune the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training, which can in general be formulated as:
-$\min_{{\bm{\theta}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}}),$ (5)
+(5) $\min_{{\bm{\theta}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}}),$
 where $\lambda_{i}$ is the weight for the $i$-th objective $f_{i}({\bm{\theta}})$ and the uniform preference vector $\bm{\alpha}$} is omitted for simplicity.
 These methods can be categorized as loss balancing methods and gradient balancing approaches.
-Some representative loss/gradient balancing methods are illustrated in Figure 3 (adapted from [19, 11]).
-3.1 Loss Balancing Methods
-This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss [9], homeostatic uncertainty of loss [10], loss scale [11, 12], and validation loss [13, 14, 16, 15].
-Dynamic Weight Average (DWA) [9] estimates each objective weight as the ratio of the training losses in the last two iterations.
+Some representative loss/gradient balancing methods are illustrated in Figure 4 (adapted from (Liu et al., 2021b, a)).
+3.1. Loss Balancing Methods
+This type of approach dynamically computes or learns the objective weights $\{\lambda_{i}\}_{i=1}^{m}$ during training using some measures on loss such as the decrease speed of loss (Liu et al., 2019a), homeostatic uncertainty of loss (Kendall et al., 2018), loss scale (Liu et al., 2021a; Lin et al., 2023), and validation loss (Ye et al., 2021, 2024b, 2024a; Liu et al., 2022).
+Dynamic Weight Average (DWA) (Liu et al., 2019a) estimates each objective weight as the ratio of the training losses in the last two iterations.
 Formally, the objective weight in $k$-th iteration ($k>2$) is computed as follows:
-$\lambda_{i}^{(k)}=\frac{m\exp(\omega^{(k-1)}_{i}/\gamma)}{\sum_{j=1}^{m}\exp(\omega_{j}^{(k-1)}/\gamma)},\quad\omega^{(k-1)}_{j}=\frac{f^{(k-1)}_{j}}{f^{(k-2)}_{j}},$ (6)
+(6) $\lambda_{i}^{(k)}=\frac{m\exp(\omega^{(k-1)}_{i}/\gamma)}{\sum_{j=1}^{m}\exp(\omega_{j}^{(k-1)}/\gamma)},\quad\omega^{(k-1)}_{j}=\frac{f^{(k-1)}_{j}}{f^{(k-2)}_{j}}$,
 where $\gamma$ is the temperature and $f^{(k)}_{i}$ is the loss value of $i$-th objective at $k$-th iteration.
-Kendall et al. [10] propose Uncertainty Weighting (UW) that considers the homoscedastic uncertainty of each objective and optimizes objective weights together with the model parameter as follows:
-$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$ (7)
+Kendall et al.
+(Kendall et al., 2018) propose Uncertainty Weighting (UW) that considers the homoscedastic uncertainty of each objective and optimizes objective weights together with the model parameter as follows:
+(7) $\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(\frac{1}{2s_{i}^{2}}f_{i}({\bm{\theta}})+\log s_{i}\right),$
 where ${\bm{s}}=[s_{1},\dots,s_{m}]^{\top}$ is learnable and $\log s_{i}$’s are regularization terms.
-Impartial Multi-Task Learning (IMTL) [11] balances losses across tasks, abbreviated as IMTL-L(oss).
+Impartial Multi-Task Learning (IMTL) (Liu et al., 2021a) balances losses across tasks, abbreviated as IMTL-L(oss).
 Specifically, IMTL-L encourages all objectives to have a similar loss scale by transforming each objective $f_{i}({\bm{\theta}})$ as $e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}$.
-Similar to UW [10], IMTL-L simultaneously learns the transformation parameter ${\bm{s}}$ and the model parameter ${\bm{\theta}}$ as follows:
-$\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}\right).$ (8)
-Similar to IMTL-L [11], Lin et al.
-[12] also aim to make all objective losses have a similar scale.
+Similar to UW (Kendall et al., 2018), IMTL-L simultaneously learns the transformation parameter $\bm{s}$} and the model parameter $\bm{\theta}$} as follows:
+(8) $\min_{{\bm{\theta}},{\bm{s}}}\sum_{i=1}^{m}\left(e^{s_{i}}f_{i}({\bm{\theta}})-s_{i}\right).$
+Similar to IMTL-L (Liu et al., 2021a), Lin et al.
+(Lin et al., 2023) also aim to make all objective losses have a similar scale.
 They achieve it by performing a logarithm transformation on each objective (i.e., $\log f_{i}({\bm{\theta}})$).
@@
 Ye et al.
-[13, 14] propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights ${\bm{\lambda}}$ based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
-$\min_{{\bm{\lambda}}}~{}\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$ (9a)
-$\ ~{}\mathrm{s.t.}\ ~{}{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$ (9b)
+(Ye et al., 2021, 2024b) propose Multi-Objective Meta Learning (MOML), which adaptively tunes the objective weights $\bm{\lambda}$} based on the validation performance, reformulating problem (5) as a multi-objective bi-level optimization problem:
+(9a) $\min_{{\bm{\lambda}}}~\big{[}f_{1}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{1}),\dots,f_{m}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{m})\big{]}^{\top}$
+(9b) $\ ~\mathrm{s.t.}\ ~{\bm{\theta}}^{*}({\bm{\lambda}})=\operatorname*{arg\,min}_{{{\bm{\theta}}}}\sum_{i=1}^{m}\lambda_{i}f_{i}({\bm{\theta}};\mathcal{D}^{\text{tr}}_{i}),$
 where $\mathcal{D}^{\text{tr}}_{i}$ and $\mathcal{D}^{\text{val}}_{i}$ are the training and validation datasets for the $i$-th objective, respectively.
@@
 Hence, the time and memory costs of MOML grow significantly fast with respect to the dimension of $\bm{\theta}$} and the number of LL iterations $T$.
-Auto-$\lambda$ [15] modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$.
+Auto-$\lambda$ (Liu et al., 2022) modifies problem (9) by replacing the multi-objective UL subproblem (9a) with a single-objective problem: $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}f_{i}({\bm{\theta}}^{*}({\bm{\lambda}});\mathcal{D}^{\text{val}}_{i})$.
 Besides, Auto-$\lambda$ approximates the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$ via the finite difference approach.
-Thus, it is more efficient than MOML [13, 14].
-Recently, FORUM [16] is proposed to improve the efficiency of MOML [13, 14].
-FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach [80] and then proposes a multi-gradient aggregation method to solve it.
+Thus, it is more efficient than MOML (Ye et al., 2021, 2024b).
+FORUM (Ye et al., 2024a) is proposed to improve the efficiency of MOML (Ye et al., 2021, 2024b).
+FORUM first reformulates problem (9) as a constrained MOO problem via the value-function approach (Liu et al., 2021c) and then proposes a multi-gradient aggregation method to solve it.
 Compared with MOML, FORUM is a first-order method and does not need to compute the hypergradient $\nabla_{{\bm{\lambda}}}{\bm{\theta}}^{*}({\bm{\lambda}})$.
 Thus, it is significantly more efficient than MOML.
-Random Weighting (RW) [17] randomly samples objective weights from a distribution (such as the standard normal distribution) and normalizes them into the simplex in each iteration.
-Compared to EW, RW has a higher probability of escaping local minima than EW due to the randomness from objective weights, resulting in better generalization performance.
-Different from the above methods which are based on linear scalarization, Lin et al.
-[18] propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
-$\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})/\mu\right),$ (10)
+Random Weighting (RW) (Lin et al., 2022c) randomly samples objective weights from a distribution (such as the standard normal distribution) and normalizes them into the simplex in each iteration.
+RW has a higher probability of escaping local minima than EW due to the randomness from objective weights, resulting in better generalization performance.
+Different from the above methods based on linear scalarization, Lin et al.
+(Lin et al., 2024b) propose Smooth Tchebycheff scalarization (STCH), which transforms the Tchebycheff scalarization into a smooth function:
+(10) $\min_{{\bm{\theta}}}f_{\bm{\alpha}}^{\mathrm{STCH}}({\bm{\theta}},\mu):=\mu\log\sum_{i=1}^{m}\exp\left(\frac{\alpha_{i}(f_{i}({\bm{\theta}})-z_{i}^{*})}{\mu}\right),$
 where $\mu>0$ is a smoothing parameter, and $\bm{z}$$^{*}$ is a reference point as in Tchebycheff scalarization (problem (4)).
@@
 It also retains the advantages of the original Tchebycheff scalarization: (1) convexity when all objective functions are convex and (2) Pareto optimality by appropriately choosing $\mu$.
-3.2 Gradient Balancing Methods
+3.2. Gradient Balancing Methods
 Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ be the gradient of the $i$-th objective at the $k$-th iteration.
@@
 To simplify notations, we omit the superscript in this section.
-Compared with loss balancing methods in Section 3.1, the gradient balancing approach usually achieves better performance.
-Figure 3: Illustration of the update direction ${\bm{d}}$ of several representative methods: EW, DWA [9] from loss balancing methods, MGDA [6] from gradient weighting methods, and PCGrad [33] from gradient manipulation methods.
+Compared with loss balancing methods in Section 3.1, gradient balancing approaches usually achieve better performance.
+Figure 4.
+Illustration of the update direction $\bm{d}$} of several representative methods: EW, DWA (Liu et al., 2019a) from loss balancing methods, MGDA (Sener and Koltun, 2018) from gradient weighting methods, and PCGrad (Yu et al., 2020a) from gradient manipulation methods.
 They are illustrated in a two-objective learning problem (two specific gradients are labeled as ${\bm{g}}_{1}$ and ${\bm{g}}_{2}$).
-3.2.1 Gradient Weighting Methods
+3.2.1. Gradient Weighting Methods
 In most gradient weighting methods, $\bm{d}$} is computed by:
-${\bm{d}}={\bm{G}}{\bm{\lambda}},$ (11)
+(11) ${\bm{d}}={\bm{G}}{\bm{\lambda}},$
 where ${\bm{G}}=[{\bm{g}}_{1},\dots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=[\lambda_{1},\dots,\lambda_{m}]^{\top}$ is the gradient weights.
-Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [73], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses.
+Sener and Koltun (Sener and Koltun, 2018) apply Multiple-Gradient Descent Algorithm (MGDA) (Désidéri, 2012), which aims to find a direction $\bm{d}$} at each iteration to maximize the minimal decrease across the losses.
 This is formulated as:
-$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
+(12) $\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$
 where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where $\bm{\lambda}$} is computed as:
-${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$ (13)
-where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun [6] use the Frank-Wolfe algorithm [81] to solve problem (13).
-Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant.
+(13) ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2},$
+where $\Delta_{m-1}$ denotes the simplex. Sener and Koltun (Sener and Koltun, 2018) use the Frank-Wolfe algorithm (Jaggi, 2013) to solve problem (13).
+Conflict-Averse Gradient descent (CAGrad) (Liu et al., 2021b) improves MGDA (Sener and Koltun, 2018) by constraining the aggregated gradient $\bm{d}$} to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant.
 Specifically, $\bm{d}$} is computed by solving the following problem:
-$\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~{}~{}\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$ (14)
+(14) $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}_{i}^{\top}{\bm{d}},\quad\mathrm{s.t.}~~\left\|{\bm{d}}-{\bm{g}}_{0}\right\|\leq c\left\|{\bm{g}}_{0}\right\|.$
 Problem (14) is equivalent to first computing $\bm{\lambda}$} by solving the following problem:
-${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$ (15)
+(15) ${\bm{\lambda}}=\operatorname*{arg\,min}_{{\bm{\lambda}}\in\Delta_{m-1}}{{\bm{g}}^{\top}_{\bm{\lambda}}}{\bm{g}}_{0}+\left\|{\bm{g}}_{0}\right\|\left\|{\bm{g}}_{{\bm{\lambda}}}\right\|,$
 where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate the update direction ${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}$.
-Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
-MGDA [6] and its variant CAGrad [19] can converge to a Pareto stationary point in the deterministic setting with full-gradient computations.
-However, in deep learning, where stochastic gradients are used, many works [20, 21, 23, 22, 24, 25, 26] explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
-IMTL-G(rad) [11] aims to find ${\bm{d}}$ that has equal projections on all objective gradients. Thus, we have:
-${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\quad 2\leq i\leq m,$ (16)
+Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA (Sener and Koltun, 2018) when $c\rightarrow+\infty$.
+MGDA (Sener and Koltun, 2018) and its variant CAGrad (Liu et al., 2021b) can converge to a Pareto stationary point in the deterministic setting with full-gradient computations.
+However, in deep learning, where stochastic gradients are used, many works (Liu and Vicente, 2021; Zhou et al., 2022; Xu et al., 2025b; Fernando et al., 2023; Chen et al., 2023a; Xiao et al., 2023; Zhang et al., 2025) explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
+IMTL-G(rad) (Liu et al., 2021a) aims to find $\bm{d}$} that has equal projections on all objective gradients. Thus, we have:
+(16) ${\bm{u}}^{\top}_{1}{\bm{d}}={\bm{u}}_{i}^{\top}{\bm{d}},\quad 2\leq i\leq m,$
 where ${\bm{u}}_{i}=\nicefrac{{{\bm{g}}_{i}}}{{\left\|{\bm{g}}_{i}\right\|}}$. If constraining $\sum_{i=1}^{m}\lambda_{i}=1$, problem (16) has a closed-form solution of $\bm{\lambda}$}:
-${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (17)
+(17) ${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$
 where ${\bm{\lambda}}_{(2,\dots,m)}=[\lambda_{2},\dots,\lambda_{m}]^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\dots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\dots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
-Nash-MTL [27] formulates problem (11) as a bargaining game [82, 83], where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$.
+Nash-MTL (Navon et al., 2022) formulates problem (11) as a bargaining game (Nash, 1953; Thomson, 1994), where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$.
 Hence, $\bm{d}$} is computed by solving the following problem:
-$\max_{{\bm{d}}}\sum_{i=1}^{m}\log\left({{\bm{g}}_{i}}^{\top}{\bm{d}}\right).$ (18)
+(18) $\max_{{\bm{d}}}\sum_{i=1}^{m}\log\left({{\bm{g}}_{i}}^{\top}{\bm{d}}\right).$
 The solution of problem (18) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where $\bm{\lambda}$} is obtained by solving the following problem:
-${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1},$ (19)
+(19) ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1},$
 which can be approximately solved using a sequence of convex optimization problems.
-FairGrad [28] extends Nash-MTL [27] by computing ${\bm{\lambda}}$ through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant.
-Unlike Nash-MTL [27], FairGrad treats the problem as a constrained nonlinear least squares problem.
-Similarly, UPGrad [29] calculates ${\bm{\lambda}}$ by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}{\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}$, a quadratic programming (QP) problem.
-Aligned-MTL [30] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$.
+FairGrad (Ban and Ji, 2024) extends Nash-MTL (Navon et al., 2022) by computing $\bm{\lambda}$} through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant.
+Unlike Nash-MTL (Navon et al., 2022), FairGrad treats the problem as a constrained nonlinear least squares problem.
+Similarly, UPGrad (Quinton and Rey, 2024) calculates $\bm{\lambda}$} by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}({\bm{G}}^{\top}{\bm{G}}){\bm{\lambda}}$, a quadratic programming (QP) problem.
+Aligned-MTL (Senushkin et al., 2023) considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of $\bm{G}$}.
 Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed.
@@
 Finally, $\bm{\lambda}$} is computed by:
-${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$ (20)
+(20) ${\bm{\lambda}}=\sqrt{\sigma_{R}}{\bm{V}}{{\bm{\Sigma}}}^{-1}{{\bm{V}}}^{\top},$
 where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\dots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
-Gradient Normalization (GradNorm) [31] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have similar magnitudes.
+Gradient Normalization (GradNorm) (Chen et al., 2018) aims to learn $\bm{\lambda}$} so that the scaled gradients have similar magnitudes.
 It updates $\bm{\lambda}$} by solving the following problem via one-step gradient descent:
-$\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$ (21)
+(21) $\min_{{\bm{\lambda}}}\sum_{i=1}^{m}\left(\left\|\lambda_{i}{\bm{g}}_{i}\right\|-c\times r_{i}^{\gamma}\right),$
 where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is treated as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ is the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
-Dual-Balancing Multi-Task Learning (DB-MTL) [12] also normalizes all objective gradients to the same magnitude.
+Dual-Balancing Multi-Task Learning (DB-MTL) (Lin et al., 2023) also normalizes all objective gradients to the same magnitude.
 $\bm{\lambda}$} is computed as $\lambda_{i}=\nicefrac{{\gamma}}{{\left\|{\bm{g}}_{i}\right\|}}$, where $\gamma=\max_{i\in[m]}\left\|{\bm{g}}_{i}\right\|$ is a scaling factor controlling the update magnitude.
-Thus, compared with GradNorm [31], this method guarantees all objective gradients have the same norm in each iteration.
+Thus, compared with GradNorm (Chen et al., 2018), this method guarantees all objective gradients have the same norm in each iteration.
 Moreover, they observe that the choice of the update magnitude $\gamma$ significantly affects performance.
-3.2.2 Gradient Manipulation Methods
-To address the conflicting gradient problem, gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
-Projecting Conflicting Gradients (PCGrad) [33] considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$.
+3.2.2. Gradient Manipulation Methods
+To address gradient conflicting , gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
+Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020a) considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$.
 Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients.
@@
 Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as:
-$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$ (22)
+(22) $\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}}{\left\|{\bm{g}}_{j}\right\|^{2}}{\bm{g}}_{j}.$
 This reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients.
-While PCGrad [33] corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) [34] extends it to a more general and adaptive formulation.
+While PCGrad (Yu et al., 2020a) corrects the gradient if and only if two objectives have a negative gradient similarity, Gradient Vaccine (GradVac) (Wang et al., 2021b) extends it to a more general and adaptive formulation.
 In each iteration, if $\phi_{ij}<\bar{\phi}_{ij}$, GradVac updates the corrected gradient $\hat{{\bm{g}}}_{i}$ as:
-$\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\left\|\hat{{\bm{g}}}_{i}\right\|\left(\bar{\phi}_{ij}\sqrt{1-\phi_{ij}^{2}}-\phi_{ij}\sqrt{1-\bar{\phi}_{ij}^{2}}\right)}{\left\|{\bm{g}}_{j}\right\|\sqrt{1-\bar{\phi}_{ij}^{2}}}{\bm{g}}_{j},$ (23)
+(23) $\hat{{\bm{g}}}_{i}=\hat{{\bm{g}}}_{i}-\frac{\left\|\hat{{\bm{g}}}_{i}\right\|\left(\bar{\phi}_{ij}\sqrt{1-\phi_{ij}^{2}}-\phi_{ij}\sqrt{1-\bar{\phi}_{ij}^{2}}\right)}{\left\|{\bm{g}}_{j}\right\|\sqrt{1-\bar{\phi}_{ij}^{2}}}{\bm{g}}_{j},$
 where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$.
-$\bar{\phi}_{ij}$ is initialized to 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter.
+$\bar{\phi}_{ij}$ is initialized to 0 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter.
 Note that GradVac simplifies to PCGrad when $\bar{\phi}_{ij}=0$.
-Gradient Sign Dropout (GradDrop) [35] considers that conflicts arise from differences in the sign of gradient values.
+Gradient Sign Dropout (GradDrop) (Chen et al., 2020) considers that conflicts arise from differences in the sign of gradient values.
 Thus, a probabilistic masking procedure is proposed to preserve gradients with consistent signs during each update.
-3.2.3 Practical Speedup Strategy
+3.2.3. Practical Speedup Strategy
 Gradient balancing methods suffer from high computational and storage costs.
@@
 This can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters.
-Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], Nash-MTL [27], and FairGrad [28]) need to solve an optimization problem to obtain the objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs.
+Moreover, many gradient balancing methods (such as MGDA (Sener and Koltun, 2018), CAGrad (Liu et al., 2021b), Nash-MTL (Navon et al., 2022), and FairGrad (Ban and Ji, 2024)) need to solve an optimization problem to obtain the objective weight $\bm{\lambda}$} in each iteration, which also increases the computational and memory costs.
 Hence, several strategies are proposed to alleviate this problem.
-Sener and Koltun [6] use feature-level gradients (i.e., gradients w.r.t.
+Sener and Koltun (Sener and Koltun, 2018) use feature-level gradients (i.e., gradients w.r.t.
 the representations from the last shared layer) to replace the parameter-level gradients (i.e., gradients w.r.t.
@@
 Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs.
-This strategy is also adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [30].
+This strategy is also adopted by some gradient balancing methods, such as IMTL-G (Liu et al., 2021a) and Aligned-MTL (Senushkin et al., 2023).
 Liu et al.
-[19] randomly select a subset of objectives to calculate the update direction in each iteration.
+(Liu et al., 2021b) randomly select a subset of objectives to calculate the update direction in each iteration.
 Navon et al.
-[27] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iterations instead of updating in each iteration.
+(Navon et al., 2022) propose to update the objective weight $\bm{\lambda}$} every $\tau$ iterations instead of updating in each iteration.
 Although this strategy speeds up training, they observe that it may cause a noticeable drop in performance.
 Liu et al.
-[32] consider optimizing the logarithm of the MGDA objective [6] and propose a speedup strategy.
+(Liu et al., 2023) consider optimizing the logarithm of the MGDA objective (Sener and Koltun, 2018) and propose a speedup strategy.
 Specifically, when solving problem (13) via one-step gradient descent, $\bm{\lambda}$} is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$.
 Note that
-$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$
-$\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right]^{\top},$ (24)
+(24) $\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right$$^{\top}$,
 where $f^{(k)}_{i}$ is the loss value of the $i$-th objective in the $k$-th iteration.
@@
 Moreover, it is only applicable to MGDA-based methods.
-3.3 Discussions
+3.3. Discussions
 Loss balancing and gradient balancing methods effectively mitigate conflicts among different objectives and achieve a Pareto-optimal solution, which is important for MOO scenarios such as multi-task learning.
 However, they are limited to finding a single Pareto-optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning.
-In Sections 4 and 5, we introduce methods to achieve more diverse and controllable solutions.
-The computational cost of loss balancing methods (except MOML [13, 14], Auto-$\lambda$ [15], and FORUM [16]) is almost the same as that of EW since they only need one backpropagation process in each training iteration.
+In Sections 4 and 5, we introduce methods to achieve more diverse and targeted solutions.
+The computational cost of loss balancing methods (except MOML (Ye et al., 2021, 2024b), Auto-$\lambda$ (Liu et al., 2022), and FORUM (Ye et al., 2024a)) is almost the same as that of EW since they only need one backpropagation process in each training iteration.
 However, gradient balancing methods are much more computationally expensive due to the need of $m$ backpropagation processes to obtain each objective gradient, where $m$ is the number of objectives.
-Although some speedup strategies are proposed to alleviate this problem, they may cause a drop in performance, as detailed in Section 3.2.3.
-Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, as detailed in Section 6.1.
-Figure 4: Distributions of solutions obtained by using representative methods: PTML [36] and EPO [37, 38] are preference vector-based methods, while GradHV [47] and MOO-SVGD [48] are methods not using preference vectors.
+While various speedup strategies can address this problem, they often come at the cost of performance, as detailed in Section 3.2.3.
+Compared to loss balancing methods, gradient balancing methods generally exhibit better convergence properties such as guaranteeing to converge to a Pareto stationary point, detailed in Section 6.1.
+Figure 5.
+Distributions of solutions obtained by using representative methods: PMTL (Lin et al., 2019b) and EPO (Mahapatra and Rajan, 2020, 2021) are preference vector-based methods, while GradHV (Wang et al., 2017) and MOO-SVGD (Liu et al., 2021d) are methods not using preference vectors.
 The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
-4 Finding a Finite Set of Solutions
+As can be seen, MGDA (Sener and Koltun, 2018) demonstrates poor diversity due to its lack of preference incorporation.
+While the preference vector-based method PTML (Lin et al., 2019b) achieves better alignment with the preference vector, it still falls short of exact alignment.
+In contrast, EPO (Mahapatra and Rajan, 2020, 2021) successfully achieves exact alignment.
+GradHV (Wang et al., 2017) and MOO-SVGD (Liu et al., 2021d) can also generate diverse solutions, despite they do not use a preference vector.
+4. Finding a Finite Set of Solutions
 In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs.
@@
 We first discuss methods based on preference vectors (Section 4.1), followed by approaches that do not require the use of preference vectors (Section 4.2).
-Solutions obtained by some representative methods on LSMOP1 [84] are shown in Figure 4, with additional details and more results provided in Appendix B.
-4.1 Methods Based on Preference Vectors
+Figure 5 shows solutions obtained by some representative methods on LSMOP1 (Cheng et al., 2016) after 5000 iterations of each algorithm.
+4.1. Methods Based on Preference Vectors
 Preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\dots,{\bm{\alpha}}^{(n)}\}$.
@@
 By solving $n$ subproblems, these methods obtain a set of $n$ solutions.
-Pareto Multi-Task Learning (PMTL) [36], inspired by the idea of decomposition-based MOO algorithms [85], incorporates preference vectors as constraints on the objectives.
+Pareto Multi-Task Learning (PMTL) (Lin et al., 2019b), inspired by the idea of decomposition-based MOO algorithms (Zhang and Li, 2007), incorporates preference vectors as constraints on the objectives.
 For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to the other preference vectors:
-$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$
-$~{}~{}\mathrm{s.t.}~{}~{}r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,\quad j\in[n].$ (25)
+(25) $\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},~~\mathrm{s.t.}~~r^{(j)}({\bm{\theta}}):=({\bm{\alpha}}^{(j)}-{\bm{\alpha}}^{(i)})^{\top}{\bm{f}}({\bm{\theta}})\leq 0,~~j\in[n].$
 Let $\bm{R}$} be the matrix containing gradients of the active constraints, i.e, ${\bm{R}}=\big{[}\nabla r^{(j)}({\bm{\theta}})\big{]}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ containing indices of the active constraints.
-Using a method akin to MGDA [6], PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$.
+Using a method akin to MGDA (Sener and Koltun, 2018), PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$.
 The descent direction is:
-${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$ (26)
-where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
-${\bm{\lambda}},{\bm{\beta}}=$ $\operatorname*{arg\,min}_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},$
-[…]
-A limitation of PMTL is its weak constraint that only requires the objective vector to lie within sectors, which lacks precise control over the position of Pareto solutions.
-Exact Pareto Optimal search (EPO) [37, 38] is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors.
+(26) ${\bm{d}}=-{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}},$
+where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the Jacobian matrix.
+$\bm{\lambda}$} and $\bm{\beta}$} are coefficients obtained by solving the following quadratic programming problem:
+(27) $\min_{{\bm{\lambda}},{\bm{\beta}}}\left\|{\bm{G}}{\bm{\lambda}}+{\bm{R}}{\bm{\beta}}\right\|^{2},~~\text{s.t.}~~{\bm{\lambda}}^{\top}\bm{1}+{\bm{\beta}}^{\top}\bm{1}=1,\;\lambda_{i}\geq 0,\;\beta_{j}\geq 0.$
+A limitation of PMTL is it can only constrain objective vectors within sectors, which lacks precise control over the position of Pareto solutions.
+Exact Pareto Optimal search (EPO) (Mahapatra and Rajan, 2020, 2021) is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors.
 EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with $\bm{\alpha}$}, achieving precise alignment and Pareto optimality.
@@
 Based on these sets, it determines $\bm{\lambda}$} for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
-${\bm{\lambda}}=$ $\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}~{}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
-$\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right.$ (30)
+(30) $\max_{{\bm{\lambda}}\in\Delta_{m-1}}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),~~\mathrm{s.t.}\left\{\begin{array}[]{ll}{\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq{\bm{a}}^{\top}{\bm{c}}_{j}\mathbbm{1}_{\mathbb{J}},\quad\forall j\in\overline{{\mathbb{J}}}\setminus{\mathbb{J}}^{*},\\ {\bm{\lambda}}^{\top}{\bm{c}}_{j}\geq 0,\quad\forall j\in{\mathbb{J}}^{*}.\end{array}\right$.
 This updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector.
-However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (30)).
-To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [39] considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem.
+However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem.
+To address these drawbacks, Weighted Chebyshev MGDA (WC-MGDA) (Momma et al., 2022) considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem.
 Optimizing Tchebycheff function addresses the previously mentioned drawbacks of EPO.
 However, its convergence rate is relatively slow.
-Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the Preference-based MGDA (PMGDA) [40] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient.
-FERERO [41] captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint.
+Two methods for improving the Tchebycheff scalarization include smooth Tchebycheff scalarization, which replaces the non-smooth max operation with a smooth approximation, and Preference-based MGDA (PMGDA), which only requires the update direction to have a negative inner product with the exact constraint gradient.
+FERERO (Chen et al., 2024a) captures preference information within the MGDA framework, addressing both objective constraints with constants and the exactness constraint.
 It achieves fast convergence rates of $\mathcal{O}(\epsilon^{-1})$ for deterministic gradients and $\mathcal{O}(\epsilon^{-2})$ for stochastic gradients, where $\epsilon$ is the error tolerance.
@@
 However, in real-world applications where the shape of the Pareto front is unknown, a predefined set of preference vectors may not always result in well-distributed solutions.
-To overcome this limitation, GMOOAR [42] formulates the problem as a bi-level optimization problem.
+To overcome this limitation, GMOOAR (Chen and Kwok, 2022) formulates the problem as a bi-level optimization problem.
 In the upper level, preference vectors are optimized to maximize either the hypervolume or uniformity of the solutions.
@@
 This approach enables the algorithm to dynamically adjust preference vectors based on the given optimization problem, ensuring the desired solution distribution.
-UMOD [43] introduces an approach that aims to maximize the minimum distance between solutions:
-$\max\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (31)
+UMOD (Zhang et al., 2024a) introduces an approach that aims to maximize the minimum distance between solutions:
+(31) $\max_{{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS}}\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),$
 where $\rho(\cdot,\cdot)$ denotes the Euclidean distance.
 UMOD achieves desirable solution distributions with theoretical guarantees: (1) For bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) As the number of solutions increases, the objective vectors asymptotically distribute on the PF.
-4.2 Methods without Using Preference Vectors
-Another line of research, directly optimize for a diverse set of Pareto-optimal solutions, does not rely on a given set of preference vectors.
-As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO [86].
-In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) [47, 44, 45, 46] have been developed.
+4.2. Methods without Using Preference Vectors
+Unlike methods that rely on preference vectors, another approach directly optimizes for a diverse set of Pareto-optimal solutions.
+As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO (Shang et al., 2020).
+In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) (Wang et al., 2017; Deist et al., 2021, 2020; Emmerich et al., 2007) have been developed.
 Let $\{{\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors.
 These algorithm first calculate the hypervolume gradient of the $i$-th solution ${\bm{\theta}}^{(i)}$ as follows:
-${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (32)
+(32) ${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$
 Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}+\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate.
@@
 The main challenge lies in estimating the first term $a_{i}$.
-Following Emmerich et al.
-[46], index sets are categorized based on differentiability: ${\mathbb{Z}}$ includes subvectors with zero partial gradients, ${\mathbb{U}}$ contains subvectors with undefined or indeterminate gradients, and ${\mathbb{P}}$ has subvectors with positive gradients.
+As proposed by Emmerich et al.
+(Emmerich et al., 2007), index sets are categorized based on differentiability: $\mathbb{Z}$} includes subvectors with zero partial gradients, $\mathbb{U}$} contains subvectors with undefined or indeterminate gradients, and $\mathbb{P}$} has subvectors with positive gradients.
 This classification leads to many zero-gradient terms, improving efficiency.
-Then, non-zero terms are computed using a fast dimension-sweeping algorithm [87, 88].
+Then, non-zero terms are computed using a fast dimension-sweeping algorithm (Beume et al., 2009; Guerreiro et al., 2012).
 This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions.
-To further enhance convergence, the Newton-hypervolume method [89] incorporates second-order information for faster updates.
+To further enhance convergence, the Newton-hypervolume method (Sosa Hernández et al., 2020) incorporates second-order information for faster updates.
 Deist et al.
-[44] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
-Inspired by Stein Variational Gradient Descent (SVGD) [90], MOO-SVGD [48] proposes a different approach to maintain diversity while enhancing convergence.
+(Deist et al., 2021) apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
+Inspired by Stein Variational Gradient Descent (SVGD) (Liu and Wang, 2016), MOO-SVGD (Liu et al., 2021d) proposes a different approach to maintain diversity while enhancing convergence.
 The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
-${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$
-$\qquad\qquad+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$ (33)
-where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)).
+${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$
+where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA (Sener and Koltun, 2018) (i.e., problem (13)).
 $\kappa(\cdot,\cdot)$ represents a kernel function, often chosen as the radial basis function (RBF) kernel.
-The first term in (33) guides the solution towards the PS, while the second term pushes the solutions away from each other to enhance the diversity.
-The methods discussed above focus on scenarios where the number of solutions $n$ exceeds the number of objectives $m$.
-Recent works [49, 51, 91, 50] explore the case with more objectives than solutions (i.e., $m>n$).
+The first term in Equation (4.2) drives the particles toward the target distribution, while the second term acts as a repulsive force that promotes diversity.
+While the methods discussed above address scenarios with more solutions than objectives ($n>m$), recent work has begun to explore the converse case, where objectives outnumber solutions ($m>n$) (Lin et al., 2025c; Li et al., 2025; Liu et al., 2024b; Ding et al., 2024).
 These approaches assign each objective to a solution ${\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}$ that achieves the smallest objective value for that particular objective.
-They then minimize either the maximum or the sum of these smallest objective values.
-Few for Many (F4M) [49, 91] uses the maximum aggregation, formulated as:
-$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$ (34)
-Sum of Minimum (SoM) [50] uses the sum aggregation, formulated as:
-$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (35)
-MosT [51] frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
-4.3 Discussions
-This section reviews two primary approaches for obtaining multiple Pareto optimal solutions.
-Preference vector-based methods decompose an MOO problem into multiple independent subproblems based on predefined preference vectors.
+The goal is then to minimize an aggregation of these smallest objective values, typically either their sum or their maximum.
+The Few for Many (F4M) framework (Lin et al., 2025c; Liu et al., 2024b), for example, employs a minimax strategy by minimizing the maximum value, which is formulated as:
+(33) $\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$
+Different from F4M, Sum of Minimum (SoM) (Ding et al., 2024) uses the sum aggregation, formulated as:
+(34) $\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$
+MosT (Li et al., 2025) frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
+4.3. Discussions
+This section discusses two primary approaches for obtaining multiple Pareto optimal solutions.
+Preference vector-based methods decompose a MOP into multiple independent subproblems based on predefined preference vectors.
 These subproblems can be solved independently, which enhances memory efficiency and enables parallel computation.
@@
 However, this approach typically requires higher memory resources and incurs greater computational costs, making it less efficient in resource-constrained scenarios.
-5 Finding an Infinite Set of Solutions
+5. Finding an Infinite Set of Solutions
 While a finite set of solutions can only provide a discrete approximation of the Pareto front, many applications require the ability to obtain solutions corresponding to any user preference, effectively representing the entire Pareto set.
@@
 Ma et al.
-[92] initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions.
+(Ma et al., 2020) initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions.
 However, this method has several drawbacks: (1) The approximation error grows when the solutions are widely spaced, (2) First-order approximations perform poorly in high-dimensional objective spaces, and (3) The approach requires solving a linear programming problem for updates, which can be computationally challenging.
@@
 Section 5.2 discusses training strategies for these network structures.
-5.1 Network Structure
-(a) Methods based on the hypernetwork.
-(b) Methods based on preference-conditioned network.
-(c) Methods based on model combination.
-Figure 5: Illustration of different structures to learn an infinite number of solutions.
-5.1.1 Hypernetwork
-Hypernetwork is a neural network that generates the parameters for another target network [93].
-This idea has been leveraged by the Pareto Hypernetwork (PHN) [52] and Controllable Pareto Multi-Task Learning (CPMTL) [53] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters.
-The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI [54] and SUHNPF [55].
+5.1. Network Structure
+5.1.1. Hypernetwork
+Hypernetwork is a neural network that generates the parameters for another target network (Ha et al., 2016).
+This idea has been leveraged by the Pareto Hypernetwork (PHN) (Navon et al., 2021) and Controllable Pareto Multi-Task Learning (CPMTL) (Lin et al., 2020) to learn the entire Pareto set, where the input is the user preference $\bm{\alpha}$} and the output is the target network’s parameters.
+The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI (Hoang et al., 2023) and SUHNPF (Gupta et al., 2022).
 Recently, Tuan et al.
-[56] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
+(Tuan et al., 2024) propose using a transformer architecture as the hypernetwork, demonstrating superior performance compared to MLP-based hypernetwork models.
 A primary limitation of hypernetwork-based algorithms is their sizes.
 Since the output dimension matches the number of parameters in the target network, hypernetworks are often much larger than the base networks, limiting their applicability to large models.
-Some methods address this limitation by employing hypernetworks with chunking [52, 53].
+Some methods address this limitation by employing hypernetworks with chunking (Navon et al., 2021; Lin et al., 2020).
 Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability.
 This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
-5.1.2 Preference-Conditioned Network
+5.1.2. Preference-Conditioned Network
 Instead of using a hypernetwork to generate weights, the original model can be directly modified to incorporate preferences.
-COSMOS [57] suggests adding preferences as an additional input to the model by combining user preference ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
+COSMOS (Ruchte and Grabocka, 2021) suggests adding preferences as an additional input to the model by combining user preference $\bm{\alpha}$} with input data $\bm{x}$}, thereby increasing the input dimension of the original model.
 However, such input-based conditioning has limited ability to produce diverse solutions.
-Some studies [58, 42, 59] employ the Feature-wise Linear Modulation (FiLM) layer [94] to condition the network.
+Some studies (Dosovitskiy and Djolonga, 2020; Chen and Kwok, 2022; Raychaudhuri et al., 2022) employ the Feature-wise Linear Modulation (FiLM) layer (Perez et al., 2018) to condition the network.
 The FiLM layer works by applying an affine transformation to the feature maps.
@@
 Raychaudhuri et al.
-[59] also propose leveraging another network to predict the network architecture.
+(Raychaudhuri et al., 2022) also propose leveraging another network to predict the network architecture.
 It dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
-5.1.3 Model Combination
+(a) Methods based on the hypernetwork.
+(b) Methods based on preference-conditioned network.
+(c) Methods based on model combination.
+Figure 6. Illustration of different structures to learn an infinite number of solutions.
+5.1.3. Model Combination
 Methods based on model combination construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity.
-PaMaL [60] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
-${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$ (36)
+PaMaL (Dimitriadis et al., 2023) achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
+(35) ${\bm{\theta}}({\bm{\alpha}})=\sum_{i=1}^{m}\alpha_{i}{\bm{\theta}}_{i}.$
 Although this incorporates more parameters than methods based on preference-conditioned networks, PaMaL enhances diversity and provides users with a broader range of choices.
 Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives.
-To mitigate this issue, LORPMAN [61] proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$.
+To mitigate this issue, LORPMAN (Chen and Kwok, 2024a) proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$.
 Given user preference $\bm{\alpha}$}, the composite model is given by:
-${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$ (37)
+(36) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s\sum_{i=1}^{m}\alpha_{i}{\bm{B}}_{i}{\bm{A}}_{i},$
 where $s$ is a scaling factor that adjusts the impact of the low-rank components.
 This approach efficiently scales Pareto set learning to manage up to $40$ objectives, achieving superior performance.
-A similar approach is also mentioned in [65].
-To further enhance parameter efficiency, Chen and Kwok [62] introduce a preference-aware tensor multiplication:
-${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$ (38)
-where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and ×$\times_{u}$ denotes mode-$u$ tensor multiplication [95].
-Zhong et al.[63] address scalability challenges for large-scale models, such as language models, using SVD:
-${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$ (39)
-where ${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices.
-${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, where $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars.
+A similar approach is also mentioned in (Dimitriadis et al., 2025).
+To further enhance parameter efficiency, Chen and Kwok (Chen and Kwok, 2024b) introduce a preference-aware tensor multiplication:
+(37) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+s{\bm{C}}\times_{1}{\bm{A}}\times_{2}{\bm{B}}\times_{3}{\bm{\alpha}},$
+where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and $\times_{u}$ denotes mode-$u$ tensor multiplication (Kolda and Bader, 2009).
+Zhong et al.
+(Zhong et al., 2024) address scalability challenges for large-scale models, such as large language models, using SVD-LoRA:
+(38) ${\bm{\theta}}({\bm{\alpha}})={\bm{\theta}}_{0}+{\bm{U}}{\bm{\Sigma}}{\bm{V}},$
+where $\bm{\Sigma}$} is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars.
+${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices.
 Instead of training new models, Tang et al.
-[64] propose to utilize a mixture of experts [96] to combine pre-existing models in order to achieve the Pareto set.
+(Tang et al., 2024) propose to utilize a mixture of experts (Cai et al., 2024) to combine pre-existing models in order to achieve the Pareto set.
 Given $n$ models, ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}$, each trained on different datasets, they are first integrated to a unified base model ${\bm{\theta}}_{0}$.
@@
 It enables the generation of diverse models by applying different weighted combinations of experts to align with specific user preferences.
-5.2 Training Strategy
+5.2. Training Strategy
 This section outlines the training approach for the parameters (denoted $\bm{\phi}$}) of the network structures introduced in Section 5.1.
@@
 The training objective can be written as:
-$\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~{}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$ (40)
-where $({\bm{x}},{\bm{y}})$’s are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot,\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$.
-In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted.
-Below, we summarize the algorithms adopted by the existing methods.
+(39) $\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$
+[…]
   • Scalarization: As outlined in Section 2.2, scalarization combines multiple objectives into one scalar objective.
-Linear scalarization is common in most algorithms, such as PHN [52], COSMOS [57], PAMAL [60], and LORPMAN [61].
-Alternatives like Tchebycheff and smooth Tchebycheff scalarization [18] can also be adopted.
+Linear scalarization is common in most algorithms, such as PHN (Navon et al., 2021), COSMOS (Ruchte and Grabocka, 2021), PAMAL (Dimitriadis et al., 2023), and LORPMAN (Chen and Kwok, 2024a).
+Alternatives like Tchebycheff and smooth Tchebycheff scalarization (Lin et al., 2024b) can also be adopted.
   • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization.
-PHN [52] employs the Exact Pareto Optimal (EPO) solver [37] to find Pareto-optimal solutions aligned with preferences.
-CPMTL [53] uses a constrained approach inspired by PMTL [36].
+PHN (Navon et al., 2021) employs the Exact Pareto Optimal (EPO) solver (Mahapatra and Rajan, 2020) to find Pareto-optimal solutions aligned with preferences.
+CPMTL (Lin et al., 2020) uses a constrained approach inspired by PMTL (Lin et al., 2019b).
   • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can optimize both diversity and convergence of a solution set.
-PHN-HVI [54] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$.
+PHN-HVI (Hoang et al., 2023) leverages hypervolume maximization to optimize the structural parameters $\bm{\phi}$}.
 In each iteration, it samples a set of preference vectors and generates solutions based on these vectors.
 Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
-5.3 Discussions
+5.3. Discussions
 Selecting the appropriate structural framework and optimization algorithms is important for effectively learning an infinite Pareto set.
@@
 Additionally, to achieve HV-optimal solutions, HV-based optimization criteria in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
-6 Theories
-6.1 Convergence of Gradient-Balancing Methods in Section 3.2
+6. Theories
+6.1. Convergence of Gradient-Balancing Methods in Section 3.2
 This section reviews the convergence analysis for gradient-balancing methods in Section 3.2.
@@
 Subsequently, it is analyzed under the stochastic gradient setting.
-Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 5).
-6.1.1 Deterministic Gradient
-With the use of deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ [97], where $K$ is the number of iterations.
-This rate is comparable to that of single-objective optimization [98].
+Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 2.5).
+6.1.1. Deterministic Gradient
+With the use of deterministic gradients, it has shown that under mild conditions, MGDA (Sener and Koltun, 2018) can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ (Fliege et al., 2019), where $K$ is the number of iterations.
+This rate is comparable to that of single-objective optimization (Nesterov, 2013).
 The fundamental idea is to assess the reduction in each individual objective function.
-When the current solution is not Pareto-stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization [97].
+When the current solution is not Pareto-stationary, it can be shown that the common descent direction $\bm{d}$} identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization (Fliege et al., 2019).
 Therefore, the method used in single-objective optimization can be applied to prove the convergence.
-Furthermore, other algorithms such as CAGrad [19], Nash-MTL [27], and PCGrad [33] also offer convergence analyses in similar approaches.
-TABLE II: Convergence of existing stochastic gradient MOO algorithms.
+Furthermore, other algorithms such as CAGrad (Liu et al., 2021b), Nash-MTL (Navon et al., 2022), and PCGrad (Yu et al., 2020a) also offer convergence analyses in similar approaches.
+6.1.2. Stochastic Gradient
+Liu and Vicente (Liu and Vicente, 2021) are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients.
+They propose a stochastic version of MGDA (Sener and Koltun, 2018) that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions.
+However, due to the inherent bias of common descent direction $\bm{d}$} introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence.
+This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
+Table 2.
+Convergence of existing stochastic gradient MOO algorithms.
 “LS” and “GS” denote the objectives are $L$-smooth and generalized $L$-smooth, respectively.
@@
 Method Batch Size Assumption Complexity Bounded CA
-SMG [20] ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
-CR-MOGM [21] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
-MoCo [22] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
-PSMGD [23] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
-MoDo [24] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
-SDMGrad [25] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
-SGSMGrad [26] ${\mathcal{O}}(1)$ GS ${\mathcal{O}}(\epsilon^{-2})$
-6.1.2 Stochastic Gradient
-Liu and Vicente [20] are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients.
-They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions.
-However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence.
-This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
+SMG (Liu and Vicente, 2021) ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
+CR-MOGM (Zhou et al., 2022) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
+MoCo (Fernando et al., 2023) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
+PSMGD (Xu et al., 2025b) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
+MoDo (Chen et al., 2023a) ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
+SDMGrad (Xiao et al., 2023) ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
+SGSMGrad (Zhang et al., 2025) ${\mathcal{O}}(1)$ GS ${\mathcal{O}}(\epsilon^{-2})$
 To overcome this limitation, Zhou et al.
-[21] propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM).
+(Zhou et al., 2022) propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM).
 It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient $\bm{\lambda}$}.
 In CR-MOGM, ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
-${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$ (41)
+(40) ${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$
 where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration.
@@
 Xu et al.
-[23] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM.
-They also match the convergence rate in stochastic single-objective optimization [99].
-Another approach to address gradient bias is MoCo [22], which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
-$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right),$ (42)
+(Xu et al., 2025b) further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM.
+They also match the convergence rate in stochastic single-objective optimization (Ghadimi and Lan, 2013).
+Another approach to address gradient bias is MoCo (Fernando et al., 2023), which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
+(41) $\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right)$,
 where $\prod_{L_{i}}$ is the projection to the set $\{{\bm{g}}\in\mathbb{R}^{d}~~|~~\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$.
@@
 To address these limitations, Chen et al.
-[24] propose the Multi-objective gradient with double sampling algorithm (MoDo).
+(Chen et al., 2023a) propose the Multi-objective gradient with double sampling algorithm (MoDo).
 MoDo mitigates bias in stochastic gradient-based MOO methods without assuming bounded function values.
@@
 Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
-${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$ (43)
+(42) ${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$
 where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ is the projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$.
@@
 Similarly, Xiao et al.
-[25] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives).
+(Xiao et al., 2023) propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives).
 In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated as:
-${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta\left[{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})^{\top}\left({\bm{G}}^{(k)}({\bm{z}}_{2}^{(k)}){\bm{\lambda}}^{(k)}\right.\right.\right.$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\left.\left.+\gamma{\bm{g}}_{0}({\bm{z}}^{(k)}_{2})\right)\right]\right),$ (44)
+[…]
 where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$.
@@
 Zhang et al.
-[26] further consider the convergence under generalized $L$-smoothness [100] and without bounded gradient assumption.
-A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table II.
-6.2 Generalization
+(Zhang et al., 2025) further consider the convergence under generalized $L$-smoothness (Li et al., 2023) and without bounded gradient assumption.
+A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table 2.
+6.2. Generalization
 The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties.
 Cortes et al.
-[101] analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations.
+(Cortes et al., 2020) analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations.
 Sukenik et al.
-[102] consider a broader class of scalarization methods.
+(Súken´ık and Lampert, 2024) consider a broader class of scalarization methods.
 They show that the generalization bounds for individual objectives extend to MOO with scalarization.
@@
 Formally, given a set of $m$ distributions $\{{\mathcal{D}}_{i}\}_{i=1}^{m}$ and a hypothesis class $\mathcal{H}$}, the goal is to find a (possibly randomized) hypothesis $h$ such that
-$\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$ (45)
+(44) $\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$
 where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss of hypothesis $h$ on ${\mathcal{D}}_{i}$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance.
 It is first formulated by Awasthi et al.
-[103] as an open problem in 2023 and has since been addressed by several works [104, 105].
+(Awasthi et al., 2023) as an open problem in 2023 and has since been addressed by several works (Peng, 2024; Zhang et al., 2024c).
 Haghtalab et al.
 are the first to show the sample complexity lower bound of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class $\mathcal{H}$}.
-By using a boosting framework, Peng [104] gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal.
+By using a boosting framework, Peng (Peng, 2024) gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal.
 In a concurrent work, Zhang et al.
-[105] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
+(Zhang et al., 2024c) provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
 In the online setting, Liu et al.
-[106] provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations.
-Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ [105].
+(Liu et al., 2024c) provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations.
+Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ (Zhang et al., 2024c).
 It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
-MoDo [24] introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
-7 Applications
-7.1 Reinforcement Learning
-Reinforcement Learning (RL) [107] involves training agents to make sequential decisions by maximizing cumulative rewards.
-Multi-Objective Reinforcement Learning (MORL) [108] extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$.
-MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
-$\min_{{\bm{\theta}}}{\bm{f}}({\bm{\theta}}):=\mathbb{E}_{a\sim\pi_{\bm{\theta}}(s)}\sum_{t=1}^{\infty}\gamma^{t}{\bm{r}}(s_{t},a_{t}),$ (46)
+MoDo (Chen et al., 2023a) introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
+7. Applications
+7.1. Reinforcement Learning
+Reinforcement Learning (RL) (Wang et al., 2024d) involves training agents to make sequential decisions by maximizing cumulative rewards.
+Multi-Objective Reinforcement Learning (MORL) (Felten et al., 2024) extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$.
+MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following objective:
+(45) $\min_{{\bm{\theta}}}{\bm{f}}({\bm{\theta}}):=\mathbb{E}_{a\sim\pi_{\bm{\theta}}(s)}\sum_{t=1}^{\infty}\gamma^{t}{\bm{r}}(s_{t},a_{t}),$
 where $t$ is the time step and $\gamma^{t}$ is the discount factor.
-Radius Algorithm (RA) [109] employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem.
+Radius Algorithm (RA) (Parisi et al., 2014) employs linear scalarization with fixed preference vectors to transform a MORL problem into a single-objective RL problem.
 However, a uniform distribution of preferences does not always lead to a uniform distribution of solutions.
 To address this, various methods focus on improving the configuration of preference vectors.
-For instance, Optimal Linear Support (OLS) [110, 111] adds new preference vectors dynamically during each iteration.
-Prediction-guided MORL (PGMORL) [112] employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
+For instance, Optimal Linear Support (OLS) (Roijers, 2016; Roijers et al., 2014) adds new preference vectors dynamically during each iteration.
+Prediction-guided MORL (PGMORL) (Xu et al., 2020) employs a hyperbolic model to map preference vectors to objective values, periodically updating preference vectors to optimize both the hypervolume and a diversity indicator.
 Although Lu et al.
-[5] demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization.
-CAPQL [113] introduces a scalarization method that combines linear scalarization with an entropy term.
-PPA [114] incorporates a term to align objective values more closely with preference vectors.
-Several studies [115, 116, 117, 118, 119, 120] have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies.
+(Lu et al., 2023) demonstrated that linear scalarization can recover the entire PS when the policy space includes all stochastic policies, this assumption may be too strong in practice, necessitating approaches beyond naive linear scalarization.
+CAPQL (Hu et al., 2023) introduces a scalarization method that combines linear scalarization with an entropy term.
+PPA (Kyriakis and Deshmukh, 2022) incorporates a term to align objective values more closely with preference vectors.
+Several studies (Peng and Fain, 2023; Van Moffaert et al., 2013b; Bai et al., 2022; Van Moffaert and Nowé, 2014; Van Moffaert et al., 2013a; Wang, 2024) have explored other scalarization functions, such as the Tchebycheff function, to identify Pareto-optimal policies.
 Beyond scalarization, gradient balancing methods (discussed in Section 3.2) also demonstrate promising performance in MORL.
-Another research direction focuses on learning the entire Pareto set [121, 122, 123, 124, 125, 126].
+Another research direction focuses on learning the entire Pareto set (Pirotta et al., 2015; Lin et al., 2022a; Basaklar et al., 2023; Zhu et al., 2023a; Shu et al., 2024; Liu et al., 2025b).
 In these approaches, the policy network $\pi_{\bm{\theta}}(s,{\bm{\alpha}})$ is conditioned on both the state $s$ and preference vector $\bm{\alpha}$}.
 Most of these methods employ training strategies similar to those described in Section 5.2, but with an RL-specific training objective.
-7.2 Bayesian Optimization
-Bayesian Optimization (BO) [127] is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model.
-When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) [128].
+7.2. Bayesian Optimization
+Bayesian Optimization (BO) (Shahriari et al., 2015) is a framework for optimizing expensive-to-evaluate black-box functions by using a probabilistic surrogate model.
+When extended to handle multiple black-box objectives simultaneously, this approach is known as Multi-Objective Bayesian Optimization (MOBO) (Knowles, 2006).
 In MOBO, acquisition functions are designed to manage the balance between exploration and exploitation across multiple objectives.
-Based on this, MOBO methods can be classified into three main categories:
-  1. 1.
-    Scalarization-based Methods.
-These methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 [128, 129].
-Zhang and Golovin [130] propose using the hypervolume indicator for scalarization.
-Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO [131], PAL [132, 133], and MOBO-RS [134] extend single-objective acquisitions to multi-objective optimization.
-  2. 2.
-    Hypervolume Improvement-Based Methods.
-These methods extend Expected Improvement (EI)  [135] to Expected Hypervolume Improvement (EHVI) [136], which selects candidate points to enhance the current solution set’s hypervolume.
-Many algorithms are proposed to improve the efficiency and allow parallel candidate generation  [137, 138, 139, 140, 141, 142, 143, 144].
-  3. 3.
-    Information Theoretic-Based Methods.
-These methods aim to maximize information gain about the PF by selecting candidate points.
-PESMO [145] reduces uncertainty in the posterior over the Pareto optimal input set.
-Some extensions [146, 147, 148, 149] refine this by considering output space.
-USeMO[150] offers an alternative uncertainty measure based on hyper-rectangle volumes.
+Based on this, MOBO methods can be classified into three main categories: (i) Scalarization-based methods convert MOO problems into single-objective ones using scalarization functions discussed in Section 2.2 (Knowles, 2006; Zhang et al., 2009).
+Zhang and Golovin (Zhang and Golovin, 2020) propose using the hypervolume indicator for scalarization.
+Instead of scalarizing first and then applying single-objective acquisition functions, methods like SMS-EGO (Ponweiser et al., 2008), PAL (Zuluaga et al., 2013, 2016), and MOBO-RS (Paria et al., 2020) extend single-objective acquisitions to multi-objective optimization.
+(ii) Hypervolume improvement-dased methods extend Expected Improvement (EI)  (Jones et al., 1998) to Expected Hypervolume Improvement (EHVI) (Emmerich et al., 2006), which selects candidate points to enhance the current solution set’s hypervolume.
+Many algorithms are proposed to improve the efficiency and allow parallel candidate generation  (Yang et al., 2019a; Daulton et al., 2020; Yang et al., 2019b; Konakovic Lukovic et al., 2020; Daulton et al., 2021, 2022, 2023; Ament et al., 2023).
+(iii) Information theoretic-based methods aim to maximize information gain about the PF by selecting candidate points.
+PESMO (Hernández-Lobato et al., 2016) reduces uncertainty in the posterior over the Pareto optimal input set.
+Some extensions (Belakaria et al., 2019; Suzuki et al., 2020; Tu et al., 2022; Hvarfner et al., 2022) refine this by considering output space.
+USeMO(Belakaria et al., 2020) offers an alternative uncertainty measure based on hyper-rectangle volumes.
 The solutions obtained by the aforementioned methods provide discrete approximations of the Pareto set.
 Lin et al.
-[151] propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1).
-Subsequent improvements have been proposed to leverage diffusion models [152] and Stein variational gradient descent [153] for efficient solution generation, as well as incorporating data augmentation for Pareto set model building [154].
-7.3 Computer Vision
-The most representative application of MOO in computer vision is multi-task dense prediction [4], which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving [155, 156].
+(Lin et al., 2022b) propose learning a continuous approximation of the Pareto set in MOBO using a hypernetwork (introduced in Section 5.1.1).
+Subsequent improvements have been proposed to leverage diffusion models (Li et al., 2024a) and Stein variational gradient descent (Nguyen et al., 2024) for efficient solution generation, as well as incorporating data augmentation for Pareto set model building (Lu et al., 2024).
+7.3. Computer Vision
+The most representative application of MOO in computer vision is multi-task dense prediction (Vandenhende et al., 2021), which aims to train a model for simultaneously dealing with multiple dense prediction tasks (such as semantic segmentation, monocular depth estimation, and surface normal estimation) and has been successfully applied in autonomous driving (Ishihara et al., 2021; Liang et al., 2023).
 Since the encoder in computer vision usually contains large number of parameters, sharing the encoder across different tasks can significantly reduce the computational cost but may cause conflicts among tasks, leading to a performance drop in some of the tasks.
-Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses [157, 158, 159, 160, 161].
-A comprehensive survey on multi-task dense prediction is in [4].
-Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud [162, 163, 164], medical image denoising [165], and pose estimation [166].
-7.4 Neural Architecture Search
-Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently [167].
+Hence, many methods are proposed to address the issue from the perspective of MOO, which use scalarization to balance multiple losses (Xu et al., 2022; Ye and Xu, 2022; Lin et al., 2024a, 2025a; Li et al., 2024b).
+A comprehensive survey on multi-task dense prediction is in (Vandenhende et al., 2021).
+Besides the dense prediction task, MOO is also useful in other computer vision tasks such as point cloud (Xie et al., 2023, 2024; Wang et al., 2024b), medical image denoising (Kyung et al., 2024), and pose estimation (Ye et al., 2024c).
+7.4. Neural Architecture Search
+Neural Architecture Search (NAS), which aims to design the architecture of neural networks automatically, has gained significant interest recently (Ren et al., 2021).
 Due to the complex application scenarios in the real world, recent works consider multiple objectives beyond just accuracy.
 For example, to search an efficient architecture for deployment on platforms with limited resources, many studies incorporate resource-constraint objectives such as the parameter size, FLOPs, and latency.
-These works are mainly based on gradient-based NAS methods (like DARTS [168]) and formulate it as a multi-objective optimization problem.
-Among them, [8, 169, 170, 171, 172] employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights.
+These works are mainly based on gradient-based NAS methods (like DARTS (Liu et al., 2019b)) and formulate it as a multi-objective optimization problem.
+Among them, (Wu et al., 2019; Cai et al., 2019; Wu et al., 2021; Wang et al., 2021a; Yue et al., 2022) employ scalarization to identify a single solution, where the task and efficiency objectives are combined with fixed weights.
 However, altering the weights of the objectives needs a complete rerun of the search, which is computationally intensive.
 Therefore, Sukthanker et al.
-[173] propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
-7.5 Recommendation System
-Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others [174].
+(Sukthanker et al., 2025) propose to learn a mapping from a preference vector to an architecture using a hypernetwork, enabling to provide the entire PS without the need for a new search.
+7.5. Recommendation Systems
+Beyond just focusing on accuracy, recommendation systems often incorporate additional quality metrics such as novelty, diversity, serendipity, popularity, and others (Zheng and Wang, 2022).
 These factors naturally frame the problem as a multi-objective optimization optimization.
-Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives [175, 176, 177].
-In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model [178, 179, 180, 181].
-Additionally, some research efforts aim to identify finite Pareto sets [182] and infinite Pareto sets [183, 184], enabling the provision of personalized recommender models that cater to varying user preferences.
-7.6 Large Language Models (LLMs)
+Traditionally, many methods have employed scalarization techniques to assign weights to the different objectives (Di Noia et al., 2017; Patil et al., 2021; Lacerda, 2017).
+In recent developments, gradient-balancing approaches are applied to address multi-objective optimization within recommendation systems to learn a single recommendation model (Milojkovic et al., 2019; Lin et al., 2019a; Mitrevski et al., 2020; Yang et al., 2024a).
+Additionally, some research efforts aim to identify finite Pareto sets (Xie et al., 2021) and infinite Pareto sets (Ge et al., 2022; Wilm et al., 2024), enabling the provision of personalized recommender models that cater to varying user preferences.
+7.6. Large Language Models (LLMs)
 Recently, there has been a growing trend of incorporating multi-objective optimization into the training of large language models (LLMs), including multi-task fine-tuning and multi-objective alignment.
@@
 This approach separates fine-tuning on each task, causing extensive costs in training and difficulties in deployment.
-MFTCoder [185] proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks.
-CoBa [186] studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
+MFTCoder (Liu et al., 2024a) proposes a multi-task fine-tuning framework that enhances the coding capabilities of LLMs by addressing data imbalance, varying difficulty levels, and inconsistent convergence speeds across tasks.
+CoBa (Gong et al., 2024) studies multi-task fine-tuning for LLMs and balances task convergence with minimal computational overhead.
 Aligning with multi-dimensional human preferences (such as helpfulness, harmfulness, humor, and conciseness) is essential for customizing responses to users’ needs, as users typically have diverse preferences for different aspects.
-MORLHF [187] and MODPO [188] train an LLM for every preference configuration by linearly combining multiple (implicit) reward models.
-To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters [189, 190] or output logits [191] to deal with different preference requirements.
+MORLHF (Wu et al., 2023) and MODPO (Zhou et al., 2024a) train an LLM for every preference configuration by linearly combining multiple (implicit) reward models.
+To avoid retraining, some methods train multiple LLMs separately for each preference dimension and merge their parameters (Rame et al., 2023; Jang et al., 2023) or output logits (Shi et al., 2024) to deal with different preference requirements.
 However, these methods need to train and store multiple LLMs, leading to huge computational and storage costs.
-To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts [192, 193, 194] or model parameters [63].
-7.7 Miscellaneous
-In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning [195, 196], federated learning [197, 198, 199], long-tailed learning [200, 201, 202], continual learning [203], diffusion models [204, 205, 206, 207], neural combinatorial optimization [208, 122, 209, 210, 211], GFlowNets [212, 213], and physics informed neural networks (PINNs) [214, 215].
+To improve efficiency, several preference-aware methods are proposed to fine-tune a single LLM for varying preferences by incorporating the relevant coefficients into the input prompts (Wang et al., 2024c; Guo et al., 2024; Yang et al., 2024b) or model parameters (Zhong et al., 2024; Lin et al., 2025b).
+7.7. Miscellaneous
+In addition to the applications mentioned above, multi-objective optimization has been applied to a variety of other deep learning scenarios, including meta-learning (Yu et al., 2023; Wang et al., 2021c), federated learning (Hu et al., 2022; Askin et al., 2024; Kang et al., 2024), long-tailed learning (Li et al., 2024c; Zhao et al., 2024; Zhou et al., 2024b), continual learning (Lai et al., 2025), diffusion models (Hang et al., 2023; Go et al., 2023; Xu et al., 2025a; Yao et al., 2024), neural combinatorial optimization (Li et al., 2020; Lin et al., 2022a; Wang et al., 2023a; Chen et al., 2023c, b), GFlowNets (Jain et al., 2023; Zhu et al., 2023b), and physics informed neural networks (PINNs) (Hwang and Lim, 2024; Liu et al., 2025a).
 These applications highlight the versatility of multi-objective optimization in addressing diverse challenges within deep learning.
-8 Resources
-In this section, we introduce some benchmark datasets and open-source libraries for gradient-based MOO in deep learning. Benchmark datasets are summarized in Table III, with details in Appendix C.
-Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL [216] and LibMOON [^2] https://github.com/xzhang2523/libmoon [217], provide unified environments for implementing and fairly evaluating MOO methods.
-Both are built on PyTorch [218] and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios.
-LibMTL [216] mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets.
-LibMOON [217] mainly focuses on exploring the whole Pareto set.
+Table 3. Summary of benchmark datasets in multi-objective deep learning.
+Dataset Description #Objectives #Samples
+NYUv2 (Silberman et al., 2012) indoor scene understanding 33 $1,449$
+Taskonomy (Zamir et al., 2018) indoor scene understanding $26$ $\approx 4$M
+Cityscapes (Cordts et al., 2016) urban scene understanding 22 $3,475$
+QM9 (Ramakrishnan et al., 2014) molecular property prediction $11$ 130130K
+Office-31 (Saenko et al., 2010) image classification 33 $4,110$
+Office-Home (Venkateswara et al., 2017) image classification 44 $15,500$
+CelebA (Liu et al., 2015) image classification $40$ $\approx 202$K
+CIFAR-100 (Krizhevsky and Hinton, 2009) image classification 2020 6060K
+MT10/MT50 (Yu et al., 2020b) reinforcement learning $10/50$ -
+XTREME (Hu et al., 2020) multilingual learning 99 $\approx 597$K
+PKU-SafeRLHF (Ji et al., 2024) two-dimensional preference data 22 $\approx 82$K
+UltraFeedBack (Cui et al., 2023) multi-dimensional preference data 44 $\approx 64$K
+HelpSteer2 (Wang et al., 2024a) multi-attributes preference data 55 $\approx 21$K
+8. Resources
+In this section, we introduce widely-used benchmark datasets and open-source libraries for gradient-based MOO in deep learning.
+8.1. Datasets
+Below, we provide details on the benchmark datasets, which are summarized in Table 3:
+  • NYUv2 dataset (Silberman et al., 2012) is for indoor scene understanding.
+It has three tasks (i.e., semantic segmentation, depth estimation, and surface normal prediction) with $795$ training and $654$ testing samples.
+  • Taskonomy dataset (Zamir et al., 2018) is obtained from 3D scans of about $600$ buildings and contains $26$ tasks (such as semantic segmentation, depth estimation, surface normal prediction, keypoint detection, and edge detection) with about $4$ million samples.
+  • Cityscapes dataset (Cordts et al., 2016) is for urban scene understanding. It has two tasks (i.e., semantic segmentation and depth estimation) with $2,975$ training and $500$ testing samples.
+  • QM9 dataset (Ramakrishnan et al., 2014) is for molecular property prediction with $11$ tasks. Each task performs regression on one property. It contains $130,000$ samples.
+  • Office-31 dataset (Saenko et al., 2010) contains $4,110$ images from three domains (tasks): Amazon, DSLR, and Webcam. Each task has 31 classes.
+  • Office-Home dataset (Venkateswara et al., 2017) contains $15,500$ images from four domains (tasks): artistic images, clipart, product images, and real-world images.
+Each task has $65$ object categories collected under office and home settings.
+  • CelebA (Liu et al., 2015) is a large-scale face attribute dataset comprising $202,599$ face images.
+Each image is annotated with $40$ binary attributes, resulting in $40$ distinct binary classification tasks.
+  • CIFAR-100 (Krizhevsky and Hinton, 2009) is an image classification dataset with $100$ classes, containing 5050K training and 1010K testing images.
+These $100$ classes are divided into $20$ tasks, each of which is a $5$-class image classification problem.
+  • MT10 and MT50 from Meta-World benchmark (Yu et al., 2020b) are two multi-task reinforcement learning settings, containing $10$ and $50$ robot manipulation tasks, respectively.
+  • XTREME benchmark (Hu et al., 2020) aims to evaluate the cross-lingual generalization abilities of multilingual representations.
+It contains $9$ tasks in $40$ languages, including $2$ classification tasks, $2$ structure prediction tasks, $3$ question-answering tasks, and $2$ sentence retrieval tasks.
+  • PKU-SafeRLHF dataset (Ji et al., 2024) focuses on safety alignment in LLMs, containing $82,118$ question-answering pairs with the annotations of helpfulness and harmlessness.
+  • UltraFeedBack dataset (Cui et al., 2023) contains $63,967$ prompts, each corresponding to four responses from different LLMs.
+Each response has $4$ aspects of annotations, namely instruction-following, truthfulness, honesty, and helpfulness, generated by GPT-4.
+  • HelpSteer2 dataset (Wang et al., 2024a) contains $21,362$ samples.
+Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 0 and $4$ where higher means better for each attribute.
+8.2. Libraries
+Two popular libraries, LibMTL [^1] https://github.com/median-research-group/LibMTL (Lin and Zhang, 2023) and LibMOON [^2] https://github.com/xzhang2523/libmoon (Zhang et al., 2024d), provide unified environments for implementing and fairly evaluating MOO methods.
+Both are built on PyTorch (Paszke et al., 2019) and feature modular designs, enabling flexible development of new methods or application of existing ones to new scenarios.
+LibMTL (Lin and Zhang, 2023) mainly focuses on finding a single Pareto-optimal solution and supports $24$ methods introduced in Section 3 and $6$ benchmark datasets.
+LibMOON (Zhang et al., 2024d) mainly focuses on exploring the whole Pareto set.
 It includes over 20 methods for obtaining a finite set of solutions (introduced in Section 4) or infinite set of solutions (introduced in Section 5).
-TABLE III: Summary of benchmark datasets in multi-objective deep learning.
-Dataset Description #Objectives #Samples
-NYUv2 [219] indoor scene understanding $3$ $1,449$
-Taskonomy [220] indoor scene understanding $26$ $\approx 4$M
-Cityscapes [221] urban scene understanding $2$ $3,475$
-QM9 [222] molecular property prediction $11$ $130$K
-Office-31 [223] image classification $3$ $4,110$
-Office-Home [224] image classification $4$ $15,500$
-CelebA [225] image classification $40$ $\approx 202$K
-CIFAR-100 [226] image classification $20$ $60$K
-MT10/MT50 [227] reinforcement learning $10/50$
-XTREME [228] multilingual learning $9$ $\approx 597$K
-PKU-SafeRLHF [229] two-dimensional preference data $2$ $\approx 82$K
-UltraFeedBack [230] multi-dimensional preference data $4$ $\approx 64$K
-HelpSteer2 [231] multi-attributes preference data $5$ $\approx 21$K
-9 Challenges and Future Directions
+9. Challenges and Future Directions
 Despite significant progress in applying MOO in deep learning, both in learning a single model and in learning a Pareto set of models, several challenges remain.
-Theoretical Understanding. While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
+Theoretical Understanding.
+While practical methods for MOO in deep learning have seen significant progress, their theoretical foundations remain relatively underexplored.
 Most existing research focuses on analyzing the convergence of MOO algorithms to stationary points, while generalization error, which is critical for evaluating performance on unseen data, has received less attention.
 For instance, Chen et al.
-[24], as discussed in Section 6.2, provide the algorithm-dependent generalization analysis.
+(Chen et al., 2023a), as discussed in Section 6.2, provide the algorithm-dependent generalization analysis.
 Extending this to a broader range of algorithms could offer deeper insights into how different techniques affect generalization.
@@
 Zhang et al.
-[232] established a generalization bound for HV-based Pareto set learning.
+(Zhang et al., 2023a) established a generalization bound for HV-based Pareto set learning.
 However, the effect of various network architectures (Section 5.1) on approximating Pareto sets remain unclear.
@@
 Leveraging MOO methods to effectively address these trade-offs presents an opportunity for further exploration.
-10 Conclusion
-In this paper, we provide the first comprehensive review of gradient-based MOO in deep learning, covering algorithms, theoretical foundations, applications, benchmark datasets, and software.
-By unifying diverse approaches and systematically identifying key challenges, this paper serves as a foundational resource to drive future advancements in this important and evolving field.
-Appendix A Illustration of Key Concepts in MOO
-Figure 6 illustrates the Pareto concepts on the two-objective problem.
-The blue circles represent Pareto optimal solutions (defined in Definition 2), which indicates that no solution can improve one objective without worsening the other.
-The blue curve connecting the blue circles denotes the Pareto front, as defined in Definition 3.
-The yellow circles indicate weak Pareto optimal solutions since they can be improved in one objective without negatively impacting the other.
-For example, comparing ${\bm{\theta}}_{2}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{2}$ is a weak Pareto optimal solution since it can be improved in the second objective without affecting the first.
-The red circles represent solutions that are not Pareto optimal because there exists another solution that strictly Pareto dominates them.
-For instance, comparing ${\bm{\theta}}_{3}$ and ${\bm{\theta}}_{1}$, ${\bm{\theta}}_{3}$ is not a Pareto optimal solution as ${\bm{\theta}}_{1}$ outperforms it in both objectives.
-Figure 6: Illustration of Pareto optimality concepts.
-The blue, yellow, and red circles denote Pareto optimal solutions, weakly Pareto optimal solutions, and dominated solutions, respectively.
-The blue curve represents the Pareto front.
-Appendix B Additional Details and Further Results for the Experiment in Fig. 4
-To demonstrate the differences in the solution distributions produced by various methods, we consider a commonly used benchmark problem, LSMOP1 [84].
-Each algorithm is run for 5000 iterations.
-The solutions obtained by PTML [36], EPO [37, 38], GradHV [47], and MOO-SVGD [48] are shown in Figure 4 in the main paper.
-Additionally, the solutions generated by MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18] are shown in Figure 7.
-From Figures 4 and 7, we can observe that MGDA, which does not incorporate preferences, exhibits poor diversity.
-The constraint-based approach in PTML produces solutions that are better aligned with the preference vector but still fall short of exact alignment.
-EPO and PMGDA achieves exact alignment with the preference vector.
-Some methods that do not use a preference vector, such as GradHV and MOO-SVGD, are also able to produce diverse solutions.
-Figure 7: Distributions of solutions obtained by using MGDA [6], PMGDA [40], UMOD [43], and Smooth Tchebycheff [18].
-The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
-Appendix C Details of Benchmark Datasets
-In this section, we list the details of benchmark datasets in Section 8.
-  • NYUv2 dataset [219] is for indoor scene understanding.
-It has three tasks (i.e., semantic segmentation, depth estimation, and surface normal prediction) with $795$ training and $654$ testing samples.
-  • Taskonomy dataset [220] is obtained from 3D scans of about $600$ buildings and contains $26$ tasks (such as semantic segmentation, depth estimation, surface normal prediction, keypoint detection, and edge detection) with about $4$ million samples.
-  • Cityscapes dataset [221] is for urban scene understanding. It has two tasks (i.e., semantic segmentation and depth estimation) with $2,975$ training and $500$ testing samples.
-  • QM9 dataset [222] is for molecular property prediction with $11$ tasks. Each task performs regression on one property. It contains $130,000$ samples.
-  • Office-31 dataset [223] contains $4,110$ images from three domains (tasks): Amazon, DSLR, and Webcam. Each task has 31 classes.
-  • Office-Home dataset [224] contains $15,500$ images from four domains (tasks): artistic images, clipart, product images, and real-world images.
-Each task has $65$ object categories collected under office and home settings.
-  • CelebA dataset [225] is a large-scale face attributes dataset, containing $202,599$ face images.
-Each image has $40$ attribute annotation.
-Hence, this dataset includes $40$ tasks and each of them is a binary classification problem for one attribute.
-  • CIFAR-100 dataset [226] an image classification with $100$ classes, containing $50,000$ training and $10,000$ testing image samples.
-These $100$ classes are divided into $20$ tasks, each of which is a $5$-class image classification problem.
-  • MT10 and MT50 from Meta-World benchmark [227] are two multi-task reinforcement learning settings, containing $10$ and $50$ robot manipulation tasks, respectively.
-  • XTREME benchmark [228] aims to evaluate the cross-lingual generalization abilities of multilingual representations.
-It contains $9$ tasks in $40$ languages, including $2$ classification tasks, $2$ structure prediction tasks, $3$ question-answering tasks, and $2$ sentence retrieval tasks.
-  • PKU-SafeRLHF dataset [229] focuses on safety alignment in LLMs, containing $82,118$ question-answering pairs with the annotations of helpfulness and harmlessness.
-  • UltraFeedBack dataset [230] contains $63,967$ prompts, each corresponding to four responses from different LLMs.
-Each response has $4$ aspects of annotations, namely instruction-following, truthfulness, honesty, and helpfulness, generated by GPT-4.
-  • HelpSteer2 dataset [231] contains $21,362$ samples.
-Each sample contains a prompt and a response with $5$ human-annotated attributes (i.e., helpfulness, correctness, coherence, complexity, and verbosity), each ranging between 00 and $4$ where higher means better for each attribute.
+10. Conclusion
+In this paper, we provide the first comprehensive review of gradient-based multi-objective deep learning, a field of growing importance as models are required to balance multiple, often conflicting, objectives.
+We have systematically surveyed the spectrum of algorithms, covering methods for finding a single balanced model, a finite set of models, and an entire infinite Pareto set of models.
+The review also delves into the theoretical foundations of these methods, summarizing key results in convergence while highlighting the need for a deeper understanding of generalization.
+The practical significance of MOO is demonstrated across diverse applications, including its emerging role in Large Language Models (LLMs).
+By unifying these approaches and identifying key challenges, this paper serves as a foundational resource to guide and inspire future advancements in this critical and rapidly evolving domain.